### v2

# footnotetext: MSC2020: 53A04, 53A70 (primary), 53A17, 70B15, 37K25, 37K10, 35Q53 (secondary)

Abstract.
We consider the configuration space of points on the two-dimensional sphere that satisfy a specific system of quadratic equations. We construct periodic orbits in this configuration space using elliptic theta functions and show that they satisfy semi-discrete analogues of mKdV and sine-Gordon equations. The configuration space we investigate corresponds to the state space of a linkage mechanism known as the Kaleidocycle, and the constructed orbits describe the characteristic motion of the Kaleidocycle. Our approach is founded on the relationship between the deformation of spatial curves and integrable systems, offering an intriguing example where an integrable system generates an orbit in the space of real solutions to polynomial equations defined by geometric constraints.
The second and third-named authors were partially supported by KAKENHI Grant-in-Aid for Scientific Research (C) 21K3229. All authors are partially supported by JST CREST Grant No. JPMJCR1911.
1. Introduction
Given a simple finite graph $(V,E)$ with edge weights $\ell:E\to\mathbb{R}$, its $d$-dimensional realisation is a map $\phi:V\to\mathbb{R}^{d}$ that satisfies the condition $|\phi(u)-\phi(v)|^{2}=\ell(uv)^{2}$ for all edges $uv\in E$. This mapping embeds the graph into $d$-dimensional Euclidean space while preserving the prescribed edge lengths. The collection of all possible realisations forms a real algebraic set in $\mathbb{R}^{d|V|}$, known as the configuration space of $(V,E,\ell)$. This space has been the subject of extensive research across various domains, including algebraic geometry, robotics, and mechanism theory [10, 21, 25]. The primary motivation for this work is rooted in the analysis of the motion of a particular class of linkage mechanisms called Kaleidocycles (Fig. 1 (Left)). These mechanisms consist of $k$ identical copies of equifacial tetrahedra connected by their opposing edges as hinge joints, forming a ring. A state of a Kaleidocycle corresponds to a realisation of the graph formed by the vertices and edges of the tetrahedra (Eq. (2.1)), where $\ell$ is determined by the shape of the tetrahedron. Its motion, disregarding self-collision, is described by a path in the configuration space. Although this object has been studied by several authors [8, 28, 30], even its existence beyond simple classical cases has not been clarified. Our work addresses this gap by providing an explicit construction of the object and its motion.
Our key observation is to associate each state of a Kaleidocycle with a closed spatial polygonal curve. This curve is formed by connecting the midpoints of the hinge edges. The angles between the hinge edges are reminiscent of the torsion angle of a curve, and the turning angles are reminiscent of the curvature angle (§2.1). In this geometric formulation using spatial curves, the motion of a Kaleidocycle is modelled by a segment-length and torsion-angle preserving deformation of the corresponding closed polygonal curve of a constant torsion angle. We can then identify the configuration space of a given Kaleidocycle with a subspace of a direct product of two-spheres (2.2). Note that it is not clear if there exists $\ell$ (the shape of the tetrahedra) for a given $k$ such that the configuration space is non-empty; that is, if a Kaleidocycle exists. We construct a one-dimensional subspace of the configuration space using elliptic theta functions, proving that there is a Kaleidocycle for any $k\geq 6$ that can move. The idea is to compare the functional identities of elliptic theta functions with the compatibility conditions of the framing of a discrete spatial curve.
It was known in [19] that the semi-discrete modified KdV (mKdV) and the sine-Gordon equations [19] are defined on the configuration space (see §3.2). Therefore, a one-dimensional orbit in the configuration space is generated through the time evolution defined by the system of ordinary differential equations. The main result of this paper can also be viewed as providing a solution to the system through an explicit construction based on the $\tau$ functions of the two-component KP hierarchy [11]. We show that our construction of the motion of the Kaleidocycle (§5) indeed satisfies the system of the ordinary differential equations described in [19]. The sweeping surface of the motion corresponds to a semi-discrete analogue of the discrete constant negative curvature surface (K-surface) studied in [2, 16].
Our work connects seemingly distant fields of integrable systems and kinematics of linkage via an interesting object, Kaleidocycle. It provides an interesting example, where a flow generated by an integrable system on a real-algebraic set is explicitly integrated using elliptic theta functions.
The structure of the paper is as follows. Section 2 introduces the definition of Kaleidocycles. As essential invariants to describe the state of a Kaleidocycle, the signed curvature angle and torsion angle are defined together with the Frenet frame. The state of a Kaleidocycle is identified with a solution to a system of quadratic equations corresponding to the geometric constraints on the binormal vectors. Section 3 reviews the connection between the theory of curves and integrable systems. The construction of discrete spatial curves and their Frenet frames in terms of the $\tau$ functions is presented. A previous result that describes the motion of a Kaleidocycle using semi-discrete integrable equations is recalled. Section 4 constructs a set of $\tau$ functions with three parameters $v,r,y$ using elliptic theta functions. The resulting curve is shown to have a constant torsion angle. A condition for the resulting curve to be closed is also discussed; for a given $k$, the appropriate choice of parameters yields a Kaleidocycle. By introducing an additional parameter, a family of time-parametrised Kaleidocycles is constructed in Section 5. We then study the time evolution equations satisfied by the vertex positions and curvature angle of the Kaleidocycles and show connections to previous research. Section 6 provides examples of Kaleidocycles and corresponding semi-discrete surfaces with figures. We present two conjectures that are suggested by numerical experiments. The Kaleidocycles constructed in this paper seem to be Möbius Kaleidocycles (Conjecture 6.1).
Most of the proofs of our theorems and propositions are deferred to the appendix since they require technically-involved computations regarding elliptic theta functions.
2. Algebraic and geometric models of Kaleidocycle
2.1. Configuration space of Kaleidocycle
A Kaleidocycle consists of (typically six) identical copies of equifacial tetrahedra connected by their opposing edges as hinge joints to form a ring. It can be folded from a sheet of paper and makes an example of rigid origami; it rotates like a bubble ring while keeping all faces rigid. The origin of Kaleidocycle is obscure but one of the first appearances in the literature is in Bricard’s paper [3], where he considers an essentially equivalent closed kinematic chain, now known as the Bricard 6R. For nearly a hundred years, the geometry and kinematics of this object have attracted the attention of researchers (see [29] and the references therein).
We first define a generalisation of Kaleidocycle following [18]. We consider $k$-copies of an equifacial tetrahedron in $\mathbb{R}^{3}$ connected by opposing edges to form a ring. Denote the vertex set by $\{0,1,\ldots,2k-1\}$ such that the vertices $(2n,2n+1)$ form a hinge edge and the vertices $(2n,2n+1,2n+2)$ form one of the triangle faces. Let $l_{01},l_{12},l_{20}$ be the lengths of the sides of the triangle face. A state of a $k$-Kaleidocycles of type $(l_{01},l_{12},l_{20})$ is identified by the real solution $(x_{0},x_{1},\ldots,x_{2k-1})\in(\mathbb{R}^{3})^{2k}$ of the quadratic equations
(2.1) $|x_{2n}-x_{2n+1}|^{2}$ $=l^{2}_{01},$
$|x_{2n}-x_{2n+2}|^{2}$ $=|x_{2n+1}-x_{2n+3}|^{2}=l^{2}_{02},$
$|x_{2n+1}-x_{2n+2}|^{2}$ $=|x_{2n}-x_{2n+3}|^{2}=l^{2}_{12}.$
Note that collision between tetrahedra is ignored. The solution space of (2.1) admits an action of the Euclidean transformation group $E(3)$ of $\mathbb{R}^{3}$, and the expected dimension of the quotient is $63\cdot 2k-5k-\dim(E(3))=k-$ 6. This simple counting estimate is often called Maxwell’s degree-of-freedom counting law [8]. The Bricard 6R is a case of $k=6$. Its configuration space has dimension one, which is greater than the expected dimension of 00, giving an example of an overconstrained  [8]. When $k\geq 7$, it is numerically observed that the configuration space degenerates to a circle when the angle between adjacent hinge edges is set to a critical value [20].Some of the findings of [20] are also published in [30]. For $k>7$, this class of Kaleidocycles, named the Möbius Kaleidocycles, provides a rare example of underconstrained linkage mechanisms whose configuration space has less dimension than the expected one.
The state of a Kaleidocycle can be equivalently described by a spatial polygon or a closed discrete spatial curve. Let $\gamma_{n}=(x_{2n}+x_{2n+1})/2\ (0\leq n<k-1)$ be the mid-point of the hinge edge. We set $\gamma_{k}=\gamma_{0}$. Since the tetrahedron is equifacial, $\gamma_{n+1}-\gamma_{n}$ is orthogonal to both $\widetilde{B}_{n}$ and $\widetilde{B}_{n+1}$, where $\widetilde{B}_{n}=\dfrac{x_{2n+1}-x_{2n}}{|x_{2n+1}-x_{2n}|}$ is the hinge direction. By reordering the vertices, we can assume the inner product $\langle\widetilde{B}_{n-1},\widetilde{B}_{n}\rangle$ is constant for $1\leq n<k$. Denote this constant by $\cos\lambda$ with $0<\lambda<\pi$, which we call the torsion angle. We have two cases $\langle\widetilde{B}_{k-1},\widetilde{B}_{0}\rangle=\pm\cos\lambda$. We say the Kaleidocycle is oriented (resp. non-oriented) when the value is $+\cos\lambda$ (resp. $-\cos\lambda$). We set $\widetilde{B}_{k}=\pm\widetilde{B}_{0}$ so that $\langle\widetilde{B}_{k-1},\widetilde{B}_{k}\rangle=\cos\lambda$. Furthermore, since $\widetilde{B}_{n+1}\times\widetilde{B}_{n}$ is parallel to $\gamma_{n+1}-\gamma_{n}$, we have $\sum_{n=0}^{k-1}(\widetilde{B}_{n+1}\times\widetilde{B}_{n})=\sum_{n=0}^{k-1}(\gamma_{n+1}-\gamma_{n})=0$. Conversely, the vertex coordinates $x_{n}$ can be reconstructed from $\{\gamma_{n}\}$, $\{\widetilde{B}_{n}\}$ and $(l_{01},l_{12},l_{20})$ up to translation. Therefore,
Definition 1.
The configuration space of the $k$-Kaleidocycles with the torsion angle $\lambda$ is the set $\{(\widetilde{B}_{0},\ldots,\widetilde{B}_{k})\in(S^{2})^{k+1}\}$ satisfying the following conditions:
(2.2) $\langle\widetilde{B}_{n-1},\widetilde{B}_{n}\rangle$ $=\cos\lambda,\quad(1\leq n\leq k)$
$\sum_{n=0}^{k-1}(\widetilde{B}_{n+1}\times\widetilde{B}_{n})$ $=0,$
$\widetilde{B}_{k}$ =±$=\pm\widetilde{B}_{0}.$
A motion of a Kaleidocycle is identified with a one-dimensional (time-parametrised) orbit in the space defined by (2.1) or (2.2).
Figure 1. (Left) The classical 6-Kaleidocycle. (Right) Kaleidocycle as a discrete curve with its framing.
2.2. Discrete spatial curve of a constant torsion angle
A discrete spatial curve with non-vanishing torsion angles can be described in terms of binormals. Let $\widetilde{B}:\mathbb{Z}\rightarrow S^{2}$ satisfy $\langle\widetilde{B}_{n},\widetilde{B}_{n-1}\rangle\neq 0$ for all $n\in\mathbb{Z}$. A discrete spatial curve of unit segment-length having $\widetilde{B}$ as the binormals can be constructed as follows. Define the tangent $T:\mathbb{Z}\rightarrow S^{2}$ and the normal $\widetilde{N}:\mathbb{Z}\rightarrow S^{2}$ by
(2.3) $T_{n}=\frac{\widetilde{B}_{n+1}\times\widetilde{B}_{n}}{\left|\widetilde{B}_{n+1}\times\widetilde{B}_{n}\right|},\quad\widetilde{N}_{n}=\widetilde{B}_{n}\times T_{n}.$
Furthermore, given $\gamma_{0}\in\mathbb{R}^{3}$, define $\gamma:\mathbb{Z}\to\mathbb{R}^{3}$ by
(2.4) $\gamma_{n+1}=\gamma_{n}+T_{n}=\gamma_{n}+\frac{\widetilde{B}_{n+1}\times\widetilde{B}_{n}}{\left|\widetilde{B}_{n+1}\times\widetilde{B}_{n}\right|}.$
The signed curvature angle $\kappa$ and the torsion angle are defined by
(2.5) $\begin{split}\langle T_{n},T_{n-1}\rangle&=\cos\kappa_{n},\quad-\pi<\kappa_{n}<\pi,\\ \langle\widetilde{B}_{n},\widetilde{B}_{n-1}\rangle&=\cos\lambda_{n},\quad 0<\lambda_{n}<\pi.\end{split}$
The discrete Frenet frame $\Phi_{n}=[T_{n},\widetilde{N}_{n},\widetilde{B}_{n}]\in\mathrm{SO}(3)$ satisfies the discrete Frenet-Serret formula
(2.6) $\Phi_{n+1}=\Phi_{n}L_{n},\quad L_{n}=M_{1}(-\lambda_{n+1})M_{3}(\kappa_{n+1}),$
where
(2.7) $M_{1}(\theta)=\left[\begin{array}[]{ccc}1&0&0\\ 0&\cos\theta&-\sin\theta\\ 0&\sin\theta&\cos\theta\\ \end{array}\right],\quad M_{3}(\theta)=\left[\begin{array}[]{ccc}\cos\theta&-\sin\theta&0\\ \sin\theta&\cos\theta&0\\ 0&0&1\end{array}\right].$
Remark 2.1.
Note that our Frenet frame coincides with the usual one only up to sign. We adopt this convention to avoid discontinuous changes in the framing during the motion of the Kaleidocycle; its motion passes the state when the curvature vanishes.
As in the case of a smooth curve, a discrete spatial curve of a unit segment length is determined by its torsion and curvature angles. Therefore, the motion of a discrete curve of a constant torsion angle is determined by the set of functions $\kappa_{n}(t)$, which can be freely chosen. The fact that a Kaleidocycle is closed enforces interesting topological constraints. For example, the self-linking number must be a half-integer, and hence, stay constant during the motion. By Călugăreanu-Fuller-White’s theorem [1], the writhing number of the curve should stay constant as well. The self-linking number also tells that the configuration space of a Kaleidocycle is not generally connected.
3. Curves and integrable systems
There are deep connections between differential geometry and the theory of integrable systems, where the integrable differential or difference equations arise as compatibility conditions of the geometric objects (see, for example, [27]). In the context of the deformation of spatial and plane curves, the Frenet frame and its deformation are described by the system of linear partial differential equations. The compatibility condition yields so-called the Ablowitz-Kaup-Newell-Segur (AKNS) hierarchy for the curvature and the torsion, including the nonlinear Schrödinger (NLS) equation as a typical example. In particular, if the torsion is a constant, the integrable deformation is governed by the mKdV equation and its hierarchy [9, 23, 27]. The connections are also investigated in a discrete setting, where curves are discretised as piecewise linear (polygonal) curves. For example, continuous deformations of discrete plane and spatial curves have been considered in [5], where the deformations are governed by the semi-discrete analogue of the mKdV and the NLS equations. Discrete deformations of the discrete plane and spatial curves have been studied in [16]. In principle, for a given solution to the nonlinear differential/difference equation for the curvature and torsion, one can reconstruct the curve by solving the associated system of linear differential/difference equations for the Frenet frame and integrating the tangent vector.
The theta functions naturally appear in so-called the quasi-periodic solutions to the nonlinear integrable equations, and the corresponding solutions to the associated linear problem are nothing but the Baker-Akhiezer functions. Reconstruction of the deformation of the spatial curve with constant torsion described by the theta function solutions to the mKdV equation has been studied in this framework in [4, 24]. Also, an explicit formula of the spatial curves in terms of the $\tau$ functions of the two-component KP hierarchy based on Hirota’s bilinear formalism has been presented in [11].
Since the dynamics of the Kaleidocycles are modelled by the integrable deformation of the closed discrete spatial curves with a constant torsion angle which preserves the segment-length and torsion angle [19], we study the theta function solutions to the corresponding integrable semi-discrete equations, where the theta functions are nothing but the $\tau$ functions. Therefore we construct the explicit formula of the Kaleidocycles based on the method of direct reconstruction of the discrete spatial curves in terms of the $\tau$ functions.
Hereafter, a (discrete) curve always indicates a polygonal spatial curve.
3.1. Expression of a curve by the $\tau$ functions
We construct a discrete curve $\gamma_{n}$ and its Frenet frame $\Phi_{n}=[T_{n},\widetilde{N}_{n},\widetilde{B}_{n}]$ indexed by $n\in\mathbb{Z}$ using a modification of the $\tau$ functions presented in [11]. Let $F_{n},\widetilde{G}_{n}:\mathbb{C}\to\mathbb{R}$ and $f_{n},g_{n},G_{n},H_{n}:\mathbb{C}\to\mathbb{C}$ be functions indexed by $n\in\mathbb{Z}$ with a complex variable $z$ that satisfy the following equations
(3.1) $f_{n}f_{n}^{*}+g_{n}g_{n}^{*}=F_{n+1}F_{n}$
(3.2) $\beta F_{n}f_{n}+G_{n}g_{n}^{*}=F_{n+1}f_{n-1}$
(3.3) $\beta F_{n}g_{n}-G_{n}f_{n}^{*}=F_{n+1}g_{n-1}$
(3.4) $H_{n+1}F_{n}-H_{n}F_{n+1}=f_{n}^{*}g_{n}$
(3.5) $D_{z}F_{n+1}\cdot F_{n}:=\frac{dF_{n+1}}{dz}F_{n}-F_{n+1}\frac{dF_{n}}{dz}=g_{n}g_{n}^{*}$
(3.6) $\left|G_{n}\right|^{2}=\widetilde{G}_{n}^{2},$
where $\beta\in\mathbb{R}$ is some constant and $D_{z}$ is Hirota’s bilinear differential operator [12]. These functions are called the $\tau$ functions. Let $\varrho_{n}=\dfrac{1}{2}\dfrac{G_{n}}{\widetilde{G}_{n}}$ and
(3.10) $T_{n}$ $=\left.\frac{1}{f_{n}f_{n}^{*}+g_{n}g_{n}^{*}}\left(\begin{array}[]{c}\vspace{2mm}f_{n}^{*}g_{n}+f_{n}g_{n}^{*}\\ \vspace{2mm}\frac{1}{i}\left(f_{n}^{*}g_{n}-f_{n}g_{n}^{*}\right)\\ f_{n}f_{n}^{*}-g_{n}g_{n}^{*}\end{array}\right)\right|_{z=0}$
(3.14) $\widetilde{N}_{n}$ $=\left.\frac{1}{f_{n}f_{n}^{*}+g_{n}g_{n}^{*}}\left(\begin{array}[]{c}\vspace{2mm}\left\{(f_{n}^{*})^{2}-(g_{n}^{*})^{2}\right\}\varrho_{n}+\left\{(f_{n})^{2}-(g_{n})^{2}\right\}\varrho_{n}^{*}\\ \vspace{2mm}\frac{1}{i}\left[\left\{(f_{n}^{*})^{2}+(g_{n}^{*})^{2}\right\}\varrho_{n}-\left\{(f_{n})^{2}+(g_{n})^{2}\right\}\varrho_{n}^{*}\right]\\ -2\left(f_{n}^{*}g_{n}^{*}\varrho_{n}+f_{n}g_{n}\varrho_{n}^{*}\right)\end{array}\right)\right|_{z=0}$
(3.18) $\widetilde{B}_{n}$ $=\left.\frac{1}{f_{n}f_{n}^{*}+g_{n}g_{n}^{*}}\left(\begin{array}[]{c}\vspace{2mm}\frac{1}{i}\left[-\left\{(f_{n}^{*})^{2}-(g_{n}^{*})^{2}\right\}\varrho_{n}+\left\{(f_{n})^{2}-(g_{n})^{2}\right\}\varrho_{n}^{*}\right]\\ \vspace{2mm}\left\{(f_{n}^{*})^{2}+(g_{n}^{*})^{2}\right\}\varrho_{n}+\left\{(f_{n})^{2}+(g_{n})^{2}\right\}\varrho_{n}^{*}\\ \frac{2}{i}\left(f_{n}^{*}g_{n}^{*}\varrho_{n}-f_{n}g_{n}\varrho_{n}^{*}\right)\end{array}\right)\right|_{z=0}$
(3.22) $\gamma_{n}$ $=\left.\left(\begin{array}[]{c}\vspace{2mm}\dfrac{H_{n}+H_{n}^{*}}{F_{n}}\\ \vspace{2mm}\dfrac{1}{i}\dfrac{H_{n}-H_{n}^{*}}{F_{n}}\\ n-2\dfrac{\partial\log F_{n}}{\partial z}\end{array}\right)\right|_{z=0}.$
Similarly to [11, §3], we obtain:
Proposition 3.1.
$\Phi_{n}=[T_{n},\widetilde{N}_{n},\widetilde{B}_{n}]$ defines the discrete Frenet frame for the curve $\gamma$ satisfying (2.6). Its curvature and torsion angles are given by
(3.23) $\lambda_{n}$ $=\left.\frac{1}{i}\log\left(\frac{G_{n}^{*}G_{n-1}}{\widetilde{G}_{n}\widetilde{G}_{n-1}}\right)\right|_{z=0}=\left.\frac{1}{2i}\log\left(\frac{G_{n}^{*}G_{n-1}}{G_{n}G_{n-1}^{*}}\right)\right|_{z=0},$
$\kappa_{n}$ $=\left.2\arctan\left(\frac{\widetilde{G}_{n}}{\beta F_{n}}\right)\right|_{z=0}.$
3.2. Motion of Kaleidocycle on the osculating plane
Let $\gamma$ be a Kaleidocycle, that is, a closed curve with a constant torsion angle and a unit segment length. A motion of a Kaleidocycle is identified with a time-parametrised family $\gamma:\mathbb{Z}\times\mathbb{R}_{\geq 0}\to\mathbb{R}^{3}$ such that $\gamma(t)$ is a Kaleidocycle for all $t\geq 0$. In [19], a particular motion of a Kaleidocycle satisfying the following conditions is considered:
(3.24) $\left\langle\dot{\gamma}_{n},\widetilde{B}_{n}\right\rangle=0,\quad\left\langle\dot{\gamma}_{n},\dot{\gamma}_{n}\right\rangle=\tilde{\rho}^{2},\quad|{\gamma}_{n+1}-{\gamma}_{n}|=1\quad(0\leq\forall n\leq k),$
where $\tilde{\rho}>0$ is a constant and $\dot{\gamma}_{n}=\dfrac{d\gamma_{n}}{dt}$. These conditions mean that the velocity of each vertex is restricted within the osculating plane and has a constant magnitude. It is shown in [19] that the Frenet-Serret formula (2.6) together with (3.24) demand that there exists a function $\Theta_{n}$, called the potential function, with $\kappa_{n}=\dfrac{\Theta_{n+1}-\Theta_{n-1}}{2}$ that satisfies either the semi-discrete potential mKdV equation:
(3.25) $\frac{d}{dt}\left(\frac{\Theta_{n+1}+\Theta_{n}}{2}\right)=\left(1+\cos\lambda\right)\tilde{\rho}\sin\left(\frac{\Theta_{n+1}-\Theta_{n}}{2}\right),$
or the semi-discrete sine-Gordon equation
(3.26) $\frac{d}{dt}\left(\frac{\Theta_{n+1}-\Theta_{n}}{2}\right)=\left(1-\cos\lambda\right)\tilde{\rho}\sin\left(\frac{\Theta_{n+1}+\Theta_{n}}{2}\right).$
The motions corresponding to these equations differ in general. However, the explicit construction discussed in §5 satisfies both equations simultaneously.
Remark 3.2.
As noted in §2.1, the dimension of the configurations of a Kaleidocycle defined by (2.2) is generally $k-6$, and hence, (3.24) may look too restrictive. However, the Möbius Kaleidocycles, which exhibit critical torsion angles, the dimension of the configurations is conjectured to be one regardless of $k$. An extensive numerical simulation suggests that the motion of the Möbius Kaleidocycle satisfies (2.2). Therefore, (2.2) specifies a particular yet essential motion of Kaleidocycles.
4. An explicit formula of closed discrete curve with a constant torsion angle in terms of elliptic theta functions
Based on the representation of curves discussed in §3.1, we define certain $\tau$ functions satisfying (3.1)-(3.6) using elliptic theta functions to construct a family of discrete curves with a constant torsion angle. Then, we give sufficient conditions for the curve to be closed and derive an explicit expression for the motion of Kaleidocycles.
4.1. A family of open curves with a constant torsion angle
Throughout this section, $j$ runs through $\{1,2,3,4\}$. Fix $w\in\mathbb{C}$ with $\mathrm{Im}(w)>0$. The elliptic theta functions $\vartheta_{j}:\mathbb{C}\to\mathbb{C}$, are defined as follows [26, 22]:
(4.1) $\vartheta_{1}(v;w)$ $=-\sum_{n=-\infty}^{\infty}e^{\pi iw\left(n+\frac{1}{2}\right)^{2}+2\pi i\left(n+\frac{1}{2}\right)\left(v+\frac{1}{2}\right)},$
$\vartheta_{2}(v;w)$ $=\sum_{n=-\infty}^{\infty}e^{\pi iw\left(n+\frac{1}{2}\right)^{2}+2\pi i\left(n+\frac{1}{2}\right)v},$
$\vartheta_{3}(v;w)$ $=\sum_{n=-\infty}^{\infty}e^{\pi iwn^{2}+2\pi inv},$
$\vartheta_{4}(v;w)$ $=\sum_{n=-\infty}^{\infty}e^{\pi iwn^{2}+2\pi in\left(v+\frac{1}{2}\right)}.$
We often denote $\vartheta_{j}(v;w)$ by $\vartheta_{j}(v)$ when $w$ is fixed.
We construct the $\tau$ functions satisfying (3.1)-(3.6) using these elliptic theta functions such that the torsion angle given by Proposition 3.1 is a constant. The following observations are useful insights for constructing solutions. In order to make $F$ a real-valued function, it is convenient to use a property of elliptic theta functions, shown in Lemma K.1. Therefore, $w$ is taken to be a pure imaginary number. Additionally, to make $\gamma$ a closed curve, it is convenient to use the pseudo-period of elliptic theta functions (A.12), and therefore the space variable is taken to be a pure imaginary number. A real parameter $r$ is introduced to prevent the curve from being a planar curve. Then, the solution to satisfy (3.1)-(3.6) can be constructed from the identities of elliptic theta functions (e.g. (A.2)), and the result is as shown in Theorem 4.1. For $x\in\mathbb{C}$ and $y>0$, we define $d_{j}(x;y)=\frac{\vartheta^{{}^{\prime}}_{j}(x;iy)}{\vartheta_{j}(x;iy)}$, where ${}^{{}^{\prime}}$ denotes the derivative with respect to the first variable of elliptic theta functions. For $y>0$, $r\in\mathbb{R}\setminus{(1/2\mathbb{Z})}$ and $v\in\mathbb{R}\setminus{(y\mathbb{Z})}$, we define
(4.2) $\alpha_{2}=\sqrt{\vartheta_{3}(2r;iy)\vartheta_{3}(0;iy)},\quad\beta=\frac{\vartheta_{3}(iv;iy)}{\vartheta_{3}(0;iy)},\quad s_{2}=\vartheta_{3}(2r;iy)\vartheta_{1}(iv;iy),$
$u_{2}=\vartheta_{1}\left(-\frac{1}{2}iv+r;iy\right)\vartheta_{3}\left(-\frac{1}{2}iv+r;iy\right)\vartheta_{1}\left(\frac{1}{2}iv+r;iy\right)\vartheta_{3}\left(\frac{1}{2}iv+r;iy\right),$
$\Delta_{j}=d_{j}\left(-\frac{1}{2}iv+r;y\right)-d_{j}\left(\frac{1}{2}iv+r;y\right),$
$R_{j}=\frac{\vartheta_{j}\left(-\frac{1}{2}iv+r;iy\right)}{\vartheta_{j}\left(\frac{1}{2}iv+r;iy\right)}.$
Noting $\vartheta_{j}(z;iy)^{*}=\vartheta_{j}(z^{*};iy)$ when $y>0$, we have
(4.3) $\begin{split}&\alpha_{2}^{*}=\alpha_{2},\quad\beta^{*}=\beta,\quad s_{2}^{*}=-s_{2},\quad u_{2}^{*}=u_{2},\\ &\Delta_{j}^{*}=-\Delta_{j},\quad R_{j}R_{j}^{*}=1.\end{split}$
Theorem 4.1.
Put $\mu_{n}=ivn+\frac{z}{\Delta_{3}-\Delta_{1}}$ and define
(4.4) $F_{n}(z)=\alpha_{2}\exp\left(\frac{n\Delta_{3}}{\Delta_{3}-\Delta_{1}}z\right)\vartheta_{2}\left(\mu_{n}-\frac{1}{2}iv\right),$
$f_{n}(z)=\vartheta_{3}\left(-\frac{1}{2}iv+r\right)R_{3}^{n}\exp\left(\frac{\left(n+\frac{1}{2}\right)\Delta_{3}}{\Delta_{3}-\Delta_{1}}z\right)\vartheta_{2}\left(\mu_{n}+r\right),$
$g_{n}(z)=\vartheta_{1}\left(\frac{1}{2}iv+r\right)R_{1}^{-n}\exp\left(\frac{\left(n+\frac{1}{2}\right)\Delta_{3}}{\Delta_{3}-\Delta_{1}}z\right)\vartheta_{4}\left(\mu_{n}-r\right),$
$G_{n}(z)=\alpha_{2}\frac{\vartheta_{1}(iv)}{\vartheta_{3}(0)}R_{1}^{-n}R_{3}^{n}\exp\left(\frac{n\Delta_{3}}{\Delta_{3}-\Delta_{1}}z\right)\vartheta_{4}\left(\mu_{n}-\frac{1}{2}iv\right),$
$H_{n}(z)=\frac{u_{2}}{\alpha_{2}s_{2}}R_{1}^{-n}R_{3}^{-n}\exp\left(\frac{n\Delta_{3}}{\Delta_{3}-\Delta_{1}}z\right)\vartheta_{4}\left(\mu_{n}-\frac{1}{2}iv-2r\right),$
$\widetilde{G}_{n}(z)=-i\alpha_{2}\frac{\vartheta_{1}(iv)}{\vartheta_{3}(0)}\exp\left(\frac{n\Delta_{3}}{\Delta_{3}-\Delta_{1}}z\right)\vartheta_{4}\left(\mu_{n}-\frac{1}{2}iv\right).$
Then, they satisfy (3.1)-(3.6) to form $\tau$ functions. The corresponding discrete curve $\gamma$ defined by (3.10) has a unit segment length $\left|\gamma_{n+1}-\gamma_{n}\right|=1$. The torsion angle is given by
(4.5) $\lambda_{n}=\frac{1}{i}\log\left(\frac{G_{n}^{*}(0)G_{n-1}(0)}{\widetilde{G}_{n}(0)\widetilde{G}_{n-1}(0)}\right)=\frac{1}{i}\log\left(R_{1}R_{3}^{-1}\right).$
In particular, this does not depend on $n$, and $\gamma$ has a constant torsion angle $\lambda_{n}=\lambda_{0}$. Moreover, the signed curvature angle is given by
(4.6) $\begin{split}&\kappa_{n}=2\arctan\left(\frac{\widetilde{G}_{n}(0)}{\beta F_{n}(0)}\right)\\ &=2\arctan\left(-i\frac{\vartheta_{1}(iv)\vartheta_{4}\left(ivn-\frac{1}{2}iv\right)}{\vartheta_{3}(iv)\vartheta_{2}\left(ivn-\frac{1}{2}iv\right)}\right),\end{split}$
where we take the primary value for $\arctan$ in $(-\pi/2,\pi/2)$.
A proof of Theorem 4.1 is given in Appendix B.
4.2. Sufficient conditions for closedness
The constant torsion-angle and unit segment-length curves given in Theorem 4.1 are not necessarily closed. We derive a sufficient condition on the parameters $v,r,y$ for the curve $\gamma$ to be closed to define a Kaleidocycle. We say a curve $\gamma$ is closed with a period $k\in\mathbb{N}$ when $\gamma_{n+k}=\gamma_{n}$ for all $n$. Rewriting this condition using the $\tau$ functions, we obtain
(4.7) $\frac{H_{n+k}}{F_{n+k}}=\frac{H_{n}}{F_{n}},$
$n+k-2\displaystyle\frac{\partial\log F_{n+k}}{\partial z}=n-2\displaystyle\frac{\partial\log F_{n}}{\partial z}.$
It is difficult to give general solutions to (4.7), but we can utilise the pseudo-period of elliptic theta functions (A.12) to obtain particular solutions. For a fixed $k$, choose $m\in\mathbb{Z}\setminus{\{0\}}$ and set $v=my/k$. Then, from (A.12), we see
(4.8) $\frac{H_{n+k}}{F_{n+k}}=\frac{H_{n}}{F_{n}}\left(R_{1}R_{3}\right)^{-k}\exp\left(4mr\pi i\right)(-1)^{m},$
$F_{n+k}=F_{n}\exp\left(k\frac{\Delta_{3}}{\Delta_{3}-\Delta_{1}}z\right)\exp\left[-\pi i\left\{2m\left(\mu_{n}-\frac{1}{2}iv\right)+m^{2}w\right\}\right],$
where $\mu_{n}=ivn+\frac{z}{\Delta_{3}-\Delta_{1}}$. Using (4.8), (4.7) can be rewritten as
(4.9) $\Lambda:=\exp\left\{\pi im\left(4r+1\right)\right\}\left(R_{1}R_{3}\right)^{-k}-1=0,$
$\Xi:=\frac{\left(\Delta_{3}+\Delta_{1}\right)k-4\pi im}{\Delta_{3}-\Delta_{1}}=0.$
Note that since
(4.10) $\frac{\partial}{\partial r}\log(\Lambda+1)=4\pi im-k\left(\Delta_{3}+\Delta_{1}\right)=-\left(\Delta_{3}-\Delta_{1}\right)\Xi$
holds, (4.9) can also be written as
(4.11) $\Lambda=0,\quad\frac{\partial}{\partial r}\Lambda=0.$
From the above discussion we obtain the following theorem.
Theorem 4.2.
For a fixed $k$, let $m\in\mathbb{Z}\setminus{\{0\}}$, $y>0$, $r\in\mathbb{R}\setminus{(1/2\mathbb{Z})}$ and $v\in\mathbb{R}\setminus{(y\mathbb{Z})}$ satisfy $kv=my$ and (4.9). Then, the curve defined by Theorem 4.1 is closed and forms a $k$-Kaleidocycle with a torsion angle given by (4.5).
Proof.
The closure condition (4.7) holds for any $n\in\mathbb{Z}$ when the parameters $v,r,y$ satisfy $kv=my$ and (4.9). Thus, the conditions $kv=my$ and (4.9) give sufficient conditions for the curve given by Theorem 4.1 to be a closed curve. ∎
Hereafter, we assume that the parameters $v,r,y$ are taken to satisfy the sufficient conditions for closedness.
Example 4.3.
The classical 6-Kaleidocycle (Figure 1 Left) with $\lambda=\pi/2$ (e.g., see [8]) is obtained by $(v,r,y,m)\approx(0.640,0.250,1.279,3)$.
Remark 4.4.
By using (4.9) and (A.12), we see that $\widetilde{B}_{0}=(-1)^{m}\widetilde{B}_{k}$. From this, we also see that $m$ and $Lk$ have the same parity. Furthermore, under the transformation $\widetilde{B}_{n}\rightarrow(-1)^{n+1}\widetilde{B}_{n}$, $\widetilde{N}_{n}\rightarrow(-1)^{n+1}\widetilde{N}_{n}$, the signed curvature angle becomes $(-1)^{n+1}\kappa_{n}$ and the torsion angle becomes $\pi+\lambda$. This indicates that this transformation corresponds to the one described in Proposition 5.16.
5. Construction of the motion of a Kaleidocycle
By introducing the deformation parameter to the curve constructed in §4.1, we construct an explicit formula for the motion of a Kaleidocycle in Theorem 5.1. We then derive the evolution equations for the vertex positions of the curve and the (potential of the) curvature. Connections to the previously-known isoperimetric deformation of a curve by the semi-discrete mKdV equation [16] and by the semi-discrete potential mKdV equation and the semi-discrete sine-Gordon equation [19] are discussed and summarised in Table 2.
5.1. Explicit formula of the discrete curve with deformation parameter
We introduce the deformation parameter so that only the signed curvature angle changes with a constant torsion angle and while satisfying (3.1)-(3.6). Additionally, the real parameters $C$ and $\Gamma$ are introduced to control the global motion of the curve. The following theorem describes a Kaleidocycle together with its motion explicitly in terms of elliptic theta functions.
Theorem 5.1.
Denote $\mu_{n}=ivn+\frac{z}{\Delta_{3}-\Delta_{1}}$. Consider the following functions parameterised by $t\in\mathbb{R}$.
(5.1) $F_{n}(t,z)=\alpha_{2}\exp\left(\frac{n\Delta_{3}}{\Delta_{3}-\Delta_{1}}z+\frac{C}{2}tz\right)\vartheta_{2}\left(\mu_{n}-\frac{1}{2}iv+it\right),$
$f_{n}(t,z)=\vartheta_{3}\left(-\frac{1}{2}iv+r\right)R_{3}^{n}\exp\left(\frac{\left(n+\frac{1}{2}\right)\Delta_{3}}{\Delta_{3}-\Delta_{1}}z+\frac{C}{2}tz-\frac{\Gamma}{2}it\right)\vartheta_{2}\left(\mu_{n}+r+it\right),$
$g_{n}(t,z)=\vartheta_{1}\left(\frac{1}{2}iv+r\right)R_{1}^{-n}\exp\left(\frac{\left(n+\frac{1}{2}\right)\Delta_{3}}{\Delta_{3}-\Delta_{1}}z+\frac{C}{2}tz+\frac{\Gamma}{2}it\right)\vartheta_{4}\left(\mu_{n}-r+it\right),$
$G_{n}(t,z)=\alpha_{2}\frac{\vartheta_{1}(iv)}{\vartheta_{3}(0)}R_{1}^{-n}R_{3}^{n}\exp\left(\frac{n\Delta_{3}}{\Delta_{3}-\Delta_{1}}z+\frac{C}{2}tz\right)\vartheta_{4}\left(\mu_{n}-\frac{1}{2}iv+it\right),$
$H_{n}(t,z)=\frac{u_{2}}{\alpha_{2}s_{2}}R_{1}^{-n}R_{3}^{-n}\exp\left(\frac{n\Delta_{3}}{\Delta_{3}-\Delta_{1}}z+\frac{C}{2}tz+\Gamma it\right)\vartheta_{4}\left(\mu_{n}-\frac{1}{2}iv-2r+it\right),$
$\widetilde{G}_{n}(t,z)=-i\alpha_{2}\frac{\vartheta_{1}(iv)}{\vartheta_{3}(0)}\exp\left(\frac{n\Delta_{3}}{\Delta_{3}-\Delta_{1}}z+\frac{C}{2}tz\right)\vartheta_{4}\left(\mu_{n}-\frac{1}{2}iv+it\right),$
where $C,\Gamma\in\mathbb{R}$ are constants. Then, they satisfy (3.1)-(3.6) to form $\tau$ functions. For each $t$, the corresponding discrete curve $\gamma$ (3.10) has a unit segment length $\left|\gamma_{n+1}-\gamma_{n}\right|=1$. The torsion angle is given by
(5.2) $\lambda_{n}=\frac{1}{i}\log\left(\frac{G_{n}^{*}(t,0)G_{n-1}(t,0)}{\widetilde{G}_{n}(t,0)\widetilde{G}_{n-1}(t,0)}\right)=\frac{1}{i}\log\left(R_{1}R_{3}^{-1}\right),$
In particular, this depends neither on $n$ nor on $t$, and $\gamma$ has a constant torsion angle $\lambda_{n}(t)=\lambda_{0}(0)$. Moreover, the signed curvature angle is given by
(5.3) $\begin{split}&\kappa_{n}\left(t\right)=2\arctan\left(\frac{\widetilde{G}_{n}(t,0)}{\beta F_{n}(t,0)}\right)\\ &=2\arctan\left(-i\frac{\vartheta_{1}(iv)\vartheta_{4}\left(ivn-\frac{1}{2}iv+it\right)}{\vartheta_{3}(iv)\vartheta_{2}\left(ivn-\frac{1}{2}iv+it\right)}\right),\end{split}$
where we take the primary value for $\arctan$ in $(-\pi/2,\pi/2)$. If $r,v,y$ are chosen as in Theorem 4.2, $\gamma$ is closed for every $t$.
A proof of Theorem 5.1 is given in Appendix B.
Proposition 5.2.
The parameter $C$ contributes only to the parallel translation of the curve along the $Z$-axis, whereas $\Gamma$ contributes only to the rotation of the curve around the $Z$-axis.
Proof.
For $A\in\mathbb{C}$ and $B:\mathbb{C}\to\mathbb{C}$, we have
(5.4) $\begin{split}&\left(\begin{array}[]{ccc}\cos\Gamma t&-\sin\Gamma t\\ \sin\Gamma t&\cos\Gamma t\\ \end{array}\right)\left(\begin{array}[]{ccc}A+A^{*}\\ \frac{1}{i}\left(A-A^{*}\right)\end{array}\right)\\ &=\frac{1}{2}\left(\begin{array}[]{ccc}e^{i\Gamma t}+e^{-i\Gamma t}&-\frac{1}{i}\left(e^{i\Gamma t}-e^{-i\Gamma t}\right)\\ \frac{1}{i}\left(e^{i\Gamma t}-e^{-i\Gamma t}\right)&e^{i\Gamma t}+e^{-i\Gamma t}\\ \end{array}\right)\left(\begin{array}[]{ccc}A+A^{*}\\ \frac{1}{i}\left(A-A^{*}\right)\end{array}\right)\\ &=\left(\begin{array}[]{ccc}e^{i\Gamma t}A+e^{-i\Gamma t}A^{*}\\ \frac{1}{i}\left(e^{i\Gamma t}A-e^{-i\Gamma t}A^{*}\right)\end{array}\right)\end{split}$
and
(5.5) $\begin{split}2\frac{\partial\log\left(e^{\frac{C}{2}tz}B\right)}{\partial z}=Ct+2\frac{\partial\log B}{\partial z}.\end{split}$
The assertion follows from the definition (3.10). ∎
In Theorem 5.9, we will make convenient choices of $C$ and $\Gamma$ to derive the deformation equation considered in the previous studies presented in §3.2.
Table 1 summarises the relationship between various variables. The first column lists the parameters that determine the curve $\gamma$ through the variables listed in the first row. A circle in the table indicates that the variable shown in the first row is dependent on the parameter shown in the first column.
Table 1. Dependency of variables on the parameters $v,r,y,k,m,t,C,\Gamma$.
$\gamma$ $\kappa$ $\lambda$ $\Lambda$ $\Xi$ $u_{2}$ $\beta$ $s_{2}$ $R_{j}$ $\Delta_{j}$ $\alpha_{2}$ $\beta_{1,2}$ $\epsilon_{1,2}$ $\omega$ $\delta_{1,2,3}$
$v$circ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
$r$circ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
$y$circ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
$k$circ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
$m$circ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
$t$circ $\circ$
$\Gamma$circ $\circ$
$C$circ
(3.10) (5.3) (5.2) (4.9) (4.9) (4.2) (4.2) (4.2) (4.2) (4.2) (4.2) (5.6) (5.6) (5.6) (5.6)
The constructed motion has a periodic curvature. The next proposition follows from (A.12).
Proposition 5.3.
We have $\kappa_{n}\left(t+y\right)=-\kappa_{n}\left(t\right)$. In other words, $2y$ is a period of the motion.
Remark 5.4.
There are certain relations between the curves constructed in Theorem 5.1 with different sets of parameters. The transformation of parameters $(v,m)\rightarrow(-v,-m)$ preserves the closure condition. We see that $\kappa_{n}\left(t;-v\right)=-\kappa_{n}\left(-t;v\right),\lambda(-v,r)=-\lambda(v,r)$. In what follows, We always choose $(v,m)$ or $(-v,-m)$ so that $\lambda\geq 0$.
5.2. Motion of vertex positions
The construction in Theorem 5.1 involves parameters $C$ and $\Gamma$ that contribute only to the global rigid transformation of the curve. In general, the velocity of the vertex position $\dot{\gamma_{n}}$ has a binormal component, and its magnitude is not constant as described in Theorem 5.6. However, with the right choice of $C$ and $\Gamma$, curves that satisfy (3.24) are obtained as described in Theorem 5.9.
We derive bilinear equations satisfied by the $\tau$ functions given by Theorem 5.1 to calculate $\dot{\gamma_{n}}$.
For $\Gamma\in\mathbb{R},y>0,r\in\mathbb{R}\setminus{(1/2\mathbb{Z})}$ and $v\in\mathbb{R}\setminus{(y\mathbb{Z})}$, we define
(5.6) $\omega=\frac{\vartheta_{1}\left(2r;iy\right)\vartheta_{3}\left(0;iy\right)}{\vartheta_{3}\left(2r;iy\right)\vartheta_{1}\left(iv;iy\right)},$
$\delta_{1}=\frac{\vartheta_{3}\left(\frac{1}{2}iv+r;iy\right)^{2}\vartheta_{3}\left(-\frac{1}{2}iv+r;iy\right)^{2}}{\vartheta_{3}\left(2r;iy\right)\vartheta_{3}\left(0;iy\right)^{3}},$
$\delta_{2}=\frac{\vartheta_{1}\left(\frac{1}{2}iv+r;iy\right)^{2}\vartheta_{3}\left(-\frac{1}{2}iv+r;iy\right)^{2}}{\vartheta_{1}\left(iv;iy\right)^{2}\vartheta_{3}\left(2r;iy\right)\vartheta_{3}\left(0;iy\right)},$
$\delta_{3}=\frac{-iu_{2}}{s_{2}\vartheta_{3}(0;iy)\vartheta_{1}(iv;iy)}\left[\frac{\vartheta_{1}^{{}^{\prime}}\left(\frac{1}{2}iv+r;iy\right)\vartheta_{3}\left(-\frac{1}{2}iv+r;iy\right)}{\vartheta_{1}\left(-\frac{1}{2}iv+r;iy\right)\vartheta_{3}\left(\frac{1}{2}iv+r;iy\right)}\right.$
$\left.+\frac{\vartheta_{1}\left(\frac{1}{2}iv+r;iy\right)\vartheta_{3}^{{}^{\prime}}\left(-\frac{1}{2}iv+r;iy\right)}{\vartheta_{1}\left(-\frac{1}{2}iv+r;iy\right)\vartheta_{3}\left(\frac{1}{2}iv+r;iy\right)}\right],$
$\epsilon_{1}=\frac{2i}{\Delta_{3}-\Delta_{1}}\frac{\vartheta_{3}^{{}^{\prime\prime}}(0;iy)}{\vartheta_{3}(0;iy)},$
$\epsilon_{2}=\frac{2i}{\Delta_{3}-\Delta_{1}}\frac{\vartheta_{1}^{{}^{\prime}}(0;iy)^{2}}{\vartheta_{1}(iv;iy)^{2}},$
$\beta_{1}=\frac{1}{2}\left(\delta_{3}+\delta_{3}^{*}\right)+\frac{1}{2}i\Gamma\left(\delta_{2}-\delta_{2}^{*}\right),$
$\beta_{2}=\frac{1}{2}i\left(\delta_{3}-\delta_{3}^{*}\right)-\frac{1}{2}\Gamma\left(\delta_{2}+\delta_{2}^{*}\right).$
Then, it follows that
(5.7) $\omega^{*}=-\omega,\quad\delta_{1}^{*}=\delta_{1},\quad\epsilon_{1}^{*}=\epsilon_{1},\quad\epsilon_{2}^{*}=\epsilon_{2},\quad\beta_{1}^{*}=\beta_{1},\quad\beta_{2}^{*}=\beta_{2}.$
A proof of the following proposition is given in Appendix C.
Proposition 5.5.
The $\tau$ functions (5.1) satisfy the following bilinear equations.
(5.8) $f_{n}f_{n-1}^{*}+g_{n}^{*}g_{n-1}=\beta F_{n}F_{n}+\omega G_{n}G_{n}^{*},$
(5.9) $g_{n}^{*}g_{n-1}-g_{n}g_{n-1}^{*}=f_{n}f_{n-1}^{*}-f_{n}^{*}f_{n-1}=\omega G_{n}G_{n}^{*},$
(5.10) $\delta_{2}f_{n}^{*}g_{n-1}-\delta_{2}^{*}f_{n-1}^{*}g_{n}=H_{n}F_{n},$
(5.11) $\delta_{1}F_{n}F_{n}+\delta_{2}G_{n}G_{n}^{*}=f_{n}f_{n-1}^{*},$
(5.12) $\left(\beta-\delta_{1}\right)F_{n}F_{n}+\left(\omega-\delta_{2}\right)G_{n}G_{n}^{*}=g_{n}^{*}g_{n-1},$
(5.13) $D_{t}H_{n}\cdot F_{n}=\left(\delta_{3}+i\Gamma\delta_{2}\right)f_{n}^{*}g_{n-1}+\left(\delta_{3}^{*}-i\Gamma\delta_{2}^{*}\right)f_{n-1}^{*}g_{n},$
(5.14) $D_{t}D_{z}F_{n}\cdot F_{n}=\left(\epsilon_{1}+C\right)F_{n}F_{n}+\epsilon_{2}G_{n}G_{n}^{*}.$
By using these equations, we obtain the evolution equation of the vertex positions of the curve.
Theorem 5.6.
The curve $\gamma$ constructed in Theorem 5.1 satisfies the following deformation equation.
(5.15) $\begin{split}\dot{\gamma}_{n}(t)=&2\beta\left[\beta_{1}T_{n}-\beta_{1}\tan\frac{\kappa_{n}}{2}\widetilde{N}_{n}+\beta_{2}\tan\frac{\kappa_{n}}{2}\widetilde{B}_{n}\right]\\ &+\left(\begin{array}[]{c}0\\ 0\\ 2\beta_{1}\left(\beta-2\delta_{1}\right)-\epsilon_{1}-C\end{array}\right).\end{split}$
A proof of Theorem 5.6 is given in Appendix D. This is a deformation equation similar to the one presented in [16]. Starting from (5.15), the compatibility condition of the Frenet frame yields the semi-discrete mKdV equation.
Proposition 5.7.
When the curve with a unit segment length and a constant torsion angle deforms according to the deformation equation (5.15), the following holds.
  1. (1)
    The segment length is preserved.
  2. (2)
    $\kappa$ satisfies the following semi-discrete mKdV equation:
    (5.16) $\dot{\kappa}_{n}=\left(2\beta\beta_{1}\cos\lambda+2\beta\beta_{2}\sin\lambda\right)\left(\tan\frac{\kappa_{n+1}}{2}-\tan\frac{\kappa_{n-1}}{2}\right).$
Furthermore, if the left-hand-side of the equation (5.27) does not depend on $n$, the following holds:
  1. (3)
    The torsion angle is preserved.
Proof.
The proofs of (i) and (ii) are given by the same calculation as in Section 5 of [16]. We prove (iii). From the deformation equation (5.15) and the discrete Frenet-Serret formula (2.6), we see that
(5.17) $\begin{split}&-\dot{\lambda}\sin\lambda=\left\langle\dot{\widetilde{B}}_{n},\widetilde{B}_{n+1}\right\rangle+\left\langle\widetilde{B}_{n},\dot{\widetilde{B}}_{n+1}\right\rangle\\ &=-\sin\lambda\left(2\beta\beta_{1}\sin\lambda-2\beta\beta_{2}\cos\lambda\right)\times\left(\frac{1}{\sin\kappa_{n+1}}\tan\frac{\kappa_{n+2}}{2}\right.\\ &\left.+\frac{1}{\tan\kappa_{n+1}}\tan\frac{\kappa_{n}}{2}-\frac{1}{\tan\kappa_{n}}\tan\frac{\kappa_{n+1}}{2}-\frac{1}{\sin\kappa_{n}}\tan\frac{\kappa_{n-1}}{2}\right),\end{split}$
holds, which vanishes because of (5.27). Therefore $\dot{\lambda}=0$.
Remark 5.8.
Proposition 5.7 (iii) does not hold in general without the assumption on (5.27). We can construct a motion of a Kaleidocycle satisfying (5.15) such that the left-hand side of (5.27) depends both on $n$ and $t$.
By adjusting the global motion of $\gamma$ by $C$ and $\Gamma$, the deformation equation of $\gamma$ can be made to satisfy the equation considered in the previous studies presented in §3.2.
Theorem 5.9.
Set $C,\Gamma$ to be
(5.18) $\begin{split}\Gamma_{+}=&\frac{-\left(\delta_{3}+\delta_{3}^{*}\right)+\left(\delta_{3}e^{i\lambda}+\delta_{3}^{*}e^{-i\lambda}\right)}{i\omega},\\ C_{+}=&-\epsilon_{1}-i\Gamma_{+}\left(\omega\left(2\delta_{1}-\beta\right)-\frac{\delta_{2}+\delta_{2}^{*}}{\omega}\right)\\ &-\left(\delta_{3}+\delta_{3}^{*}\right)\left(2\delta_{1}-\beta\right)+\frac{\delta_{3}-\delta_{3}^{*}}{\omega},\\ \rm or\\ \Gamma_{-}=&\frac{-\left(\delta_{3}+\delta_{3}^{*}\right)-\left(\delta_{3}e^{i\lambda}+\delta_{3}^{*}e^{-i\lambda}\right)}{i\omega},\\ C_{-}=&-\epsilon_{1}-i\Gamma_{-}\left(\omega\left(2\delta_{1}-\beta\right)-\frac{\delta_{2}+\delta_{2}^{*}}{\omega}\right)\\ &-\left(\delta_{3}+\delta_{3}^{*}\right)\left(2\delta_{1}-\beta\right)+\frac{\delta_{3}-\delta_{3}^{*}}{\omega}.\end{split}$
Then, the curve constructed in Theorem 5.1 satisfies the following deformation equations.
(5.19) $\begin{split}&\left\langle\dot{\gamma}_{n},\widetilde{B}_{n}\right\rangle=0,\quad\langle\dot{\gamma}_{n},\dot{\gamma}_{n}\rangle=\rho^{2},\end{split}$
where $\rho^{2}>0$ is given by
(5.20) $\rho^{2}=\rho_{+}^{2}:=\left(i\frac{\vartheta_{1}^{{}^{\prime}}(0)}{\vartheta_{1}(iv)}\frac{u_{2}}{\vartheta_{3}(r)^{2}\vartheta_{1}(r)^{2}}\right)^{2}$
when $(C,\Gamma)=(C_{+},\Gamma_{+})$ and
(5.21) $\rho^{2}=\rho_{-}^{2}:=\left(-i\frac{\vartheta_{1}^{{}^{\prime}}(0)}{\vartheta_{1}(iv)}\frac{u_{2}}{\vartheta_{4}(r)^{2}\vartheta_{2}(r)^{2}}\right)^{2}$
when $(C,\Gamma)=(C_{-},\Gamma_{-})$.
A proof of Theorem 5.9 is given in Appendix D.
Remark 5.10.
A discrete analogue of the constant negative curvature surface, called the discrete K-surface, is introduced in [2]. It is discussed that certain discrete-time deformation of discrete curves with a constant torsion angle generates discrete K-surfaces in [16]. Deformation equation (3.24) can be seen as a continuous-time analogue of the discrete-time deformation. Theorem 5.9 shows that specialising $(C,\Gamma)$ to $(C_{-},\Gamma_{-})$ or $(C_{+},\Gamma_{+})$, the curve constructed in Theorem 5.1 deforms according to (3.24). Therefore, its sweeping surface forms a semi-discrete analogue of the discrete K-surface (see figures in §6).
Since the curvature and torsion angles uniquely determine a curve up to a rigid transformation, Proposition 5.3 shows that the motion constructed in Theorem 5.1 is periodic up to a global rigid transform. Here, we control the global orientation by choosing the parameters $C$ and $\Gamma$ appropriately so that the motion becomes periodic in terms of vertex positions.
Proposition 5.11.
Let $l\in\mathbb{Z}$. If we set
(5.22) $C$ $=\frac{4\pi i}{y}\frac{1}{\Delta_{3}-\Delta_{1}},$
(5.23) $\Gamma$ $=-\frac{\pi}{y}\left(4r+2l+1\right),$
we have
(5.24) ${\gamma_{n}}(t+y)={\gamma_{n}}(t),$
for any $n\in\mathbb{Z}$ and $t\in\mathbb{R}$.
Proof.
First, we prove the periodicity for the $Z$-component. Denote the $Z$-component of $\gamma_{n}$ by ${\gamma_{n}}_{Z}$. From the pseudo-period of elliptic theta functions (A.12), we see
(5.25) $\begin{split}&{\gamma_{n}}_{Z}(t+y)-{\gamma_{n}}_{Z}(t)\\ &=-2\frac{\partial}{\partial z}\log\left\{\exp\left(\frac{C}{2}yz\right)\frac{\vartheta_{2}\left(iv\left(n-\frac{1}{2}\right)+\frac{z}{\Delta_{3}-\Delta_{1}}+it+iy\right)}{\vartheta_{2}\left(iv\left(n-\frac{1}{2}\right)+\frac{z}{\Delta_{3}-\Delta_{1}}+it\right)}\right\}\\ &=-2\frac{\partial}{\partial z}\log\exp\left\{\frac{C}{2}yz-\pi i\left(2iv\left(n-\frac{1}{2}\right)+\frac{2z}{\Delta_{3}-\Delta_{1}}+2it\right)+\pi y\right\}\\ &=-2\left(\frac{C}{2}y-\frac{2\pi i}{\Delta_{3}-\Delta_{1}}\right)=0.\end{split}$
For the $X$ and $Y$ components, from (3.10) we have to show $\frac{H_{n}(t+y)}{F_{n}(t+y)}\frac{F_{n}(t)}{H_{n}(t)}=1$.
If we write $P_{n}=iv\left(n-\frac{1}{2}\right)+\frac{z}{\Delta_{3}-\Delta_{1}}+it$, the left-hand side can be computed using (A.12) as
(5.26) $\begin{split}&\exp\left(i\Gamma y\right)\frac{\vartheta_{4}\left(P_{n}-2r+iy\right)\vartheta_{2}\left(P_{n}\right)}{\vartheta_{2}\left(P_{n}+iy\right)\vartheta_{4}\left(P_{n}-2r\right)}\\ &=-\exp\left(i\Gamma y\right)\frac{\exp\left\{-\pi i\left(2P_{n}-4r+iy\right)\right\}}{\exp\left\{-\pi i\left(2P_{n}+iy\right)\right\}}\\ &=-\exp\left\{i\left(\Gamma y+4\pi r\right)\right\}=1.\end{split}$
5.3. Evolution of curvature
The curve constructed in Theorem 5.1 reveals an interesting recurrence equation for $\kappa$.
Proposition 5.12.
The signed curvature angle of the curve obtained in Theorem 5.1 satisfies the following recurrence equation for any $t$ and $n$:
(5.27) $\tan\frac{\kappa_{n+1}}{2}\tan\frac{\kappa_{n-1}}{2}-\tan^{2}\frac{\kappa_{n}}{2}+\tan\frac{\kappa_{n+1}}{2}\tan^{2}\frac{\kappa_{n}}{2}\tan\frac{\kappa_{n-1}}{2}=-\frac{\vartheta_{1}\left(iv\right)^{4}}{\vartheta_{3}\left(iv\right)^{4}}.$
The right-hand side of this equation depends neither on $n$ nor on $t$. Therefore, the quantity on the left-hand side can be considered a conserved quantity of the signed curvature angle.
A proof of Proposition 5.12 is given in Appendix E. The evolution of the signed curvature angle satisfies the semi-discrete mKdV equation in Proposition 5.7.
Theorem 5.13.
The signed curvature angle (5.3) satisfies
(5.28) $\dot{\kappa}_{n}=2\beta^{2}\eta_{3}\left(\tan\frac{\kappa_{n+1}}{2}-\tan\frac{\kappa_{n-1}}{2}\right),$
where
(5.29) $\eta_{3}=i\frac{\vartheta_{1}\left(iv\right)\vartheta_{3}\left(iv\right)}{\vartheta_{1}\left(2iv\right)\vartheta_{3}\left(0\right)}\left(d_{1}(iv)-d_{3}(iv)\right).$
Combined with Proposition 5.7, (5.6), and the fact that $e^{i\lambda}\delta_{2}\in\mathbb{R}$, we see
(5.30) $2\beta^{2}\eta_{3}=\left(2\beta\beta_{1}\cos\lambda+2\beta\beta_{2}\sin\lambda\right).$
A proof of Theorem 5.13 is given in Appendix F.
5.4. Potential function of curvature
A family of functions $\Theta_{n}$ parameterised by $n\in\mathbb{Z}$ is said to be a potential of the curvature $\kappa$ if the equation $2\kappa_{n}=\Theta_{n+1}-\Theta_{n-1}$ holds. When $\gamma$ deforms by (5.19), the potential function $\Theta$ is known to satisfy the semi-discrete potential mKdV equation or the semi-discrete sine-Gordon equation [19]. Here, we derive the evolution equations satisfied by the potential function for the curve constructed by Theorem 5.1.
We construct a potential function which simultaneously satisfy (3.25) and (3.26) in terms of elliptic theta functions.
Theorem 5.14.
For fixed $v,y$, define the function $\Theta$ as follows.
(5.31) $\Theta_{n}\left(t\right)=\Theta_{n}\left(t;v,y\right)=4\arctan\frac{\vartheta_{1}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y};\frac{2i}{y}\right)}{\vartheta_{4}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y};\frac{2i}{y}\right)}.$
Then we see that $\kappa$ defined in (5.3) satisfies
(5.32) $\kappa_{n}=\frac{\Theta_{n+1}-\Theta_{n-1}}{2}.$
A proof of Theorem 5.14 is given in Appendix G. The potential function simultaneously solves the semi-discrete potential mKdV equation and the semi-discrete sine-Gordon equation.
Theorem 5.15.
Let $y>0$ and $v\in\mathbb{R}\setminus{(y\mathbb{Z})}$. The potential function $\Theta$ defined by (5.31) simultaneously satisfies the semi-discrete potential mKdV and sine-Gordon equations
(5.33) $\frac{d}{dt}\left(\frac{\Theta_{n+1}+\Theta_{n}}{2}\right)$ $=\frac{L_{1}(v,y)}{y}\sin\left(\frac{\Theta_{n+1}-\Theta_{n}}{2}\right),$
(5.34) $\frac{d}{dt}\left(\frac{\Theta_{n+1}-\Theta_{n}}{2}\right)$ $=\frac{L_{2}(v,y)}{y}\sin\left(\frac{\Theta_{n+1}+\Theta_{n}}{2}\right),$
where $Q_{1},Q_{2},L_{1},L_{2}\in\mathbb{R}$ are defined by
(5.35) $Q_{1}(v,y)=\frac{\vartheta^{{}^{\prime}}_{1}\left(0;\frac{2i}{y}\right)\vartheta_{1}\left(\frac{v}{y};\frac{2i}{y}\right)}{\vartheta_{4}\left(0;\frac{2i}{y}\right)\vartheta_{4}\left(\frac{v}{y};\frac{2i}{y}\right)},\quad Q_{2}(v,y)=\frac{\vartheta^{{}^{\prime}}_{1}\left(0;\frac{2i}{y}\right)\vartheta_{4}\left(\frac{v}{y};\frac{2i}{y}\right)}{\vartheta_{1}\left(\frac{v}{y};\frac{2i}{y}\right)\vartheta_{4}\left(0;\frac{2i}{y}\right)},$
(5.36) $L_{1}(v,y)=d_{1}\left(\frac{v}{y};\frac{2}{y}\right)-d_{4}\left(\frac{v}{y};\frac{2}{y}\right)+Q_{1}(v,y)+Q_{2}(v,y),$
(5.37) $L_{2}(v,y)=d_{1}\left(\frac{v}{y};\frac{2}{y}\right)-d_{4}\left(\frac{v}{y};\frac{2}{y}\right)-Q_{1}(v,y)-Q_{2}(v,y).$
A proof of Theorem 5.15 is given in Appendix H.
The following shows two equations are related by a parameter shift.
Proposition 5.16.
By the transformation of parameters $v\rightarrow v+y$ and $r\rightarrow\frac{1}{2}-r$, the signed curvature angle $\kappa$, the torsion angles $\lambda$, the potential functions $\Theta$, the coefficients of the semi-discrete mKdV equation $L_{1}/y$ and of the semi-discrete sine-Gordon equation $L_{2}/y$, the normal $\widetilde{N}$ and the binormal $\widetilde{B}$ are transformed as
(5.38) $\kappa_{n}\left(t\right)$ $\rightarrow(-1)^{n+1}\kappa_{n}\left(t-\frac{y}{2}\right),\quad\lambda\rightarrow\pi+\lambda,$
(5.39) $\Theta_{n}\left(t\right)$ $\rightarrow(-1)^{n}\Theta_{n}\left(t-\frac{y}{2}\right),$
(5.40) $\frac{L_{1}}{y}$ $\rightarrow\frac{L_{2}}{y},\quad\frac{L_{2}}{y}\rightarrow\frac{L_{1}}{y},$
(5.41) $\widetilde{N}_{n}(t)$ $\rightarrow(-1)^{n+1}\widetilde{N}_{n}\left(t-\frac{y}{2}\right),\quad\widetilde{B}_{n}(t)\rightarrow(-1)^{n+1}\widetilde{B}_{n}\left(t-\frac{y}{2}\right).$
A proof of Proposition 5.16 is given in Appendix I. The transformation $\Theta_{n}\rightarrow\widetilde{\Theta}_{n}:=(-1)^{n}\Theta_{n}$ exchanges a solution of the semi-discrete potential mKdV equation (5.33) with one of the semi-discrete sine-Gordon equation (5.34). Proposition 5.16 shows that this equation-level transformation is also achieved at the solution level by transforming the parameters.
Remark 5.17.
Rewriting (5.28) in terms of the potential function $\Theta$, we obtain
(5.42) $\frac{d}{dt}\Theta_{n}=4\beta^{2}\eta_{3}\tan\left(\frac{\Theta_{n+1}-\Theta_{n-1}}{4}\right),$
which is also called the semi-discrete potential mKdV equation [14, 13, 6]. In other words, the potential function $\Theta$ defined by (5.31) is a simultaneous solution of three differential-difference equations, (5.33), (5.34) and (5.42). It does not usually happen that these equations are simultaneously solved. For example, it can be checked that the 2 and 3-soliton solutions of the semi-discrete sine-Gordon equation constructed in [6, 7] are not solutions of the semi-discrete potential mKdV equation.
The following describes an explicit relation between our construction and the previous work [19] reviewed in §3.2.
Proposition 5.18.
When $C_{+}$ and $\Gamma_{+}$ are chosen as in (5.18), we have
(5.43) $\frac{L_{1}(v,y)}{y}\frac{1}{1+\cos\lambda}=i\frac{\vartheta_{1}^{{}^{\prime}}(0)}{\vartheta_{1}(iv)}\frac{u_{2}}{\vartheta_{3}(r)^{2}\vartheta_{1}(r)^{2}}=\rho_{+}.$
Similarly, when $C_{-}$ and $\Gamma_{-}$ are chosen as in (5.18), we have
(5.44) $\frac{L_{2}(v,y)}{y}\frac{1}{1-\cos\lambda}=-i\frac{\vartheta_{1}^{{}^{\prime}}(0)}{\vartheta_{1}(iv)}\frac{u_{2}}{\vartheta_{4}(r)^{2}\vartheta_{2}(r)^{2}}=\rho_{-}.$
A proof of Proposition 5.18 is given in Appendix J. Note the difference in approaches of [19] and this work; [19] starts with a geometric property observed in the motion of Kaleidocycles, and derive evolution equations, whereas we aim to construct a particular motion of Kaleidocycles and investigate the equations satisfied by them. They turn out to arrive at the same place; the latter gives an explicit solution to the equations derived in the former.
Table 2 summarises the relationships among the three deformation equations.
Table 2. Relation between the three deformation equations
Values of
$C,\Gamma$
Deformation equation of the curve Equation for potential function
$C,\Gamma\in\mathbb{R}$
$\dot{\gamma}_{n}=\footnotesize\left(\begin{array}[]{c}0\\ 0\\ 2\beta_{1}\left(\beta-2\delta_{1}\right)-\epsilon_{1}-C\end{array}\right)\normalsize$
$+2\beta\left(\beta_{1}T_{n}-\beta_{1}\tan\frac{\kappa_{n}}{2}\widetilde{N}_{n}+\beta_{2}\tan\frac{\kappa_{n}}{2}\widetilde{B}_{n}\right)$
$\dot{\Theta}_{n}=\left(4\beta\beta_{1}\cos\lambda+4\beta\beta_{2}\sin\lambda\right)$
×$\times\tan\left(\frac{\Theta_{n+1}-\Theta_{n-1}}{4}\right)$
$\downarrow$ Specialise the deformation equation by choosing rigid transformation parameters $C,\Gamma$.
$C_{+},\Gamma_{+}$ $\left\langle\dot{\gamma}_{n},\widetilde{B}_{n}\right\rangle=0,\quad\langle\dot{\gamma}_{n},\dot{\gamma}_{n}\rangle=\rho^{2}_{+}$
$\frac{\dot{\Theta}_{n+1}+\dot{\Theta}_{n}}{2}=\left(1+\cos\lambda\right)\rho_{+}$
×$\times\sin\left(\frac{\Theta_{n+1}-\Theta_{n}}{2}\right)$
$\updownarrow$ By replacing the parameters of the solution with $v\rightarrow v+y,r\rightarrow 1/2-r$,
these two deformation equations transform to each other.
$C_{-},\Gamma_{-}$ $\left\langle\dot{\gamma}_{n},\widetilde{B}_{n}\right\rangle=0,\quad\langle\dot{\gamma}_{n},\dot{\gamma}_{n}\rangle=\rho^{2}_{-}$
$\frac{\dot{\Theta}_{n+1}-\dot{\Theta}_{n}}{2}=\left(1-\cos\lambda\right)\rho_{-}$
×$\times\sin\left(\frac{\Theta_{n+1}+\Theta_{n}}{2}\right)$
6. Examples
In this section, we present examples of Kaleidocycles constructed by Theorem 4.2 and the semi-discrete K-surfaces obtained as their motion trajectory as explained in Remark 5.10.
Numerical experiments using our computer codes [17] suggest that the $8$-, $9$- and $15$-Kaleidocycles in Figs. 2, 3, and 4 have the minimum torsion angles among non-orientable $8$-, $9$-, and $15$-Kaleidocycles, respectively [18]. In other words, they are the Möbius Kaleidocycles [20].
For Fig. 5, the torsion angle is larger than the above minimum one. However, it has a self-linking number $7.5$. As the self-linking number of a Kaleidocycle is a half-integer, these $15$-Kaleidocycles depicted in Figs. 4 and 5 belong to different connected components of the configuration space of $15$-Kaleidocycles. We numerically find that the latter attains the local minima of the torsion angle. We present the following two conjectures based on our extensive numerical experiments:
Conjecture 6.1.
Any Kaleidocycle obtained by Theorem 4.2 has a minimal or maximal torsion angle. In other words, the projection from the space defined by (2.2) onto the $\lambda$-coordinate attains a critical value.
Conjecture 6.2.
The self-linking number of the closed curve $\gamma$ constructed in Theorem 4.2 is $m/2$ when the curve is unknotted.
As briefly discussed in §1, these Kaleidocycles with critical torsion angles numerically exhibit special properties, including the single degree of freedom [18, 20, 30]. Our construction produces this interesting family of Kaleidocycles and may shed light on the rigorous proof for some of their unique properties.
Figure 2. (Left) An example of 8-Kaleidocycle with $(v,r,y,m)\approx(0.400,0.717,1.067,3)$ and $\cos\lambda\approx 0.4700$. (Middle) 8-Kaleidocycle as a discrete curve. (Right) the semi-discrete K-surface obtained as its trajectory.
Figure 3. (Left) An example of 9-Kaleidocycle with $(v,r,y,m)\approx(0.342,0.289,1.027,3)$ and $\cos\lambda\approx 0.5852$. (Middle) 9-Kaleidocycle as a discrete curve. (Right) the semi-discrete K-surface obtained as its trajectory.
Figure 4. (Left) An example of 15-Kaleidocycle with $(v,r,y,m)\approx(0.189,0.300,0.947,3)$ and $\cos\lambda\approx 0.8533$. (Middle) 15-Kaleidocycle as a discrete curve. (Right) the semi-discrete K-surface obtained as its trajectory.
Figure 5. (Left) An example of 15-Kaleidocycle with $(v,r,y,m)\approx(0.401,0.689,1.204,5)$ and $\cos\lambda\approx 0.6497$. (Middle) 15-Kaleidocycle as a discrete curve. (Right) the semi-discrete K-surface obtained as its trajectory.
Appendix
In the appendix, we prove the theorems and propositions in the paper by using the properties of elliptic theta functions listed in Appendix A.
Appendix A Functional identities of elliptic theta functions
We collect some properties of elliptic theta functions which we rely heavily on our calculations.
A.1. Four term identities
Let
(A.1) $\begin{split}&X_{1}=\frac{1}{2}(X+Y+U+V),\quad Y_{1}=\frac{1}{2}(X+Y-U-V),\\ &U_{1}=\frac{1}{2}(X-Y+U-V),\quad V_{1}=\frac{1}{2}(X-Y-U+V),\end{split}$
then for any $X,Y,U,V\in\mathbb{C}$, elliptic theta functions satisfy the following identities [26, 22].
(A.2) $\vartheta_{2}(X_{1})\vartheta_{2}(Y_{1})\vartheta_{3}(U_{1})\vartheta_{3}(V_{1})-\vartheta_{4}(X_{1})\vartheta_{4}(Y_{1})\vartheta_{1}(U_{1})\vartheta_{1}(V_{1})$
$=\vartheta_{2}(X)\vartheta_{2}(Y)\vartheta_{3}(U)\vartheta_{3}(V)-\vartheta_{4}(X)\vartheta_{4}(Y)\vartheta_{1}(U)\vartheta_{1}(V),$
(A.3) $\vartheta_{4}(X_{1})\vartheta_{3}(Y_{1})\vartheta_{1}(U_{1})\vartheta_{2}(V_{1})-\vartheta_{2}(X_{1})\vartheta_{1}(Y_{1})\vartheta_{3}(U_{1})\vartheta_{4}(V_{1})$
$=\vartheta_{4}(X)\vartheta_{3}(Y)\vartheta_{1}(U)\vartheta_{2}(V)-\vartheta_{2}(X)\vartheta_{1}(Y)\vartheta_{3}(U)\vartheta_{4}(V),$
(A.4) $\vartheta_{2}(X_{1})\vartheta_{2}(Y_{1})\vartheta_{4}(U_{1})\vartheta_{4}(V_{1})+\vartheta_{1}(X_{1})\vartheta_{1}(Y_{1})\vartheta_{3}(U_{1})\vartheta_{3}(V_{1})$
$=\vartheta_{4}(X)\vartheta_{4}(Y)\vartheta_{2}(U)\vartheta_{2}(V)-\vartheta_{3}(X)\vartheta_{3}(Y)\vartheta_{1}(U)\vartheta_{1}(V),$
(A.5) $\vartheta_{3}(X_{1})\vartheta_{3}(Y_{1})\vartheta_{2}(U_{1})\vartheta_{2}(V_{1})+\vartheta_{1}(X_{1})\vartheta_{1}(Y_{1})\vartheta_{4}(U_{1})\vartheta_{4}(V_{1})$
$=\vartheta_{3}(X)\vartheta_{3}(Y)\vartheta_{2}(U)\vartheta_{2}(V)+\vartheta_{1}(X)\vartheta_{1}(Y)\vartheta_{4}(U)\vartheta_{4}(V),$
(A.6) $\vartheta_{4}(X_{1})\vartheta_{4}(Y_{1})\vartheta_{4}(U_{1})\vartheta_{4}(V_{1})-\vartheta_{1}(X_{1})\vartheta_{1}(Y_{1})\vartheta_{1}(U_{1})\vartheta_{1}(V_{1})$
$=\vartheta_{4}(X)\vartheta_{4}(Y)\vartheta_{4}(U)\vartheta_{4}(V)-\vartheta_{1}(X)\vartheta_{1}(Y)\vartheta_{1}(U)\vartheta_{1}(V),$
(A.7) $\vartheta_{4}(X_{1})\vartheta_{4}(Y_{1})\vartheta_{1}(U_{1})\vartheta_{1}(V_{1})-\vartheta_{1}(X_{1})\vartheta_{1}(Y_{1})\vartheta_{4}(U_{1})\vartheta_{4}(V_{1})$
$=\vartheta_{4}(X)\vartheta_{4}(Y)\vartheta_{1}(U)\vartheta_{1}(V)-\vartheta_{1}(X)\vartheta_{1}(Y)\vartheta_{4}(U)\vartheta_{4}(V),$
(A.8) $\vartheta_{4}(X_{1})\vartheta_{4}(Y_{1})\vartheta_{1}(U_{1})\vartheta_{1}(V_{1})+\vartheta_{2}(X_{1})\vartheta_{2}(Y_{1})\vartheta_{3}(U_{1})\vartheta_{3}(V_{1})$
$=\vartheta_{3}(X)\vartheta_{3}(Y)\vartheta_{2}(U)\vartheta_{2}(V)-\vartheta_{1}(X)\vartheta_{1}(Y)\vartheta_{4}(U)\vartheta_{4}(V),$
(A.9) $\vartheta_{3}(X_{1})\vartheta_{3}(Y_{1})\vartheta_{1}(U_{1})\vartheta_{1}(V_{1})-\vartheta_{1}(X_{1})\vartheta_{1}(Y_{1})\vartheta_{3}(U_{1})\vartheta_{3}(V_{1})$
$=\vartheta_{3}(X)\vartheta_{3}(Y)\vartheta_{1}(U)\vartheta_{1}(V)-\vartheta_{1}(X)\vartheta_{1}(Y)\vartheta_{3}(U)\vartheta_{3}(V),$
(A.10) $\vartheta_{3}(X_{1})\vartheta_{3}(Y_{1})\vartheta_{4}(U_{1})\vartheta_{4}(V_{1})-\vartheta_{1}(X_{1})\vartheta_{1}(Y_{1})\vartheta_{2}(U_{1})\vartheta_{2}(V_{1})$
$=\vartheta_{3}(X)\vartheta_{3}(Y)\vartheta_{4}(U)\vartheta_{4}(V)-\vartheta_{1}(X)\vartheta_{1}(Y)\vartheta_{2}(U)\vartheta_{2}(V).$
A.2. Shifts by periods and half-periods
For any $X\in\mathbb{C}$, $n\in\mathbb{Z}$ and $y>0$, the following holds [26, 22].
(A.11) $\vartheta_{1}\left(X+n\right)=(-1)^{n}\vartheta_{1}\left(X\right),\quad\vartheta_{4}\left(X+n\right)=\vartheta_{4}\left(X\right).$
(A.12) $\begin{split}&\vartheta_{1}\left(X+niy;iy\right)=(-1)^{n}\exp\{-\pi i\left(2nX+n^{2}iy\right)\}\vartheta_{1}\left(X;iy\right),\\ &\vartheta_{2}\left(X+niy;iy\right)=\exp\{-\pi i\left(2nX+n^{2}iy\right)\}\vartheta_{2}\left(X;iy\right),\\ &\vartheta_{3}\left(X+niy;iy\right)=\exp\{-\pi i\left(2nX+n^{2}iy\right)\}\vartheta_{3}\left(X;iy\right),\\ &\vartheta_{4}\left(X+niy;iy\right)=(-1)^{n}\exp\{-\pi i\left(2nX+n^{2}iy\right)\}\vartheta_{4}\left(X;iy\right).\end{split}$
(A.13) $\begin{split}&\vartheta_{1}\left(X+\frac{1}{2}\right)=\vartheta_{2}\left(X\right),\quad\vartheta_{2}\left(X+\frac{1}{2}\right)=-\vartheta_{1}\left(X\right),\\ &\vartheta_{3}\left(X+\frac{1}{2}\right)=\vartheta_{4}\left(X\right),\quad\vartheta_{4}\left(X+\frac{1}{2}\right)=\vartheta_{3}\left(X\right).\end{split}$
(A.14) $\begin{split}&\vartheta_{1}\left(X+\frac{iy}{2};iy\right)=i\exp\left\{-\pi i\left(X+\frac{iy}{4}\right)\right\}\vartheta_{4}\left(X;iy\right),\\ &\vartheta_{2}\left(X+\frac{iy}{2};iy\right)=\exp\left\{-\pi i\left(X+\frac{iy}{4}\right)\right\}\vartheta_{3}\left(X;iy\right),\\ &\vartheta_{3}\left(X+\frac{iy}{2};iy\right)=\exp\left\{-\pi i\left(X+\frac{iy}{4}\right)\right\}\vartheta_{2}\left(X;iy\right),\\ &\vartheta_{4}\left(X+\frac{iy}{2};iy\right)=i\exp\left\{-\pi i\left(X+\frac{iy}{4}\right)\right\}\vartheta_{1}\left(X;iy\right).\end{split}$
A.3. Identity for theta constants
For any $y>0$, elliptic theta functions satisfy the following identity [26, 22].
(A.15) $\vartheta_{1}^{{}^{\prime}}\left(0;iy\right)=\pi\vartheta_{2}\left(0;iy\right)\vartheta_{3}\left(0;iy\right)\vartheta_{4}\left(0;iy\right).$
A.4. Modular transformations
For any $X\in\mathbb{C}$ and $y>0$, elliptic theta functions satisfy the following identities [26, 22].
(A.16) $\begin{split}&\vartheta_{1}\left(\frac{X}{y};i\frac{2}{y}\right)=-i\sqrt{\frac{y}{2}}\exp\left(-\pi\frac{X^{2}}{2y}\right)\vartheta_{1}\left(\frac{iX}{2};i\frac{y}{2}\right),\\ &\vartheta_{2}\left(\frac{X}{y};i\frac{2}{y}\right)=\sqrt{\frac{y}{2}}\exp\left(-\pi\frac{X^{2}}{2y}\right)\vartheta_{4}\left(\frac{iX}{2};i\frac{y}{2}\right),\\ &\vartheta_{3}\left(\frac{X}{y};i\frac{2}{y}\right)=\sqrt{\frac{y}{2}}\exp\left(-\pi\frac{X^{2}}{2y}\right)\vartheta_{3}\left(\frac{iX}{2};i\frac{y}{2}\right),\\ &\vartheta_{4}\left(\frac{X}{y};i\frac{2}{y}\right)=\sqrt{\frac{y}{2}}\exp\left(-\pi\frac{X^{2}}{2y}\right)\vartheta_{2}\left(\frac{iX}{2};i\frac{y}{2}\right).\end{split}$
(A.17) $\frac{\vartheta_{1}\left(X;iy\right)}{\vartheta_{2}\left(X;iy\right)}=-i\frac{\vartheta_{1}\left(\frac{iX}{y};\frac{i}{y}\right)}{\vartheta_{4}\left(\frac{iX}{y};\frac{i}{y}\right)}.$
A.5. Bilinear identities connecting theta functions with $y$ and $y/2$
For any $X,Y\in\mathbb{C}$, and $y>0$, elliptic theta functions satisfy the following identities [26, 22].
$2\vartheta_{1}\left(X+Y;iy\right)\vartheta_{4}\left(X-Y;iy\right)$
(A.18) $=\vartheta_{1}\left(X;\frac{iy}{2}\right)\vartheta_{2}\left(Y;\frac{iy}{2}\right)+\vartheta_{2}\left(X;\frac{iy}{2}\right)\vartheta_{1}\left(Y;\frac{iy}{2}\right),$
$2\vartheta_{2}\left(X+Y;iy\right)\vartheta_{3}\left(X-Y;iy\right)$
(A.19) $=\vartheta_{2}\left(X;\frac{iy}{2}\right)\vartheta_{2}\left(Y;\frac{iy}{2}\right)-\vartheta_{1}\left(X;\frac{iy}{2}\right)\vartheta_{1}\left(Y;\frac{iy}{2}\right).$
(A.20) $\begin{split}&\vartheta_{3}\left(\frac{X}{2};\frac{iy}{2}\right)\vartheta_{4}\left(\frac{X}{2};\frac{iy}{2}\right)=\vartheta_{4}\left(X;iy\right)\vartheta_{4}\left(0;iy\right),\\ &\vartheta_{1}\left(\frac{X}{2};\frac{iy}{2}\right)\vartheta_{2}\left(\frac{X}{2};\frac{iy}{2}\right)=\vartheta_{1}\left(X;iy\right)\vartheta_{4}\left(0;iy\right),\\ -&\vartheta_{1}\left(\frac{X}{2};\frac{iy}{2}\right)^{2}+\vartheta_{2}\left(\frac{X}{2};\frac{iy}{2}\right)^{2}=2\vartheta_{2}\left(X;iy\right)\vartheta_{3}\left(0;iy\right).\\ \end{split}$
Appendix B Proofs of Theorem 4.1 and Theorem 5.1
Here we prove that the $\tau$ functions given in Theorem 5.1 satisfy (3.1)-(3.6). Theorem 4.1 follows from Theorem 5.1 by setting $t=0$. We assume $z=0$ without loss of generality except for the proof of (3.5). Putting
(B.5) $\left\{\begin{array}[]{l}\vspace{2mm}X=ivn+r+it,\\ \vspace{2mm}Y=ivn-r+it,\\ \vspace{2mm}U=\frac{1}{2}iv+r,\\ V=\frac{1}{2}iv-r,\\ \end{array}\right.$
in (A.2) yields (3.1). Putting
(B.10) $\left\{\begin{array}[]{l}\vspace{2mm}X=iv\left(n-\frac{1}{2}\right)+it,\\ \vspace{2mm}Y=ivn+r+it,\\ \vspace{2mm}U=iv,\\ V=\frac{1}{2}iv-r,\\ \end{array}\right.$
in (A.2) yields (3.2). Putting
(B.15) $\left\{\begin{array}[]{l}\vspace{2mm}X=iv\left(n-\frac{1}{2}\right)+it,\\ \vspace{2mm}Y=\frac{1}{2}iv+r,\\ \vspace{2mm}U=iv,\\ V=ivn-r+it,\\ \end{array}\right.$
in (A.3) yields (3.3). Putting
(B.20) $\left\{\begin{array}[]{l}\vspace{2mm}X=ivn-r+it,\\ \vspace{2mm}Y=-\frac{1}{2}iv+r,\\ \vspace{2mm}U=ivn-r+it,\\ V=-\frac{1}{2}iv+r,\\ \end{array}\right.$
in (A.4) yields
(B.21) $\begin{split}&\frac{\vartheta_{2}\left(0\right)\vartheta_{4}\left(0\right)}{\vartheta_{1}\left(-\frac{1}{2}iv+r\right)\vartheta_{3}\left(-\frac{1}{2}iv+r\right)}\vartheta_{2}\left(iv\left(n-\frac{1}{2}\right)+it\right)\vartheta_{4}\left(iv\left(n+\frac{1}{2}\right)-2r+it\right)\\ &=\frac{\vartheta_{2}\left(-\frac{1}{2}iv+r\right)\vartheta_{4}\left(-\frac{1}{2}iv+r\right)}{\vartheta_{1}\left(-\frac{1}{2}iv+r\right)\vartheta_{3}\left(-\frac{1}{2}iv+r\right)}\vartheta_{2}\left(ivn-r+it\right)\vartheta_{4}\left(ivn-r+it\right)\\ &-\vartheta_{3}\left(ivn-r+it\right)\vartheta_{1}\left(ivn-r+it\right).\end{split}$
Also, putting
(B.26) $\left\{\begin{array}[]{l}\vspace{2mm}X=ivn-r+it,\\ \vspace{2mm}Y=\frac{1}{2}iv+r,\\ \vspace{2mm}U=ivn-r+it,\\ V=\frac{1}{2}iv+r,\\ \end{array}\right.$
in (A.4) yields
(B.27) $\begin{split}&\frac{\vartheta_{2}\left(0\right)\vartheta_{4}\left(0\right)}{\vartheta_{1}\left(\frac{1}{2}iv+r\right)\vartheta_{3}\left(\frac{1}{2}iv+r\right)}\vartheta_{2}\left(iv\left(n+\frac{1}{2}\right)+it\right)\vartheta_{4}\left(iv\left(n-\frac{1}{2}\right)-2r+it\right)\\ &=\frac{\vartheta_{2}\left(\frac{1}{2}iv+r\right)\vartheta_{4}\left(\frac{1}{2}iv+r\right)}{\vartheta_{1}\left(\frac{1}{2}iv+r\right)\vartheta_{3}\left(\frac{1}{2}iv+r\right)}\vartheta_{2}\left(ivn-r+it\right)\vartheta_{4}\left(ivn-r+it\right)\\ &-\vartheta_{3}\left(ivn-r+it\right)\vartheta_{1}\left(ivn-r+it\right).\end{split}$
Subtracting (B.27) from (B.21) and substituting
(B.28) $\begin{split}&\vartheta_{3}\left(2r\right)\vartheta_{4}\left(0\right)\vartheta_{2}\left(0\right)\vartheta_{1}\left(iv\right)\\ &=\vartheta_{1}\left(\frac{1}{2}iv+r\right)\vartheta_{3}\left(\frac{1}{2}iv+r\right)\vartheta_{2}\left(-\frac{1}{2}iv+r\right)\vartheta_{4}\left(-\frac{1}{2}iv+r\right)\\ &-\vartheta_{1}\left(-\frac{1}{2}iv+r\right)\vartheta_{3}\left(-\frac{1}{2}iv+r\right)\vartheta_{2}\left(\frac{1}{2}iv+r\right)\vartheta_{4}\left(\frac{1}{2}iv+r\right),\end{split}$
in the equation, we obtain
(B.29) $\begin{split}&\frac{\vartheta_{1}\left(-\frac{1}{2}iv+r\right)\vartheta_{3}\left(-\frac{1}{2}iv+r\right)\vartheta_{1}\left(\frac{1}{2}iv+r\right)\vartheta_{3}\left(\frac{1}{2}iv+r\right)}{\vartheta_{3}(2r)\vartheta_{1}(iv)}\\ &\times\left\{\frac{\vartheta_{1}\left(\frac{1}{2}iv+r\right)\vartheta_{3}\left(\frac{1}{2}iv+r\right)}{\vartheta_{1}\left(-\frac{1}{2}iv+r\right)\vartheta_{3}\left(-\frac{1}{2}iv+r\right)}\vartheta_{2}\left(iv\left(n-\frac{1}{2}\right)+it\right)\vartheta_{4}\left(iv\left(n+\frac{1}{2}\right)-2r+it\right)\right.\\ &\left.-\vartheta_{2}\left(iv\left(n+\frac{1}{2}\right)+it\right)\vartheta_{4}\left(iv\left(n-\frac{1}{2}\right)-2r+it\right)\right\}\\ &=\vartheta_{3}\left(\frac{1}{2}iv+r\right)\vartheta_{2}\left(ivn-r+it\right)\vartheta_{1}\left(\frac{1}{2}iv+r\right)\vartheta_{4}\left(ivn-r+it\right),\end{split}$
from which we see (3.4) holds. Putting
(B.34) $\left\{\begin{array}[]{l}\vspace{2mm}X=ivn+z+r+it,\\ \vspace{2mm}Y=ivn+z-r+it,\\ \vspace{2mm}U=\frac{1}{2}iv+r+h,\\ V=\frac{1}{2}iv-r+h,\\ \end{array}\right.$
in (A.2) yields
(B.35) $\begin{split}&\vartheta_{2}\left(iv\left(n+\frac{1}{2}\right)+z+it+h\right)\vartheta_{2}\left(iv\left(n-\frac{1}{2}\right)+z+it-h\right)\vartheta_{3}(2r)\vartheta_{3}(0)\\ &=\vartheta_{2}\left(ivn+z+r+it\right)\vartheta_{2}\left(ivn+z-r+it\right)\vartheta_{3}\left(\frac{1}{2}iv+r+h\right)\vartheta_{3}\left(\frac{1}{2}iv-r+h\right)\\ &-\vartheta_{4}\left(ivn+z+r+it\right)\vartheta_{4}\left(ivn+z-r+it\right)\vartheta_{1}\left(\frac{1}{2}iv+r+h\right)\vartheta_{1}\left(\frac{1}{2}iv-r+h\right).\end{split}$
Differentiating (B.35) with $h$ and then putting $h=0$, we obtain
(B.36) $\begin{split}&\vartheta_{3}(2r)\vartheta_{3}(0)D_{z}\vartheta_{2}\left(iv\left(n+\frac{1}{2}\right)+z+it\right)\cdot\vartheta_{2}\left(iv\left(n-\frac{1}{2}\right)+z+it\right)\\ &=\left(\frac{\vartheta^{{}^{\prime}}_{3}\left(\frac{1}{2}iv+r\right)}{\vartheta_{3}\left(\frac{1}{2}iv+r\right)}-\frac{\vartheta^{{}^{\prime}}_{3}\left(-\frac{1}{2}iv+r\right)}{\vartheta_{3}\left(-\frac{1}{2}iv+r\right)}\right)\\ &\times\vartheta_{3}\left(-\frac{1}{2}iv+r\right)\vartheta_{3}\left(\frac{1}{2}iv+r\right)\vartheta_{2}\left(ivn+z+r+it\right)\vartheta_{2}\left(ivn+z-r+it\right)\\ &+\left(\frac{\vartheta^{{}^{\prime}}_{1}\left(\frac{1}{2}iv+r\right)}{\vartheta_{1}\left(\frac{1}{2}iv+r\right)}-\frac{\vartheta^{{}^{\prime}}_{1}\left(-\frac{1}{2}iv+r\right)}{\vartheta_{1}\left(-\frac{1}{2}iv+r\right)}\right)\\ &\times\vartheta_{1}\left(-\frac{1}{2}iv+r\right)\vartheta_{1}\left(\frac{1}{2}iv+r\right)\vartheta_{4}\left(ivn+z+r+it\right)\vartheta_{4}\left(ivn+z-r+it\right).\end{split}$
Using the equation (B.35) with $h=0$ and substituting (B.36), we obtain
(B.37) $\begin{split}&\vartheta_{3}(2r)\vartheta_{3}(0)\left\{D_{z}\vartheta_{2}\left(iv\left(n+\frac{1}{2}\right)+z+it\right)\cdot\vartheta_{2}\left(iv\left(n-\frac{1}{2}\right)+z+it\right)\right.\\ &\left.+\left(\frac{\vartheta^{{}^{\prime}}_{3}\left(-\frac{1}{2}iv+r\right)}{\vartheta_{3}\left(-\frac{1}{2}iv+r\right)}-\frac{\vartheta^{{}^{\prime}}_{3}\left(\frac{1}{2}iv+r\right)}{\vartheta_{3}\left(\frac{1}{2}iv+r\right)}\right)\right.\\ &\left.\times\vartheta_{2}\left(iv\left(n+\frac{1}{2}\right)+z+it\right)\vartheta_{2}\left(iv\left(n-\frac{1}{2}\right)+z+it\right)\right\}\\ &=\left(\frac{\vartheta^{{}^{\prime}}_{1}\left(\frac{1}{2}iv+r\right)}{\vartheta_{1}\left(\frac{1}{2}iv+r\right)}-\frac{\vartheta^{{}^{\prime}}_{1}\left(-\frac{1}{2}iv+r\right)}{\vartheta_{1}\left(-\frac{1}{2}iv+r\right)}+\frac{\vartheta^{{}^{\prime}}_{3}\left(-\frac{1}{2}iv+r\right)}{\vartheta_{3}\left(-\frac{1}{2}iv+r\right)}-\frac{\vartheta^{{}^{\prime}}_{3}\left(\frac{1}{2}iv+r\right)}{\vartheta_{3}\left(\frac{1}{2}iv+r\right)}\right)\\ &\times\vartheta_{1}\left(-\frac{1}{2}iv+r\right)\vartheta_{4}\left(ivn+z+r+it\right)\vartheta_{1}\left(\frac{1}{2}iv+r\right)\vartheta_{4}\left(ivn+z-r+it\right),\end{split}$
from which we see (3.5) holds. By the definition of $G$ and $\widetilde{G}$, (3.6) follows.
Appendix C Proof of Proposition 5.5
For the proof of Proposition 5.5, we use the following lemma:
Lemma C.1 ([11, §3]).
The following equation holds:
(C.1) $f_{n}f_{n-1}^{*}+g_{n}g_{n-1}^{*}=f_{n}^{*}f_{n-1}+g_{n}^{*}g_{n-1}=\beta F_{n}F_{n}.$
C.1. Proof of Proposition 5.5
Here we prove that the $\tau$ functions given in Theorem 5.1 satisfy (5.8)-(5.14). We assume $z=0$ without loss of generality except for the proof of (5.14). Putting
(C.6) $\left\{\begin{array}[]{l}\vspace{2mm}X=-\frac{1}{2}iv+r,\\ \vspace{2mm}Y=\frac{1}{2}iv-r,\\ \vspace{2mm}U=ivn+r+it,\\ V=iv\left(n-1\right)-r+it,\\ \end{array}\right.$
in (A.8) yields (5.8). Also, (5.9) follows from (C.1) and (5.8). Putting
(C.11) $\left\{\begin{array}[]{l}\vspace{2mm}X=ivn-r+it,\\ \vspace{2mm}Y=-\frac{1}{2}iv-r+h,\\ \vspace{2mm}U=\frac{1}{2}iv-r+h,\\ V=iv\left(n-1\right)-r+it,\\ \end{array}\right.$
in (A.3) yields
(C.12) $\begin{split}&\vartheta_{4}\left(iv\left(n-\frac{1}{2}\right)-2r+it+h\right)\vartheta_{3}\left(0\right)\vartheta_{1}\left(iv\right)\vartheta_{2}\left(iv\left(n-\frac{1}{2}\right)+it-h\right)\\ &=\vartheta_{4}\left(ivn-r+it\right)\vartheta_{3}\left(-\frac{1}{2}iv-r+h\right)\\ &\times\vartheta_{1}\left(\frac{1}{2}iv-r+h\right)\vartheta_{2}\left(iv\left(n-1\right)-r+it\right)\\ &-\vartheta_{2}\left(ivn-r+it\right)\vartheta_{1}\left(-\frac{1}{2}iv-r+h\right)\\ &\times\vartheta_{3}\left(\frac{1}{2}iv-r+h\right)\vartheta_{4}\left(iv\left(n-1\right)-r+it\right),\end{split}$
After differentiating (C.12) by $h$ and putting $h=0$, we obtain (5.13) by rearrangement. Also, rearranging the equation with $h=0$ in (C.12) yields (5.10). Putting
(C.17) $\left\{\begin{array}[]{l}\vspace{2mm}X=\frac{1}{2}iv+r,\\ \vspace{2mm}Y=-\frac{1}{2}iv-r,\\ \vspace{2mm}U=iv\left(n-\frac{1}{2}\right)+it,\\ V=-iv\left(n-\frac{1}{2}\right)-it,\\ \end{array}\right.$
in (A.5) yields (5.11). Also, (5.12) follows from (5.8) and (5.11). Putting
(C.22) $\left\{\begin{array}[]{l}\vspace{2mm}X=iv\left(n-\frac{1}{2}\right)+it+\frac{z}{\Delta_{3}-\Delta_{1}},\\ \vspace{2mm}Y=iv\left(n-\frac{1}{2}\right)+it+\frac{z}{\Delta_{3}-\Delta_{1}},\\ \vspace{2mm}U=h,\\ V=h,\\ \end{array}\right.$
in (A.2) yields
(C.23) $\begin{split}&\vartheta_{2}\left(iv\left(n-\frac{1}{2}\right)+it+\frac{z}{\Delta_{3}-\Delta_{1}}+h\right)\vartheta_{2}\left(iv\left(n-\frac{1}{2}\right)+it+\frac{z}{\Delta_{3}-\Delta_{1}}-h\right)\\ &=-\vartheta_{4}\left(iv\left(n-\frac{1}{2}\right)+it+\frac{z}{\Delta_{3}-\Delta_{1}}\right)^{2}\frac{\vartheta_{1}\left(h\right)^{2}}{\vartheta_{3}\left(0\right)^{2}}\\ &+\vartheta_{2}\left(iv\left(n-\frac{1}{2}\right)+it+\frac{z}{\Delta_{3}-\Delta_{1}}\right)^{2}\frac{\vartheta_{3}\left(h\right)^{2}}{\vartheta_{3}\left(0\right)^{2}}.\end{split}$
After differentiating (C.23) twice with $h$ and putting $h=0$, we obtain (5.14).
Appendix D Proofs of Theorem 5.6 and Theorem 5.9
For the proofs of Theorem 5.6 and Theorem 5.9, we need the following lemmas.
Lemma D.1 ([11, §3]).
Using (3.2) and (3.3), we see that the following equation holds.
(D.1) $g_{n}f_{n-1}-f_{n}g_{n-1}=G_{n}F_{n}.$
Lemma D.2.
The following relation holds.
(D.2) $\omega=\delta_{2}-\delta_{2}^{*}.$
Proof.
Putting
(D.7) $\left\{\begin{array}[]{l}\vspace{2mm}X=-\frac{1}{2}iv+r,\\ \vspace{2mm}Y=-\frac{1}{2}iv+r,\\ \vspace{2mm}U=\frac{1}{2}iv+r,\\ V=\frac{1}{2}iv+r,\\ \end{array}\right.$
in (A.9) yields (D.2). ∎
Lemma D.3.
The following identity holds.
(D.8) $02\beta_{1}\left(\delta_{2}+\delta_{2}^{*}\right)+2i\omega\beta_{2}+\epsilon_{2}=2\delta_{2}\delta_{3}^{*}+2\delta_{2}^{*}\delta_{3}+\epsilon_{2}=$ 0 .
Proof.
Putting
(D.13) $\left\{\begin{array}[]{l}\vspace{2mm}X=\frac{1}{2}iv+r+h,\\ \vspace{2mm}Y=-\frac{1}{2}iv+r-h,\\ \vspace{2mm}U=\frac{1}{2}iv+r-h,\\ V=-\frac{1}{2}iv+r+h,\\ \end{array}\right.$
in (A.9) yields
(D.14) $\begin{split}&\vartheta_{3}\left(\frac{1}{2}iv+r+h\right)\vartheta_{3}\left(-\frac{1}{2}iv+r-h\right)\vartheta_{1}\left(\frac{1}{2}iv+r-h\right)\vartheta_{1}\left(-\frac{1}{2}iv+r+h\right)\\ &-\vartheta_{1}\left(\frac{1}{2}iv+r+h\right)\vartheta_{1}\left(-\frac{1}{2}iv+r-h\right)\vartheta_{3}\left(\frac{1}{2}iv+r-h\right)\vartheta_{3}\left(-\frac{1}{2}iv+r+h\right)\\ &=\vartheta_{3}\left(2r\right)\vartheta_{3}\left(0\right)\vartheta_{1}\left(iv\right)\vartheta_{1}\left(2h\right).\end{split}$
After differentiating (D.14) by $h$ and setting $h=0$, we obtain (D.8). ∎
For simplicity, we define $p$ and $q$ as follows
(D.15) $p_{n}=g_{n}^{*}f_{n-1}-f_{n}g_{n-1}^{*},$
(D.16) $q_{n}=g_{n}^{*}f_{n-1}+f_{n}g_{n-1}^{*}.$
Then we rewrite $D_{t}H_{n}\cdot F_{n}$ by using (5.6), (D.15) and (D.16) as follows
(D.17) $D_{t}H_{n}\cdot F_{n}=\beta_{1}q_{n}^{*}+i\beta_{2}p_{n}^{*}.$
Also, we see the following hold.
(D.18) $\tan\frac{\kappa_{n}}{2}\widetilde{N}_{n}=T_{n}-\frac{1}{2\beta F_{n}^{2}}\left(\begin{array}[]{c}\vspace{2mm}q_{n}^{*}+q_{n}\\ \vspace{2mm}\frac{1}{i}\left(q_{n}^{*}-q_{n}\right)\\ -2\left(\beta-2\delta_{1}\right)F_{n}^{2}+2\left(\delta_{2}+\delta_{2}^{*}\right)|G_{n}|^{2}\end{array}\right),$
(D.19) $\tan\frac{\kappa_{n}}{2}\widetilde{B}_{n}=\frac{1}{2\beta F_{n}^{2}}\left(\begin{array}[]{c}\vspace{2mm}\frac{1}{i}\left(p_{n}-p_{n}^{*}\right)\\ \vspace{2mm}p_{n}+p_{n}^{*}\\ -\frac{2}{i}\omega|G_{n}|^{2}\end{array}\right).$
We have the following lemma.
Lemma D.4.
The following relations hold.
(D.20) $p_{n}p_{n}^{*}=F_{n}^{2}|G_{n}|^{2}+\omega^{2}|G_{n}|^{4},$
(D.21) $\begin{split}p_{n}q_{n}^{*}-p_{n}^{*}q_{n}=2\left(2\delta_{1}-\beta\right)\omega F_{n}^{2}|G_{n}|^{2}+2\omega\left(\delta_{2}+\delta_{2}^{*}\right)|G_{n}|^{4},\end{split}$
(D.22) $\begin{split}&q_{n}q_{n}^{*}=4\delta_{1}\left(\beta-\delta_{1}\right)F_{n}^{4}+\left\{1+2\left(\beta-2\delta_{1}\right)\left(\delta_{2}+\delta_{2}^{*}\right)\right\}F_{n}^{2}|G_{n}|^{2}\\ &-\left(\delta_{2}+\delta_{2}^{*}\right)^{2}|G_{n}|^{4}.\end{split}$
Proof.
By definition of $p$, we see that
(D.23) $\begin{split}p_{n}p_{n}^{*}&=\left(g_{n}^{*}f_{n-1}-f_{n}g_{n-1}^{*}\right)\left(g_{n}f_{n-1}^{*}-f_{n}^{*}g_{n-1}\right)\\ &=\left(f_{n}^{*}g_{n-1}^{*}-g_{n}^{*}f_{n-1}^{*}\right)\left(f_{n}g_{n-1}-g_{n}f_{n-1}\right)\\ &+\left(f_{n}f_{n-1}^{*}-f_{n}^{*}f_{n-1}\right)\left(g_{n}^{*}g_{n-1}-g_{n}g_{n-1}^{*}\right)\end{split}$
holds. Substituting (D.1) and (5.9) into (D.23) yields (D.20). By definition of $p$ and $q$, we see that
(D.24) $\begin{split}&p_{n}q_{n}^{*}-p_{n}^{*}q_{n}\\ &=\left(g_{n}^{*}f_{n-1}-f_{n}g_{n-1}^{*}\right)\left(g_{n}f_{n-1}^{*}+f_{n}^{*}g_{n-1}\right)\\ &-\left(g_{n}f_{n-1}^{*}-f_{n}^{*}g_{n-1}\right)\left(g_{n}^{*}f_{n-1}+f_{n}g_{n-1}^{*}\right)\\ &=2\left(g_{n}^{*}f_{n-1}f_{n}^{*}g_{n-1}-g_{n}f_{n-1}^{*}f_{n}g_{n-1}^{*}\right)\\ \end{split}$
holds. On the other hand, from (C.1) and (5.9), the following holds.
(D.25) $\begin{split}&\beta\omega F_{n}^{2}|G_{n}|^{2}\\ &=\left(f_{n}f_{n-1}^{*}+g_{n}g_{n-1}^{*}\right)\left(g_{n}^{*}g_{n-1}-g_{n}g_{n-1}^{*}\right)\\ &=\left(f_{n}^{*}f_{n-1}+g_{n}^{*}g_{n-1}\right)\left(g_{n}^{*}g_{n-1}-g_{n}g_{n-1}^{*}\right)\\ &=\left(f_{n}f_{n-1}^{*}+g_{n}g_{n-1}^{*}\right)\left(f_{n}f_{n-1}^{*}-f_{n}^{*}f_{n-1}\right)\\ &=\left(f_{n}^{*}f_{n-1}+g_{n}^{*}g_{n-1}\right)\left(f_{n}f_{n-1}^{*}-f_{n}^{*}f_{n-1}\right).\\ \end{split}$
Using (D.25), we obtain
(D.26) $\begin{split}&p_{n}q_{n}^{*}-p_{n}^{*}q_{n}\\ &=f_{n}f_{n-1}^{*}f_{n}f_{n-1}^{*}-f_{n}^{*}f_{n-1}f_{n}^{*}f_{n-1}+g_{n}g_{n-1}^{*}g_{n}g_{n-1}^{*}-g_{n}^{*}g_{n-1}g_{n}^{*}g_{n-1}\\ &=\left(f_{n}f_{n-1}^{*}\right)^{2}-\left(f_{n}^{*}f_{n-1}\right)^{2}+\left(g_{n}g_{n-1}^{*}\right)^{2}-\left(g_{n}^{*}g_{n-1}\right)^{2}\\ &=\left(f_{n}f_{n-1}^{*}-f_{n}^{*}f_{n-1}\right)\left(f_{n}f_{n-1}^{*}+f_{n}^{*}f_{n-1}\right)\\ &+\left(g_{n}g_{n-1}^{*}-g_{n}^{*}g_{n-1}\right)\left(g_{n}g_{n-1}^{*}+g_{n}^{*}g_{n-1}\right).\end{split}$
Substituting (5.9), (5.11) and (5.12) into (D.26) gives
(D.27) $\begin{split}&p_{n}q_{n}^{*}-p_{n}^{*}q_{n}\\ &=\omega|G_{n}|^{2}\left(f_{n}f_{n-1}^{*}+f_{n}^{*}f_{n-1}-g_{n}g_{n-1}^{*}-g_{n}^{*}g_{n-1}\right)\\ &=2\left(2\delta_{1}-\beta\right)\omega F_{n}^{2}|G_{n}|^{2}+2\omega\left(\delta_{2}+\delta_{2}^{*}\right)|G_{n}|^{4},\end{split}$
thus (D.21). Also from (D.1), (5.11) and (5.12), we see that
(D.28) $\begin{split}&q_{n}q_{n}^{*}-F_{n}^{2}|G_{n}|^{2}\\ &=\left(g_{n}^{*}f_{n-1}+f_{n}g_{n-1}^{*}\right)\left(g_{n}f_{n-1}^{*}+f_{n}^{*}g_{n-1}\right)\\ &-\left(g_{n}f_{n-1}-f_{n}g_{n-1}\right)\left(g_{n}^{*}f_{n-1}^{*}-f_{n}^{*}g_{n-1}^{*}\right)\\ &=\left(f_{n}f_{n-1}^{*}+f_{n}^{*}f_{n-1}\right)\left(g_{n}g_{n-1}^{*}+g_{n}^{*}g_{n-1}\right)\\ &=\left(2\delta_{1}F_{n}^{2}+\left(\delta_{2}+\delta_{2}^{*}\right)|G_{n}|^{2}\right)\left(2\left(\beta-\delta_{1}\right)F_{n}^{2}-\left(\delta_{2}+\delta_{2}^{*}\right)|G_{n}|^{2}\right)\\ &=4\delta_{1}\left(\beta-\delta_{1}\right)F_{n}^{4}+2\left(\beta-2\delta_{1}\right)\left(\delta_{2}+\delta_{2}^{*}\right)F_{n}^{2}|G_{n}|^{2}-\left(\delta_{2}+\delta_{2}^{*}\right)^{2}|G_{n}|^{4}\end{split}$
holds. From (D.28), we get (D.22). ∎
Using Proposition 5.5, Lemma D.2, Lemma D.3 and Lemma D.4, Theorem 5.6 and Theorem 5.9 can be proved as follows.
D.1. Proof of Theorem 5.6
By differentiating $\gamma$ with respect to $t$, we get
(D.29) $\begin{split}&\dot{\gamma}_{n}(t)=\frac{1}{F_{n}^{2}}\left(\begin{array}[]{c}\vspace{2mm}\displaystyle D_{t}H_{n}\cdot F_{n}+D_{t}H_{n}^{*}\cdot F_{n}\\ \vspace{2mm}\displaystyle\frac{1}{i}\left(D_{t}H_{n}\cdot F_{n}-D_{t}H_{n}^{*}\cdot F_{n}\right)\\ -D_{t}D_{z}F_{n}\cdot F_{n}\end{array}\right)\\ &=\frac{1}{F_{n}^{2}}\left(\begin{array}[]{c}\vspace{2mm}\displaystyle\beta_{1}q_{n}^{*}+i\beta_{2}p_{n}^{*}+\beta_{1}q_{n}-i\beta_{2}p_{n}\\ \vspace{2mm}\displaystyle\frac{1}{i}\left(\beta_{1}q_{n}^{*}+i\beta_{2}p_{n}^{*}-\beta_{1}q_{n}+i\beta_{2}p_{n}\right)\\ -\left(\epsilon_{1}+C\right)F_{n}^{2}-\epsilon_{2}|G_{n}|^{2}\end{array}\right).\end{split}$
On the other hand, from (D.18) and (D.19), we see
(D.30) $\begin{split}&\beta_{1}\left(T_{n}-\tan\frac{\kappa_{n}}{2}\widetilde{N}_{n}\right)+\beta_{2}\tan\frac{\kappa_{n}}{2}\widetilde{B}_{n}\\ &=\frac{1}{2\beta F_{n}^{2}}\left(\begin{array}[]{c}\vspace{2mm}\displaystyle\beta_{1}q_{n}^{*}+i\beta_{2}p_{n}^{*}+\beta_{1}q_{n}-i\beta_{2}p_{n}\\ \vspace{2mm}\displaystyle\frac{1}{i}\left(\beta_{1}q_{n}^{*}+i\beta_{2}p_{n}^{*}-\beta_{1}q_{n}+i\beta_{2}p_{n}\right)\\ -2\beta_{1}\left(\beta-2\delta_{1}\right)F_{n}^{2}+2\beta_{1}\left(\delta_{2}+\delta_{2}^{*}\right)|G_{n}|^{2}+2i\beta_{2}\omega|G_{n}|^{2}\end{array}\right)\end{split}$
holds. From (D.8), (D.30) can be rewritten as follows
(D.31) $\begin{split}&\beta_{1}\left(T_{n}-\tan\frac{\kappa_{n}}{2}\widetilde{N}_{n}\right)+\beta_{2}\tan\frac{\kappa_{n}}{2}\widetilde{B}_{n}\\ &=\frac{1}{2\beta F_{n}^{2}}\left(\begin{array}[]{c}\vspace{2mm}\displaystyle\beta_{1}q_{n}^{*}+i\beta_{2}p_{n}^{*}+\beta_{1}q_{n}-i\beta_{2}p_{n}\\ \vspace{2mm}\displaystyle\frac{1}{i}\left(\beta_{1}q_{n}^{*}+i\beta_{2}p_{n}^{*}-\beta_{1}q_{n}+i\beta_{2}p_{n}\right)\\ -2\beta_{1}\left(\beta-2\delta_{1}\right)F_{n}^{2}-\epsilon_{2}|G_{n}|^{2}\end{array}\right).\end{split}$
Thus (5.15) is obtained.
D.2. Proof of Theorem 5.9
We compute the condition satisfied by $C$ and $\Gamma$ when
(D.32) $\left\langle\dot{\gamma}_{n},\widetilde{B}_{n}\right\rangle=0,$
(D.33) $\left\langle\dot{\gamma}_{n},\dot{\gamma}_{n}\right\rangle=\rho^{2},$
where $\rho\in\mathbb{R}$ is a constant. From (D.19) and (D.29) we see that
(D.34) $\begin{split}2iF_{n}^{3}\widetilde{G}_{n}\left\langle\dot{\gamma}_{n},\widetilde{B}_{n}\right\rangle=&\left(p_{n}-p_{n}^{*}\right)\left(\beta_{1}q_{n}^{*}+i\beta_{2}p_{n}^{*}+\beta_{1}q_{n}-i\beta_{2}p_{n}\right)\\ &+\left(p_{n}+p_{n}^{*}\right)\left(\beta_{1}q_{n}^{*}+i\beta_{2}p_{n}^{*}-\beta_{1}q_{n}+i\beta_{2}p_{n}\right)\\ &+2\omega|G_{n}|^{2}\left\{\left(\epsilon_{1}+C\right)F_{n}^{2}+\epsilon_{2}|G_{n}|^{2}\right\}\\ =&2F_{n}^{2}|G_{n}|^{2}\left\{2\beta_{1}\left(2\delta_{1}-\beta\right)\omega+2i\beta_{2}+\omega\left(\epsilon_{1}+C\right)\right\}\\ &+2\omega|G_{n}|^{4}\left\{2\beta_{1}\left(\delta_{2}+\delta_{2}^{*}\right)+2i\omega\beta_{2}+\epsilon_{2}\right\}\end{split}$
holds. For (D.32) to be satisfied, it is sufficient that the following equations hold.
(D.35) $2\beta_{1}\left(2\delta_{1}-\beta\right)\omega+2i\beta_{2}+\omega\left(\epsilon_{1}+C\right)=0,\$
(D.36) $02\beta_{1}\left(\delta_{2}+\delta_{2}^{*}\right)+2i\omega\beta_{2}+\epsilon_{2}=$ 0 .
By comparing (D.36) and (D.8) we see that (D.35) is a necessary and sufficient condition for the realization of (D.32). By substituting (5.6) into (D.35), it can be rewritten in terms of $C$ and $\Gamma$ as follows.
(D.37) $i\Gamma\left(\omega\left(2\delta_{1}-\beta\right)-\frac{\delta_{2}+\delta_{2}^{*}}{\omega}\right)+\epsilon_{1}+C+\left(\delta_{3}+\delta_{3}^{*}\right)\left(2\delta_{1}-\beta\right)-\frac{\delta_{3}-\delta_{3}^{*}}{\omega}=0.$
Also, since
(D.38) $\begin{split}&F_{n}^{4}\langle\dot{\gamma}_{n},\dot{\gamma}_{n}\rangle\\ =&\left(\beta_{1}q_{n}^{*}+i\beta_{2}p_{n}^{*}+\beta_{1}q_{n}-i\beta_{2}p_{n}\right)^{2}\\ &-\left(\beta_{1}q_{n}^{*}+i\beta_{2}p_{n}^{*}-\beta_{1}q_{n}+i\beta_{2}p_{n}\right)^{2}\\ &+\left\{\left(\epsilon_{1}+C\right)F_{n}^{2}+\epsilon_{2}|G_{n}|^{2}\right\}^{2}\\ =&F_{n}^{4}\left\{16\beta_{1}^{2}\delta_{1}\left(\beta-\delta_{1}\right)+\left(\epsilon_{1}+C\right)^{2}\right\}\\ &+F_{n}^{2}|G_{n}|^{2}\left\{4\beta_{1}^{2}\left(1+2\left(\beta-2\delta_{1}\right)\left(\delta_{2}+\delta_{2}^{*}\right)\right)\right.\\ &\left.-8i\beta_{1}\beta_{2}\omega\left(2\delta_{1}-\beta\right)+4\beta_{2}^{2}+2\epsilon_{2}\left(\epsilon_{1}+C\right)\right\}\\ &+|G_{n}|^{4}\left(-4\beta_{1}^{2}\left(\delta_{2}+\delta_{2}^{*}\right)^{2}-8i\beta_{1}\beta_{2}\omega\left(\delta_{2}+\delta_{2}^{*}\right)+4\beta_{2}^{2}\omega^{2}+\epsilon_{2}^{2}\right),\end{split}$
holds, we see that for (D.33) holds, it is sufficient that
(D.39) $4\beta_{1}^{2}\left(1+2\left(\beta-2\delta_{1}\right)\left(\delta_{2}+\delta_{2}^{*}\right)\right)-8i\beta_{1}\beta_{2}\omega\left(2\delta_{1}-\beta\right)+4\beta_{2}^{2}+2\epsilon_{2}\left(\epsilon_{1}+C\right)=0,$
(D.40) $-4\beta_{1}^{2}\left(\delta_{2}+\delta_{2}^{*}\right)^{2}-8i\beta_{1}\beta_{2}\omega\left(\delta_{2}+\delta_{2}^{*}\right)+4\beta_{2}^{2}\omega^{2}+\epsilon_{2}^{2}=0,$
hold. Since
(D.41) $\begin{split}&4\beta_{1}^{2}\left(\delta_{2}+\delta_{2}^{*}\right)^{2}+8i\beta_{1}\beta_{2}\omega\left(\delta_{2}+\delta_{2}^{*}\right)-4\beta_{2}^{2}\omega^{2}-\epsilon_{2}^{2}\\ =&\left\{2\beta_{1}\left(\delta_{2}+\delta_{2}^{*}\right)+2i\omega\beta_{2}+\epsilon_{2}\right\}\left\{2\beta_{1}\left(\delta_{2}+\delta_{2}^{*}\right)+2i\omega\beta_{2}-\epsilon_{2}\right\},\end{split}$
holds, (D.40) is identically satisfied. Therefore, (D.39) is a necessary and sufficient condition for the realization of (D.33). By using (D.8), (5.6) and (D.37), (D.39) can be written in terms of $\Gamma$ as follows.
(D.42) $i\delta_{2}\delta_{2}^{*}\left(\delta_{2}-\delta_{2}^{*}\right)\Gamma^{2}+2\delta_{2}\delta_{2}^{*}\left(\delta_{3}+\delta_{3}^{*}\right)\Gamma+i\left(\delta_{2}{\delta_{3}^{*}}^{2}-\delta_{2}^{*}\delta_{3}^{2}\right)=0.$
(D.42) has the following solutions.
(D.43) $\Gamma=\frac{-\delta_{2}\delta_{2}^{*}\left(\delta_{3}+\delta_{3}^{*}\right)\pm\sqrt{\delta_{2}\delta_{2}^{*}\left(\delta_{2}\delta_{3}^{*}+\delta_{2}^{*}\delta_{3}\right)^{2}}}{i\delta_{2}\delta_{2}^{*}\left(\delta_{2}-\delta_{2}^{*}\right)}.$
Since
(D.44) $\begin{split}\delta_{2}\delta_{2}^{*}\left(\delta_{2}\delta_{3}^{*}+\delta_{2}^{*}\delta_{3}\right)^{2}=\left(\delta_{2}\delta_{2}^{*}\right)^{2}\left(\delta_{3}e^{i\lambda}+\delta_{3}^{*}e^{-i\lambda}\right)^{2}=\left(\delta_{2}\delta_{2}^{*}\right)^{2}\left(i\frac{\vartheta_{1}^{{}^{\prime}}(0)}{\vartheta_{1}(iv)}\right)^{2},\end{split}$
holds, (D.43) can be rewritten as
(D.45) $\Gamma=\frac{-\left(\delta_{3}+\delta_{3}^{*}\right)\pm\left(\delta_{3}e^{i\lambda}+\delta_{3}^{*}e^{-i\lambda}\right)}{i\omega}.$
Now, we prove equations (5.20) and (5.21). From (4.2) and (5.6), we see that when $C_{+}$ and $\Gamma_{+}$ are chosen as in (5.18), the following relation holds
(D.46) $16\beta_{1}^{2}\left(\beta-\delta_{1}\right)\delta_{1}=4\left(i\frac{\vartheta_{1}^{{}^{\prime}}(0)}{\vartheta_{1}(iv)}\right)^{2}\frac{u_{2}^{2}}{\vartheta_{3}(2r)^{2}\vartheta_{3}(0)^{6}}.$
Also, from (4.2), (5.6), (D.35) and (D.36), we see that when $C_{+}$ and $\Gamma_{+}$ are chosen as in (5.18), the following relation holds
(D.47) $\left(\epsilon_{1}+C\right)^{2}=\left(i\frac{\vartheta_{1}^{{}^{\prime}}(0)}{\vartheta_{1}(iv)}\right)^{2}\frac{u_{2}^{2}}{\vartheta_{3}(2r)^{2}\vartheta_{3}(0)^{6}}\frac{\left(\vartheta_{3}(r)^{4}-\vartheta_{1}(r)^{4}\right)^{2}}{\vartheta_{3}(r)^{4}\vartheta_{1}(r)^{4}}.$
From (D.38), (D.46) and (D.47), we see that
(D.48) $\rho^{2}=16\beta_{1}^{2}\left(\beta-\delta_{1}\right)\delta_{1}+\left(\epsilon_{1}+C\right)^{2}=\left(i\frac{\vartheta_{1}^{{}^{\prime}}(0)}{\vartheta_{1}(iv)}\right)^{2}\frac{u_{2}^{2}}{\vartheta_{3}(r)^{4}\vartheta_{1}(r)^{4}}=\rho_{+}^{2}$
holds. (D.48) is equivalent to (5.20). From (4.2) and (5.6), we see that when $C_{-}$ and $\Gamma_{-}$ are chosen as in (5.18), the following relation holds
(D.49) $16\beta_{1}^{2}\left(\beta-\delta_{1}\right)\delta_{1}=4\left(i\frac{\vartheta_{1}^{{}^{\prime}}(0)}{\vartheta_{1}(iv)}\right)^{2}\frac{u_{2}^{2}}{\vartheta_{3}(2r)^{2}\vartheta_{3}(0)^{6}}.$
Also, from (4.2), (5.6), (D.35) and (D.36), we see that when $C_{-}$ and $\Gamma_{-}$ are chosen as in (5.18), the following relation holds
(D.50) $\left(\epsilon_{1}+C\right)^{2}=\left(i\frac{\vartheta_{1}^{{}^{\prime}}(0)}{\vartheta_{1}(iv)}\right)^{2}\frac{u_{2}^{2}}{\vartheta_{3}(2r)^{2}\vartheta_{3}(0)^{6}}\frac{\left(\vartheta_{4}(r)^{4}-\vartheta_{2}(r)^{4}\right)^{2}}{\vartheta_{4}(r)^{4}\vartheta_{2}(r)^{4}}.$
From (D.38), (D.49) and (D.50), we see that
(D.51) $\rho^{2}=16\beta_{1}^{2}\left(\beta-\delta_{1}\right)\delta_{1}+\left(\epsilon_{1}+C\right)^{2}=\left(i\frac{\vartheta_{1}^{{}^{\prime}}(0)}{\vartheta_{1}(iv)}\right)^{2}\frac{u_{2}^{2}}{\vartheta_{4}(r)^{4}\vartheta_{2}(r)^{4}}=\rho_{-}^{2}$
holds. (D.51) is equivalent to (5.21).
Appendix E Proof of Proposition 5.12
Here we prove that the signed curvature angle $\kappa$ satisfies (5.27). First, we see
Lemma E.1.
The $\tau$ functions (5.1) satisfy the following bilinear equations.
(E.1) $\widetilde{G}_{n}^{2}+\beta^{2}F_{n}^{2}=F_{n+1}F_{n-1},$
(E.2) $\beta^{2}\widetilde{G}_{n}^{2}-\frac{\vartheta_{1}\left(iv\right)^{4}}{\vartheta_{3}\left(iv\right)^{4}}\beta^{4}F_{n}^{2}=\widetilde{G}_{n+1}\widetilde{G}_{n-1}.$
Proof.
Using (3.1)-(3.3), we see that (E.1) hold. Putting
(E.7) $\left\{\begin{array}[]{l}\vspace{2mm}X=0,\\ \vspace{2mm}Y=0,\\ \vspace{2mm}U=iv\left(n+\frac{1}{2}\right)+it+\frac{z}{\Delta_{3}-\Delta_{1}},\\ V=-iv\left(n-\frac{3}{2}\right)-it-\frac{z}{\Delta_{3}-\Delta_{1}},\\ \end{array}\right.$
in (A.10), we obtain (E.2).
E.1. Proof of Proposition 5.12
Using (E.1) and (E.2), we see that (5.27) holds.
Appendix F Proof of Theorem 5.13
Here, we show that the signed curvature angle satisfies the semi-discrete mKdV equation (5.28). For the proof of Theorem 5.13, we first see
Lemma F.1.
The $\tau$ functions (5.1) satisfy the following bilinear equation.
(F.1) $D_{t}\widetilde{G}_{n}\cdot F_{n}=\eta_{3}\left(\widetilde{G}_{n+1}F_{n-1}-\widetilde{G}_{n-1}F_{n+1}\right).$
Proof.
We assume $z=0$ without loss of generality. After differentiating the equation (A.3) with
(F.6) $\left\{\begin{array}[]{l}\vspace{2mm}X=iv\left(n+\frac{1}{2}\right)+it,\\ \vspace{2mm}Y=iv+h,\\ \vspace{2mm}U=-iv+h,\\ V=iv\left(n-\frac{3}{2}\right)+it\\ \end{array}\right.$
by $h$ and putting $h=0$, we obtain
(F.7) $\begin{split}&D_{t}\left\{-i\vartheta_{1}(iv)\vartheta_{4}\left(iv\left(n-\frac{1}{2}\right)+it\right)\right\}\cdot\left\{\vartheta_{3}(iv)\vartheta_{2}\left(iv\left(n-\frac{1}{2}\right)+it\right)\right\}\\ &=\frac{\vartheta_{1}(iv)\vartheta_{3}(iv)}{\vartheta_{1}(2iv)\vartheta_{3}(0)}\frac{\vartheta^{{}^{\prime}}_{1}\left(iv\right)\vartheta_{3}\left(iv\right)-\vartheta_{1}\left(iv\right)\vartheta^{{}^{\prime}}_{3}\left(iv\right)}{i\vartheta_{1}\left(iv\right)\vartheta_{3}\left(iv\right)}\\ &\times\left\{-i\vartheta_{1}(iv)\vartheta_{4}\left(iv\left(n-\frac{3}{2}\right)+it\right)\right\}\left\{\vartheta_{3}(iv)\vartheta_{2}\left(iv\left(n+\frac{1}{2}\right)+it\right)\right\}\\ &-\frac{\vartheta_{1}(iv)\vartheta_{3}(iv)}{\vartheta_{1}(2iv)\vartheta_{3}(0)}\frac{\vartheta^{{}^{\prime}}_{1}\left(iv\right)\vartheta_{3}\left(iv\right)-\vartheta_{1}\left(iv\right)\vartheta^{{}^{\prime}}_{3}\left(iv\right)}{i\vartheta_{1}\left(iv\right)\vartheta_{3}\left(iv\right)}\\ &\times\left\{-i\vartheta_{1}(iv)\vartheta_{4}\left(iv\left(n+\frac{1}{2}\right)+it\right)\right\}\left\{\vartheta_{3}(iv)\vartheta_{2}\left(iv\left(n-\frac{3}{2}\right)+it\right)\right\}.\end{split}$
After rearranging (F.7), we obtain (F.1).
F.1. Proof of Theorem 5.13
Using (E.1) and (F.1), we see that the signed curvature angle satisfies the semi-discrete mKdV equation (5.28).
Appendix G Proof of Theorem 5.14
From (A.17), (A.18) and (A.19), we see that
(G.1) $\begin{split}&\kappa_{n}=2\arctan\left(-i\frac{\vartheta_{1}(iv;iy)\vartheta_{4}\left(iv\left(n-\frac{1}{2}\right)+it;iy\right)}{\vartheta_{3}(iv;iy)\vartheta_{2}\left(iv\left(n-\frac{1}{2}\right)+it;iy\right)}\right)\\ &=2\arctan\left(-i\frac{\frac{\vartheta_{1}\left(i\frac{v}{2}\left(n+\frac{1}{2}\right)+\frac{it}{2};\frac{iy}{2}\right)}{\vartheta_{2}\left(i\frac{v}{2}\left(n+\frac{1}{2}\right)+\frac{it}{2};\frac{iy}{2}\right)}-\frac{\vartheta_{1}\left(i\frac{v}{2}\left(n-\frac{3}{2}\right)+\frac{it}{2};\frac{iy}{2}\right)}{\vartheta_{2}\left(i\frac{v}{2}\left(n-\frac{3}{2}\right)+\frac{it}{2};\frac{iy}{2}\right)}}{1-\frac{\vartheta_{1}\left(i\frac{v}{2}\left(n+\frac{1}{2}\right)+\frac{it}{2};\frac{iy}{2}\right)\vartheta_{1}\left(i\frac{v}{2}\left(n-\frac{3}{2}\right)+\frac{it}{2};\frac{iy}{2}\right)}{\vartheta_{2}\left(i\frac{v}{2}\left(n+\frac{1}{2}\right)+\frac{it}{2};\frac{iy}{2}\right)\vartheta_{2}\left(i\frac{v}{2}\left(n-\frac{3}{2}\right)+\frac{it}{2};\frac{iy}{2}\right)}}\right)\\ &=2\arctan\left(\frac{\frac{\vartheta_{1}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y};\frac{2i}{y}\right)}{\vartheta_{4}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y};\frac{2i}{y}\right)}-\frac{\vartheta_{1}\left(\frac{v}{y}\left(n-\frac{3}{2}\right)+\frac{t}{y};\frac{2i}{y}\right)}{\vartheta_{4}\left(\frac{v}{y}\left(n-\frac{3}{2}\right)+\frac{t}{y};\frac{2i}{y}\right)}}{1+\frac{\vartheta_{1}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y};\frac{2i}{y}\right)\vartheta_{1}\left(\frac{v}{y}\left(n-\frac{3}{2}\right)+\frac{t}{y};\frac{2i}{y}\right)}{\vartheta_{4}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y};\frac{2i}{y}\right)\vartheta_{4}\left(\frac{v}{y}\left(n-\frac{3}{2}\right)+\frac{t}{y};\frac{2i}{y}\right)}}\right)\\ &=2\arctan\left(\frac{\vartheta_{1}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y};\frac{2i}{y}\right)}{\vartheta_{4}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y};\frac{2i}{y}\right)}\right)\\ &-2\arctan\left(\frac{\vartheta_{1}\left(\frac{v}{y}\left(n-\frac{3}{2}\right)+\frac{t}{y};\frac{2i}{y}\right)}{\vartheta_{4}\left(\frac{v}{y}\left(n-\frac{3}{2}\right)+\frac{t}{y};\frac{2i}{y}\right)}\right)\\ &=\frac{\Theta_{n+1}-\Theta_{n-1}}{2}\end{split}$
holds.
Appendix H Proof of Theorem 5.15
We first construct the solution to the semi-discrete potential mKdV equation (5.33) in two stages by proving Lemma H.1 and Lemma H.2. Then we construct the solution to the semi-discrete sine-Gordon equation (5.34) in two stages by proving Lemma H.3 and Lemma H.4. This is based on the construction of the solution to the mKdV equation using the theory of $\tau$ functions [15].
For fixed $v,y$, define the following real-valued functions
(H.1) $\begin{split}&A_{n}\left(t\right)=\vartheta_{4}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y};\frac{2i}{y}\right),\\ &B_{n}\left(t\right)=\vartheta_{1}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y};\frac{2i}{y}\right).\end{split}$
Then the potential function can be written as follows
(H.2) $\Theta_{n}\left(t\right)=4\arctan\frac{B_{n}}{A_{n}}=\frac{2}{i}\log\frac{A_{n}+iB_{n}}{A_{n}-iB_{n}}=\frac{2}{i}\log\frac{\widetilde{\tau}_{n}}{\widetilde{\tau}^{*}_{n}},$
where $\widetilde{\tau}_{n}=A_{n}+iB_{n}$. $\widetilde{\tau}$ corresponds to the $\tau$ functions of the one-component KP hierarchy, while $A$ and $B$ correspond to the $\tau$ functions of the two-component KP hierarchy.
Lemma H.1.
Assuming that the solution of (5.33) is expressed as (H.2), we obtain the following bilinear equations as a sufficient condition for satisfying (5.33).
(H.3) $D_{t}A_{n+1}\cdot A_{n}+D_{t}B_{n+1}\cdot B_{n}=\frac{L_{1}}{2y}\left(\eta_{1}-1\right)\left(A_{n+1}A_{n}+B_{n+1}B_{n}\right),$
(H.4) $D_{t}A_{n+1}\cdot B_{n}-D_{t}B_{n+1}\cdot A_{n}=\frac{L_{1}}{2y}\left(\eta_{1}+1\right)\left(A_{n+1}B_{n}-B_{n+1}A_{n}\right),$
where $\eta_{1}$ is a real constant.
Proof.
Since
(H.5) $\begin{split}&\frac{d}{dt}\left(\frac{\Theta_{n+1}+\Theta_{n}}{2}\right)=\frac{1}{i}\frac{d}{dt}\left(\log\frac{\widetilde{\tau}_{n+1}}{\widetilde{\tau}^{*}_{n+1}}+\log\frac{\widetilde{\tau}_{n}}{\widetilde{\tau}^{*}_{n}}\right)\\ &=\frac{1}{i}\frac{d}{dt}\left(\log\frac{\widetilde{\tau}_{n+1}}{\widetilde{\tau}^{*}_{n}}-\log\frac{\widetilde{\tau}^{*}_{n+1}}{\widetilde{\tau}_{n}}\right)\\ &=\frac{1}{i}\left(\frac{D_{t}\widetilde{\tau}_{n+1}\cdot\widetilde{\tau}^{*}_{n}}{\widetilde{\tau}_{n+1}\widetilde{\tau}^{*}_{n}}-\frac{D_{t}\widetilde{\tau}^{*}_{n+1}\cdot\widetilde{\tau}_{n}}{\widetilde{\tau}^{*}_{n+1}\widetilde{\tau}_{n}}\right),\end{split}$
(H.6) $\begin{split}&\sin\left(\frac{\Theta_{n+1}-\Theta_{n}}{2}\right)=\frac{1}{2i}\left(\frac{\widetilde{\tau}_{n+1}\widetilde{\tau}^{*}_{n}}{\widetilde{\tau}^{*}_{n+1}\widetilde{\tau}_{n}}-\frac{\widetilde{\tau}^{*}_{n+1}\widetilde{\tau}_{n}}{\widetilde{\tau}_{n+1}\widetilde{\tau}^{*}_{n}}\right)\end{split}$
hold, it can be seen that a sufficient condition for (H.2) to be a solution of (5.33) is that
(H.7) $\begin{split}&D_{t}\widetilde{\tau}_{n+1}\cdot\widetilde{\tau}^{*}_{n}=-\frac{L_{1}}{2y}\widetilde{\tau}^{*}_{n+1}\widetilde{\tau}_{n}+\frac{L_{1}}{2y}\eta_{1}\widetilde{\tau}_{n+1}\widetilde{\tau}_{n}^{*},\end{split}$
is satisfied, where $\eta_{1}$ is a real constant. Substituting $\widetilde{\tau}_{n}=A_{n}+iB_{n}$ for this, we get
(H.8) $\begin{split}&D_{t}\left(A_{n+1}\cdot A_{n}+B_{n+1}\cdot B_{n}\right)-iD_{t}\left(A_{n+1}\cdot B_{n}-B_{n+1}\cdot A_{n}\right)\\ &=\frac{L_{1}}{2y}\left(\eta_{1}-1\right)\left(A_{n+1}A_{n}+B_{n+1}B_{n}\right)-i\frac{L_{1}}{2y}\left(\eta_{1}+1\right)\left(A_{n+1}B_{n}-B_{n+1}A_{n}\right).\end{split}$
The real part of this equation yields (H.3), and the imaginary part yields (H.4). ∎
In Lemma H.2, we construct a solution of (H.3) and (H.4) in terms of elliptic theta functions.
Lemma H.2.
The real functions $A$ and $B$ defined by (H.1) are the solutions of (H.3) and (H.4) with
(H.9) $\eta_{1}=\frac{d_{1}\left(\frac{v}{y};\frac{2}{y}\right)+d_{4}\left(\frac{v}{y};\frac{2}{y}\right)-Q_{1}(v,y)+Q_{4}(v,y)}{L_{1}(v,y)}.$
Proof.
In this proof, all elliptic theta functions have $w=2i/y$ and we denote $\vartheta_{j}(v;2i/y)$ by $\vartheta_{j}(v)$. After differentiating the equation in (A.6) with
(H.14) $\left\{\begin{array}[]{l}\vspace{2mm}X=\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y},\\ \vspace{2mm}Y=\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y},\\ \vspace{2mm}U=\frac{v}{y}+h,\\ V=h,\\ \end{array}\right.$
by $h$ and putting $h=0$, we obtain
(H.15) $\begin{split}&D_{t}\vartheta_{4}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\cdot\vartheta_{4}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &=\frac{1}{y}\frac{\vartheta^{{}^{\prime}}_{4}\left(\frac{v}{y}\right)}{\vartheta_{4}\left(\frac{v}{y}\right)}\vartheta_{4}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{4}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &-\frac{1}{y}\frac{\vartheta_{1}\left(\frac{v}{y}\right)\vartheta^{{}^{\prime}}_{1}\left(0\right)}{\vartheta_{4}\left(\frac{v}{y}\right)\vartheta_{4}\left(0\right)}\vartheta_{1}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{1}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &=\frac{1}{y}d_{4}\left(\frac{v}{y};\frac{2}{y}\right)\vartheta_{4}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{4}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &-\frac{1}{y}Q_{1}(v,y)\vartheta_{1}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{1}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right).\end{split}$
Also, after differentiating the equation in (A.7) with
(H.20) $\left\{\begin{array}[]{l}\vspace{2mm}X=\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y},\\ \vspace{2mm}Y=\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y},\\ \vspace{2mm}U=\frac{v}{y}+h,\\ V=h,\\ \end{array}\right.$
by $h$ and putting $h=0$, we obtain
(H.21) $\begin{split}&D_{t}\vartheta_{1}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\cdot\vartheta_{1}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &=\frac{1}{y}\frac{\vartheta^{{}^{\prime}}_{4}\left(\frac{v}{y}\right)}{\vartheta_{4}\left(\frac{v}{y}\right)}\vartheta_{1}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{1}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &-\frac{1}{y}\frac{\vartheta_{1}\left(\frac{v}{y}\right)\vartheta^{{}^{\prime}}_{1}\left(0\right)}{\vartheta_{4}\left(\frac{v}{y}\right)\vartheta_{4}\left(0\right)}\vartheta_{4}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{4}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &=\frac{1}{y}d_{4}\left(\frac{v}{y};\frac{2}{y}\right)\vartheta_{1}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{1}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &-\frac{1}{y}Q_{1}(v,y)\vartheta_{4}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{4}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right).\end{split}$
Also, after differentiating the equation in (A.7) with
(H.26) $\left\{\begin{array}[]{l}\vspace{2mm}X=\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y},\\ \vspace{2mm}Y=\frac{v}{y}+h,\\ \vspace{2mm}U=\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y},\\ V=h,\\ \end{array}\right.$
by $h$ and putting $h=0$, we obtain
(H.27) $\begin{split}&D_{t}\vartheta_{1}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\cdot\vartheta_{4}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &=\frac{1}{y}\frac{\vartheta^{{}^{\prime}}_{1}\left(\frac{v}{y}\right)}{\vartheta_{1}\left(\frac{v}{y}\right)}\vartheta_{1}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{4}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &-\frac{1}{y}\frac{\vartheta_{4}\left(\frac{v}{y}\right)\vartheta^{{}^{\prime}}_{1}\left(0\right)}{\vartheta_{1}\left(\frac{v}{y}\right)\vartheta_{4}\left(0\right)}\vartheta_{4}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{1}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &=\frac{1}{y}d_{1}\left(\frac{v}{y};\frac{2}{y}\right)\vartheta_{1}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{4}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &-\frac{1}{y}Q_{2}(v,y)\vartheta_{4}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{1}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right).\end{split}$
Also, after differentiating the equation in (A.7) with
(H.32) $\left\{\begin{array}[]{l}\vspace{2mm}X=\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y},\\ \vspace{2mm}Y=\frac{v}{y}+h,\\ \vspace{2mm}U=\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y},\\ V=h,\\ \end{array}\right.$
by $h$ and putting $h=0$, we obtain
(H.33) $\begin{split}&D_{t}\vartheta_{4}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\cdot\vartheta_{1}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &=\frac{1}{y}\frac{\vartheta^{{}^{\prime}}_{1}\left(\frac{v}{y}\right)}{\vartheta_{1}\left(\frac{v}{y}\right)}\vartheta_{4}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{1}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &-\frac{1}{y}\frac{\vartheta_{4}\left(\frac{v}{y}\right)\vartheta^{{}^{\prime}}_{1}\left(0\right)}{\vartheta_{1}\left(\frac{v}{y}\right)\vartheta_{4}\left(0\right)}\vartheta_{1}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{4}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &=\frac{1}{y}d_{1}\left(\frac{v}{y};\frac{2}{y}\right)\vartheta_{4}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{1}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &-\frac{1}{y}Q_{2}(v,y)\vartheta_{1}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{4}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right).\end{split}$
Adding (H.15) and (H.21) yields (H.3), and subtracting (H.27) from (H.33) yields (H.4). ∎
Next, we construct the solution to the semi-discrete sine-Gordon equation (5.34).
Lemma H.3.
Assuming that the solution of (5.34) is expressed as (H.2), we obtain the following bilinear equations as a sufficient condition for satisfying (5.34).
(H.34) $D_{t}A_{n+1}\cdot A_{n}-D_{t}B_{n+1}\cdot B_{n}=\frac{L_{2}}{2y}\left(\eta_{2}-1\right)\left(A_{n+1}A_{n}-B_{n+1}B_{n}\right),$
(H.35) $D_{t}A_{n+1}\cdot B_{n}+D_{t}B_{n+1}\cdot A_{n}=\frac{L_{2}}{2y}\left(\eta_{2}+1\right)\left(A_{n+1}B_{n}+B_{n+1}A_{n}\right),$
where $\eta_{2}$ is a real constant.
Proof.
Since
(H.36) $\begin{split}&\frac{d}{dt}\left(\frac{\Theta_{n+1}-\Theta_{n}}{2}\right)=\frac{1}{i}\frac{d}{dt}\left(\log\frac{\widetilde{\tau}_{n+1}}{\widetilde{\tau}^{*}_{n+1}}-\log\frac{\widetilde{\tau}_{n}}{\widetilde{\tau}^{*}_{n}}\right)\\ &=\frac{1}{i}\frac{d}{dt}\left(\log\frac{\widetilde{\tau}_{n+1}}{\widetilde{\tau}_{n}}-\log\frac{\widetilde{\tau}^{*}_{n+1}}{\widetilde{\tau}^{*}_{n}}\right)\\ &=\frac{1}{i}\left(\frac{D_{t}\widetilde{\tau}_{n+1}\cdot\widetilde{\tau}_{n}}{\widetilde{\tau}_{n+1}\widetilde{\tau}_{n}}-\frac{D_{t}\widetilde{\tau}^{*}_{n+1}\cdot\widetilde{\tau}^{*}_{n}}{\widetilde{\tau}^{*}_{n+1}\widetilde{\tau}^{*}_{n}}\right),\end{split}$
(H.37) $\begin{split}&\sin\left(\frac{\Theta_{n+1}+\Theta_{n}}{2}\right)=\frac{1}{2i}\left(\frac{\widetilde{\tau}_{n+1}\widetilde{\tau}_{n}}{\widetilde{\tau}^{*}_{n+1}\widetilde{\tau}^{*}_{n}}-\frac{\widetilde{\tau}^{*}_{n+1}\widetilde{\tau}^{*}_{n}}{\widetilde{\tau}_{n+1}\widetilde{\tau}_{n}}\right)\end{split}$
hold, it can be seen that a sufficient condition for (H.2) to be a solution of (5.34) is that
(H.38) $\begin{split}&D_{t}\widetilde{\tau}_{n+1}\cdot\widetilde{\tau}_{n}=-\frac{L_{2}}{2y}\widetilde{\tau}^{*}_{n+1}\widetilde{\tau}^{*}_{n}+\frac{L_{2}}{2y}\eta_{2}\widetilde{\tau}_{n+1}\widetilde{\tau}_{n},\end{split}$
is satisfied, where $\eta_{2}$ is a real constant. Substituting $\widetilde{\tau}_{n}=A_{n}+iB_{n}$ for this, we get
(H.39) $\begin{split}&D_{t}\left(A_{n+1}\cdot A_{n}-B_{n+1}\cdot B_{n}\right)+iD_{t}\left(A_{n+1}\cdot B_{n}+B_{n+1}\cdot A_{n}\right)\\ &=\frac{L_{2}}{2y}\left(\eta_{2}-1\right)\left(A_{n+1}A_{n}-B_{n+1}B_{n}\right)\\ &+i\frac{L_{2}}{2y}\left(\eta_{2}+1\right)\left(A_{n+1}B_{n}+B_{n+1}A_{n}\right).\end{split}$
The real part of (H.39) yields (H.34), and the imaginary part yields (H.35). ∎
In Lemma H.4, we construct a solution of (H.34) and (H.35) in terms of elliptic theta functions.
Lemma H.4.
The real functions $A$ and $B$ defined by (H.1) are the solutions of (H.34) and (H.35) with
(H.40) $\eta_{2}=\frac{d_{1}\left(\frac{v}{y};\frac{2}{y}\right)+d_{4}\left(\frac{v}{y};\frac{2}{y}\right)+Q_{1}(v,y)-Q_{4}(v,y)}{L_{2}(v,y)}.$
Proof.
In this proof, we set $w=2i/y$ similarly to Lemma H.2. Subtracting (H.21) from (H.15) yields (H.34), and adding (H.33) and (H.27) yields (H.35). ∎
Appendix I Proof of Proposition 5.16
We first show (5.40) and (5.39) by using the periodicity of elliptic theta functions (A.11). Consider replacing the parameter with $\tilde{v}=v+y$. Then we see
(I.1) $\begin{split}&Q_{1}(\tilde{v},y)=-Q_{1}(v,y),\quad Q_{2}(\tilde{v},y)=-Q_{2}(v,y),\\ &d_{1}(\tilde{v},y)=d_{1}(v,y),\quad d_{4}(\tilde{v},y)=d_{4}(v,y),\\ &L_{1}(\tilde{v},y)=L_{2}(v,y).\end{split}$
and
(I.2) $\begin{split}&\Theta_{n}\left(t;\tilde{v},y\right)=4\arctan\frac{\vartheta_{1}\left(\frac{\tilde{v}}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)}{\vartheta_{4}\left(\frac{\tilde{v}}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)}\\ &=4\arctan\frac{\vartheta_{1}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}+n-\frac{1}{2}\right)}{\vartheta_{4}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}+n-\frac{1}{2}\right)}\\ &=(-1)^{n}4\arctan\frac{\vartheta_{1}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}-\frac{1}{2}\right)}{\vartheta_{4}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}-\frac{1}{2}\right)}\\ &=(-1)^{n}\Theta_{n}\left(t-\frac{y}{2};v,y\right).\end{split}$
From the relation between the signed curvature angle $\kappa$ and the potential function $\Theta$, we see
(I.3) $\begin{split}&\kappa_{n}\left(t;\tilde{v}\right)=(-1)^{n+1}\kappa_{n}\left(t-\frac{y}{2};v\right).\end{split}$
Next, we check the transformation of $\lambda$. From (A.13) and (A.14), we have
(I.4) $\begin{split}&\lambda\left(\tilde{v},\frac{1}{2}-r\right)\\ &=\frac{1}{i}\log\left(\frac{\vartheta_{1}\left(-\frac{1}{2}iv+\frac{1}{2}-r-\frac{1}{2}iy\right)}{\vartheta_{1}\left(\frac{1}{2}iv+\frac{1}{2}-r+\frac{1}{2}iy\right)}\frac{\vartheta_{3}\left(\frac{1}{2}iv+\frac{1}{2}-r+\frac{1}{2}iy\right)}{\vartheta_{3}\left(-\frac{1}{2}iv+\frac{1}{2}-r-\frac{1}{2}iy\right)}\right)\\ &=\frac{1}{i}\log\left(\frac{\vartheta_{2}\left(-\frac{1}{2}iv-r-\frac{1}{2}iy\right)}{\vartheta_{2}\left(\frac{1}{2}iv-r+\frac{1}{2}iy\right)}\frac{\vartheta_{4}\left(\frac{1}{2}iv-r+\frac{1}{2}iy\right)}{\vartheta_{4}\left(-\frac{1}{2}iv-r-\frac{1}{2}iy\right)}\right)\\ &=\frac{1}{i}\log\left(-\frac{\vartheta_{3}\left(-\frac{1}{2}iv-r\right)}{\vartheta_{3}\left(\frac{1}{2}iv-r\right)}\frac{\vartheta_{1}\left(\frac{1}{2}iv-r\right)}{\vartheta_{1}\left(-\frac{1}{2}iv-r\right)}\right)\\ &=\pi+\frac{1}{i}\log\left(\frac{\vartheta_{1}\left(-\frac{1}{2}iv+r\right)}{\vartheta_{1}\left(\frac{1}{2}iv+r\right)}\frac{\vartheta_{3}\left(\frac{1}{2}iv+r\right)}{\vartheta_{3}\left(-\frac{1}{2}iv+r\right)}\right)\\ &=\pi+\lambda\left(v,r\right).\\ \end{split}$
Finally, we see that the transformation of the normal $\widetilde{N}$ and the binormal $\widetilde{B}$ follows from the Frenet-Serret formula (2.6).
Appendix J Proof of Proposition 5.18
First, we calculate $L_{1}(v,y)/y$ and $L_{2}(v,y)/y$. From (5.36), (A.15), (A.16) and (A.20), we see
(J.1) $\begin{split}\frac{L_{1}(v,y)}{y}&=\frac{1}{y}\frac{\vartheta_{1}^{{}^{\prime}}\left(\frac{v}{y};i\frac{2}{y}\right)\vartheta_{4}\left(\frac{v}{y};i\frac{2}{y}\right)-\vartheta_{1}\left(\frac{v}{y};i\frac{2}{y}\right)\vartheta_{4}^{{}^{\prime}}\left(\frac{v}{y};i\frac{2}{y}\right)}{\vartheta_{1}\left(\frac{v}{y};i\frac{2}{y}\right)\vartheta_{4}\left(\frac{v}{y};i\frac{2}{y}\right)}\\ &+\frac{1}{y}\frac{\vartheta_{1}^{{}^{\prime}}\left(0;i\frac{2}{y}\right)\vartheta_{1}\left(\frac{v}{y};i\frac{2}{y}\right)}{\vartheta_{4}\left(0;i\frac{2}{y}\right)\vartheta_{4}\left(\frac{v}{y};i\frac{2}{y}\right)}+\frac{1}{y}\frac{\vartheta_{1}^{{}^{\prime}}\left(0;i\frac{2}{y}\right)\vartheta_{4}\left(\frac{v}{y};i\frac{2}{y}\right)}{\vartheta_{4}\left(0;i\frac{2}{y}\right)\vartheta_{1}\left(\frac{v}{y};i\frac{2}{y}\right)}\\ &=i\frac{\vartheta_{1}^{{}^{\prime}}\left(0;iy\right)}{\vartheta_{4}\left(0;iy\right)}\frac{\vartheta_{4}\left(iv;iy\right)}{\vartheta_{1}\left(iv;iy\right)}+i\frac{\vartheta_{1}^{{}^{\prime}}\left(0;iy\right)}{\vartheta_{2}\left(0;iy\right)}\frac{\vartheta_{2}\left(iv;iy\right)}{\vartheta_{1}\left(iv;iy\right)}\\ &=i\frac{\vartheta_{1}^{{}^{\prime}}\left(0;iy\right)}{\vartheta_{1}\left(iv;iy\right)}\frac{\vartheta_{4}\left(iv;iy\right)\vartheta_{2}\left(0;iy\right)+\vartheta_{2}\left(iv;iy\right)\vartheta_{4}\left(0;iy\right)}{\vartheta_{4}\left(0;iy\right)\vartheta_{2}\left(0;iy\right)}.\end{split}$
Similarly, from (5.37), (A.15), (A.16) and (A.20), we see
(J.2) $\begin{split}\frac{L_{2}(v,y)}{y}&=\frac{1}{y}\frac{\vartheta_{1}^{{}^{\prime}}\left(\frac{v}{y};i\frac{2}{y}\right)\vartheta_{4}\left(\frac{v}{y};i\frac{2}{y}\right)-\vartheta_{1}\left(\frac{v}{y};i\frac{2}{y}\right)\vartheta_{4}^{{}^{\prime}}\left(\frac{v}{y};i\frac{2}{y}\right)}{\vartheta_{1}\left(\frac{v}{y};i\frac{2}{y}\right)\vartheta_{4}\left(\frac{v}{y};i\frac{2}{y}\right)}\\ &-\frac{1}{y}\frac{\vartheta_{1}^{{}^{\prime}}\left(0;i\frac{2}{y}\right)\vartheta_{1}\left(\frac{v}{y};i\frac{2}{y}\right)}{\vartheta_{4}\left(0;i\frac{2}{y}\right)\vartheta_{4}\left(\frac{v}{y};i\frac{2}{y}\right)}-\frac{1}{y}\frac{\vartheta_{1}^{{}^{\prime}}\left(0;i\frac{2}{y}\right)\vartheta_{4}\left(\frac{v}{y};i\frac{2}{y}\right)}{\vartheta_{4}\left(0;i\frac{2}{y}\right)\vartheta_{1}\left(\frac{v}{y};i\frac{2}{y}\right)}\\ &=i\frac{\vartheta_{1}^{{}^{\prime}}\left(0;iy\right)}{\vartheta_{4}\left(0;iy\right)}\frac{\vartheta_{4}\left(iv;iy\right)}{\vartheta_{1}\left(iv;iy\right)}-i\frac{\vartheta_{1}^{{}^{\prime}}\left(0;iy\right)}{\vartheta_{2}\left(0;iy\right)}\frac{\vartheta_{2}\left(iv;iy\right)}{\vartheta_{1}\left(iv;iy\right)}\\ &=i\frac{\vartheta_{1}^{{}^{\prime}}\left(0;iy\right)}{\vartheta_{1}\left(iv;iy\right)}\frac{\vartheta_{4}\left(iv;iy\right)\vartheta_{2}\left(0;iy\right)-\vartheta_{2}\left(iv;iy\right)\vartheta_{4}\left(0;iy\right)}{\vartheta_{4}\left(0;iy\right)\vartheta_{2}\left(0;iy\right)}.\end{split}$
On the other hand, from (4.2) and (5.2), we see
(J.3) $\begin{split}&1+\cos\lambda\\ &=\frac{\left\{\vartheta_{1}\left(\frac{iv}{2}+r;iy\right)\vartheta_{3}\left(-\frac{iv}{2}+r;iy\right)+\vartheta_{1}\left(-\frac{iv}{2}+r;iy\right)\vartheta_{3}\left(\frac{iv}{2}+r;iy\right)\right\}^{2}}{2u_{2}},\end{split}$
(J.4) $\begin{split}&1-\cos\lambda\\ &=-\frac{\left\{\vartheta_{1}\left(\frac{iv}{2}+r;iy\right)\vartheta_{3}\left(-\frac{iv}{2}+r;iy\right)-\vartheta_{1}\left(-\frac{iv}{2}+r;iy\right)\vartheta_{3}\left(\frac{iv}{2}+r;iy\right)\right\}^{2}}{2u_{2}}.\end{split}$
From (J.1) and (J.3), we see that (5.43) holds. Similarly, from (J.2) and (J.4), we see that (5.44) holds.
Appendix K Special property of elliptic theta functions when $w$ is a pure imaginary number
Lemma K.1.
The following holds for any $z\in\mathbb{C}$ when $w=iy$.
(K.1) $\vartheta_{j}(z)^{*}=\vartheta_{j}(z^{*}),\quad j=1,2,3,4.$
Proof.
Consider the following theta function with characteristics.
(K.2) $\vartheta_{a,b}(z):=\sum_{n\in\mathbb{Z}}\exp\left(2\pi i(n+a)(z+b)+\pi i(n+a)^{2}\omega\right),\quad a,b\in\left\{0,\frac{1}{2}\right\},$
then,
(K.3) $\vartheta_{1}(z)=-\vartheta_{\frac{1}{2},\frac{1}{2}}(z),\quad\vartheta_{2}(z)=\vartheta_{\frac{1}{2},0}(z),\quad\vartheta_{3}(z)=\vartheta_{0,0}(z),\quad\vartheta_{4}(z)=\vartheta_{0,\frac{1}{2}}(z)$
hold. Since $w^{*}=-w$, we have
(K.4) $\begin{split}\vartheta_{a,b}(z)^{*}=&\sum_{n\in\mathbb{Z}}\exp\left(-2\pi i(n+a)(z^{*}+b)-\pi i(n+a)^{2}w^{*}\right)\\ =&\sum_{n\in\mathbb{Z}}\exp\left(2\pi i(-n-a)(z^{*}+b)+\pi i(n+a)^{2}w\right)\\ =&\sum_{n\in\mathbb{Z}}\exp\left(2\pi i(n+a)(z^{*}+b)+\pi i(n+a)^{2}w\right)\\ =&\vartheta_{a,b}(z^{*}).\end{split}$
This lemma guarantees that some $\tau$ functions, such as $F_{n}$, are real-valued functions.