### v3

# An explicit construction of Kaleidocycles by elliptic theta functions

Abstract.
We consider the configuration space of ordered points on the two-dimensional sphere that satisfy a specific system of quadratic equations. We construct periodic orbits in this configuration space using elliptic theta functions and show that they simultaneously satisfy semi-discrete analogues of mKdV and sine-Gordon equations. The configuration space we investigate corresponds to the state space of a linkage mechanism known as the Kaleidocycle, and the constructed orbits describe the characteristic motion of the Kaleidocycle. A key consequence of our construction is the proof that Kaleidocycles exist for any number of tetrahedra greater than five. Our approach is founded on the relationship between the deformation of spatial curves and integrable systems, offering an intriguing example where an integrable system is explicitly solved to generate an orbit in the space of real solutions to polynomial equations defined by geometric constraints.
Key words and phrases:
Kinematic chain, Configuration space, Discrete differential geometry, Integrable systems, Elliptic theta function
2000 Mathematics Subject Classification:
Primary 53A04, 53A70; Secondary 53A17, 70B15, 37K25, 37K10, 35Q53
The first-named author was partially supported by KAKENHI, Grant-in-Aid for Scientific Research (B) 25K00921 and (S) 25H00399. The second and third-named authors were partially supported by KAKENHI Grant-in-Aid for Scientific Research (C) 21K3229.
1. Introduction
Given a natural number $k$ and a real number $1<\cos\lambda<1$, we consider the following space of configurations of $k+1$ ordered points on the 2-sphere $S^{2}$ satisfying certain quadratic constraints:
(1.1) $\mathcal{C}^{\pm}_{k,\lambda}=\{(b_{0},\ldots,b_{k})\in(S^{2})^{k+1}\mid$ $\langle b_{n-1},b_{n}\rangle=\cos\lambda,\,(1\leq n\leq k),$
$\sum_{n=0}^{k-1}(b_{n+1}\times b_{n})=0,$
$b_{k}=\pm b_{0}\}.$
Our main result states that for any $k\geq 6$, there exists $\lambda$ such that a circle can be embedded in $\mathcal{C}^{\pm}_{k,\lambda}$. Moreover, we explicitly construct a circle in terms of elliptic theta functions. This result implies the existence of a particular class of linkage mechanisms called Kaleidocycles (Fig. 1 (Left)). These mechanisms consist of $k$ identical copies of equifacial tetrahedra connected by their opposing edges as hinge joints, forming a ring. Although Kaleidocycles have been studied by several authors [5, 9, 30, 31], the rigorous proof of their existence beyond trivial ones has remained an open problem. Since an element of $\mathcal{C}^{\pm}_{k,\lambda}$ corresponds to a Kaleidocycle with $k$-tetrahedra in a specific position (Section 2.1) and its motion is described by a path in the configuration space, our main result provides an explicit construction and its periodic motion for Kaleidocycles of any $k\geq 6$.
The key observation is to identify a Kaleidocycle with a closed spatial polygonal curve formed by connecting the midpoints of the hinge edges. Then, the angles between the hinge edges are reminiscent of the torsion angle of a curve, and the turning angles are reminiscent of the signed curvature angle (§2.1). An element $(b_{0},\ldots,b_{k})\in\mathcal{C}^{\pm}_{k,\lambda}$ corresponds to a sequence of binormals. In this geometric formulation using spatial curves, the motion of a Kaleidocycle is modelled by a segment-length and torsion-angle preserving deformation of the corresponding closed polygonal curve of a constant torsion angle $\lambda$. In our construction, it plays the central role to compare the functional identities of elliptic theta functions with the compatibility conditions of the framing of a discrete spatial curve.
It was known in [21] that the semi-discrete modified KdV (mKdV) and the sine-Gordon equations are defined on the configuration space (see §3.2). The main result of this paper can also be viewed as providing a solution to the system through an explicit construction based on the $\tau$ functions of the two-component KP hierarchy [12]5). Moreover, the sweeping surface of the motion corresponds to a semi-discrete analogue of the discrete constant negative curvature surface (K-surface) studied in [2, 18].
Both views combined, a flow generated by an integrable system on a real-algebraic set $\mathcal{C}^{\pm}_{k,\lambda}$ is explicitly integrated using elliptic theta functions. Our work connects seemingly distant fields of integrable systems and kinematics of linkage via an interesting object, the Kaleidocycle.
The structure of the paper is as follows. Section 2 introduces the definition of Kaleidocycles. They are identified with a certain class of discrete spatial curves. As essential invariants to describe the state of a Kaleidocycle, the signed curvature angle and torsion angle are defined together with the Frenet frame. Section 3 reviews the connection between the theory of curves and integrable systems. The construction of discrete spatial curves and their Frenet frames in terms of the $\tau$ functions is presented. A previous result that describes the motion of a Kaleidocycle using semi-discrete integrable equations is recalled. Section 4 constructs a set of $\tau$ functions with three parameters $v,r,y$ using elliptic theta functions. The resulting curve is shown to have a constant torsion angle. In Section 5, a family of time-parametrised curves is constructed by introducing an additional parameter $t$. We then study the time evolution equations satisfied by the vertex positions and curvature angle of the curve and show connections to previous research. Section 6 investigates a necessary and sufficient condition for the resulting curve to be closed for every $t$. We prove there exists a set of parameters that satisfies the closure condition for a given $k\geq 6$, which proves the existence of Kaleidocycles. Section 7 provides examples of Kaleidocycles and corresponding semi-discrete surfaces with figures. We present a conjecture (Conjecture 7.1) suggested by numerical experiments: the Kaleidocycles constructed in this paper seem to be so-called Möbius Kaleidocycles.
2. Algebraic and geometric models of Kaleidocycle
2.1. Configuration space of a Kaleidocycle
A Kaleidocycle consists of (typically six) identical copies of equifacial tetrahedra connected by their opposing edges as hinge joints to form a ring. It can be folded from a sheet of paper and is an example of rigid origami; it rotates like a bubble ring while keeping all faces rigid. The origin of Kaleidocycle is obscure, but one of the first appearances in the literature is in Bricard’s paper [3], where he considers an essentially equivalent closed kinematic chain, now known as the Bricard 6R. For nearly a hundred years, the geometry and kinematics of this object have attracted the attention of researchers (see [30] and the references therein). However, there is no literature that discusses the existence of such objects for arbitrary $k$.
We first define a generalisation of the Kaleidocycle following [20]. We consider $k$-copies of an equifacial tetrahedron in $\mathbb{R}^{3}$ connected by opposing edges to form a ring. Denote the vertex set by $\{0,1,\ldots,2k-1\}$ such that the vertices $(2n,2n+1)$ form a hinge edge and the vertices $(2n,2n+1,2n+2)$ form one of the triangle faces. Let $l_{01},l_{12},l_{02}$ be the lengths of the sides of the triangle face, with the edge of length $l_{01}$ serving as a hinge (see Figure 1 Right). Now, a natural way to model a Kaleidocycle is as a realisation of an edge-weighted graph. Let $(V,E)$ be the graph formed by the vertices and edges of the tetrahedra in a Kaleidocycle and $\ell:E\to\mathbb{R}$ be the edge length. A realisation of $(V,E,\ell)$ in $\mathbb{R}^{3}$ is a map $\phi:V\to\mathbb{R}^{3}$ satisfying $|\phi(u)-\phi(v)|^{2}=\ell(uv)^{2}$ for all edges $uv\in E$.
In general, the collection of all possible realisations of an edge-weighted finite graph $(V,E,\ell)$ in $\mathbb{R}^{d}$ forms a real algebraic set in $\mathbb{R}^{d|V|}$, known as the configuration space of $(V,E,\ell)$. This space has been the subject of extensive research across various domains, including algebraic geometry, robotics, and mechanism theory [11, 23, 27].
More concretely, a state of a $k$-Kaleidocycle of type $(l_{01},l_{12},l_{02})$ is identified by the real solution $(x_{0},x_{1},\ldots,x_{2k+1})\in(\mathbb{R}^{3})^{2k+2}$ of the quadratic equations
(2.1) $|x_{2n}-x_{2n+1}|^{2}$ $=l^{2}_{01},$
$|x_{2n}-x_{2n+2}|^{2}$ $=|x_{2n+1}-x_{2n+3}|^{2}=l^{2}_{02},$
$|x_{2n+1}-x_{2n+2}|^{2}$ $=|x_{2n}-x_{2n+3}|^{2}=l^{2}_{12},$
for $0\leq n\leq k-2$ and $(x_{2k},x_{2k+1})=(x_{0},x_{1})$ or $(x_{1},x_{0})$. To remove the translational freedom, we demand $x_{0}+x_{1}=0$.
Notice that we have two combinatorially distinct cases: length $l_{02}$ edges form a $k$-cycle ($x_{2k}=x_{0}$) or $2k$-cycle, going around the ring twice ($x_{2k}=x_{1}$).
Remark 2.1.
The solution space of (2.1) admits an action of the orthogonal group $O(3)$ of $\mathbb{R}^{3}$, and the expected dimension of the quotient is $3\cdot 2k-5k-3-\dim(O(3))=k-6$, where $5k$ is the number of edges and $3$ comes from the equation $x_{0}+x_{1}=0$. This simple counting estimate is often called Maxwell’s degree-of-freedom counting law [9]. However, this estimate is sometimes violated in interesting examples. The Bricard 6R is a case of $k=6$. Its configuration space has dimension one, which is greater than the expected dimension of 0, giving an example of an overconstrained system  [9]. When $k\geq 7$, it is numerically observed that the configuration space degenerates to a circle when the angle between adjacent hinge edges is set to a critical value [22]. [^1] Some of the findings of [22] are also published in [31]. For $k>7$, this class of Kaleidocycles, named the Möbius Kaleidocycles, provides a rare example of an underconstrained linkage mechanism whose configuration space has a smaller dimension than the expected one.
We use the following elementary geometric property of equifacial tetrahedron:
Lemma 2.2.
An equifacial tetrahedron whose faces are congruent triangles with edge-lengths $(l_{01},l_{12},l_{02})$ may be realised within a rectangular cuboid of side lengths
$\sqrt{\frac{l_{01}^{2}+l_{12}^{2}-l_{02}^{2}}{2}},\quad\sqrt{\frac{l_{12}^{2}+l_{02}^{2}-l_{01}^{2}}{2}},\quad\sqrt{\frac{l_{02}^{2}+l_{01}^{2}-l_{12}^{2}}{2}}$
so that tetrahedron edges are diagonals of the faces. Moreover, the twist angle between the pair of opposite edges of length $l_{01}$ is given by the relation
$\cos(\lambda)=\frac{|l_{12}^{2}-l_{02}^{2}|}{l_{01}^{2}}.$
We will see the solution space of Equation 2.1 is identified with $\mathcal{C}^{\pm}_{k,\lambda}$.
Proposition 2.3.
The space of real solutions to Equation 2.1 is $O(3)$-equivariantly homeomorphic to $\mathcal{C}^{\pm}_{k,\lambda}$ with $\lambda=\arccos\left(\frac{|l_{12}^{2}-l_{02}^{2}|}{l_{01}^{2}}\right)$.
Proof.
Given $(x_{0},\ldots,x_{2k+1})$ satisfying Equation 2.1, we define $b_{n}=\dfrac{x_{2n+1}-x_{2n}}{|x_{2n+1}-x_{2n}|}$. By 2.2, the inner product $\langle b_{n-1},b_{n}\rangle$ is constant $\cos\lambda$ for $1\leq n<k$.
Let $\gamma_{n}=(x_{2n}+x_{2n+1})/2\ (0\leq n<k)$ be the mid-point of the hinge edge. There exists $l>0$ such that $l|b_{n+1}\times b_{n}|=|\gamma_{n+1}-\gamma_{n}|$. By 2.2, $\gamma_{n+1}-\gamma_{n}$ is orthogonal to both $b_{n}$ and $b_{n+1}$. Therefore, $b_{n+1}\times b_{n}$ is parallel to $\gamma_{n+1}-\gamma_{n}$, and we have $l\sum_{n=0}^{k-1}(b_{n+1}\times b_{n})=\sum_{n=0}^{k-1}(\gamma_{n+1}-\gamma_{n})=0$. This means $(b_{0},\ldots,b_{k})\in\mathcal{C}^{\pm}_{k,\lambda}$.
Conversely, given $(b_{0},\ldots,b_{k})\in\mathcal{C}^{\pm}_{k,\lambda}$, define $\gamma_{n}=l\sum_{m=0}^{n-1}(b_{m+1}\times b_{m})$, where $l$ is the distance of the opposing hinge edges of the tetrahedron. We set $x_{2n}=\gamma_{n}-\frac{l_{01}}{2}b_{n}$ and $x_{2n+1}=\gamma_{n}+\frac{l_{01}}{2}b_{n}$. Then, $(x_{0},\ldots,x_{2k+1})$ satisfies Equation 2.1.
It is clear that the above constructions are mutually inverse. ∎
While the shape of the tetrahedron is determined by the triangle edge lengths $(l_{01},l_{12},l_{02})$, the configuration space depends only on the hinge angle $\lambda$. We call $\lambda$ the torsion angle.
We call the (possible non-flat) polygon, or closed discrete spatial curve, defined by $\gamma_{n}$ the central curve of the Kaleidocycle.
Depending on the sign of $b_{k}=\pm b_{0}$, we say the Kaleidocycle is oriented (resp. non-oriented).
Figure 1. (Left) The classical 6-Kaleidocycle. (Centre) Kaleidocycle as a discrete curve with its framing (Right) Net of a Kaleidocycle with labelled vertices and edge lengths.
2.2. Discrete spatial curve of a constant torsion angle
We saw that a discrete spatial curve $\gamma_{n}$ is associated with a Kaleidocycle. We review some basics on discrete curves.
A discrete spatial curve with non-vanishing torsion angles can be described in terms of binormals. Let $\widetilde{B}:\mathbb{Z}\rightarrow S^{2}$ satisfy $\left|\widetilde{B}_{n}\times\widetilde{B}_{n-1}\right|\neq 0$ for all $n\in\mathbb{Z}$. A discrete spatial curve of unit segment-length having $\widetilde{B}$ as the binormals can be constructed as follows. Define the tangent $T:\mathbb{Z}\rightarrow S^{2}$ and the normal $\widetilde{N}:\mathbb{Z}\rightarrow S^{2}$ by
(2.2) $T_{n}=\frac{\widetilde{B}_{n+1}\times\widetilde{B}_{n}}{\left|\widetilde{B}_{n+1}\times\widetilde{B}_{n}\right|},\quad\widetilde{N}_{n}=\widetilde{B}_{n}\times T_{n}.$
Furthermore, given $\gamma_{0}\in\mathbb{R}^{3}$, define $\gamma:\mathbb{Z}\to\mathbb{R}^{3}$ by
(2.3) $\gamma_{n+1}=\gamma_{n}+T_{n}=\gamma_{n}+\frac{\widetilde{B}_{n+1}\times\widetilde{B}_{n}}{\left|\widetilde{B}_{n+1}\times\widetilde{B}_{n}\right|}.$
The signed curvature angle $\kappa$ and the torsion angle are defined by
(2.4) $\begin{split}\langle T_{n},T_{n-1}\rangle=\cos\kappa_{n},\quad\langle\widetilde{N}_{n},T_{n-1}\rangle&=-\sin\kappa_{n},\quad-\pi<\kappa_{n}<\pi,\\ \langle\widetilde{B}_{n},\widetilde{B}_{n-1}\rangle&=\cos\lambda_{n},\quad 0<\lambda_{n}<\pi.\end{split}$
The discrete Frenet frame $\Phi_{n}=[T_{n},\widetilde{N}_{n},\widetilde{B}_{n}]\in\mathrm{SO}(3)$ satisfies the discrete Frenet-Serret formula
(2.5) $\Phi_{n+1}=\Phi_{n}L_{n},\quad L_{n}=M_{1}(-\lambda_{n+1})M_{3}(\kappa_{n+1}),$
where
(2.6) $M_{1}(\theta)=\left[\begin{array}[]{ccc}1&0&0\\ 0&\cos\theta&-\sin\theta\\ 0&\sin\theta&\cos\theta\\ \end{array}\right],\quad M_{3}(\theta)=\left[\begin{array}[]{ccc}\cos\theta&-\sin\theta&0\\ \sin\theta&\cos\theta&0\\ 0&0&1\end{array}\right].$
Remark 2.4.
Note that our Frenet frame coincides with the usual one only up to sign. We adopt this convention to avoid discontinuous changes in the framing during the motion of the Kaleidocycle; its motion passes through the state where the curvature vanishes.
As in the case of a smooth curve, a discrete spatial curve of a unit segment length is determined by its torsion and curvature angles. Therefore, the motion of a discrete curve of a constant torsion angle is determined by the set of functions $\kappa_{n}(t)$, which can be freely chosen.
In this language, a Kaleidocycle is equivalent to a discrete closed ($\gamma_{k}=\gamma_{0})$ spatial curve with constant torsion angle $\lambda$ determined by the binormal $\widetilde{B}_{n}=b_{n}$. Our approach is to utilise this identification.
The fact that a Kaleidocycle is closed enforces interesting topological constraints. For example, the self-linking number must be a half-integer, and hence, stay constant during the motion. By the Călugăreanu-Fuller-White theorem [1], the writhing number of the curve must remain constant as well. The self-linking number also tells that the configuration space of a Kaleidocycle is not generally connected.
3. Curves and integrable systems
There are deep connections between differential geometry and the theory of integrable systems, where the integrable differential or difference equations arise as compatibility conditions of the geometric objects (see, for example, [29]). In the context of the deformation of spatial and plane curves, the Frenet frame and its deformation are described by the system of linear partial differential equations. The compatibility condition yields the so-called Ablowitz-Kaup-Newell-Segur (AKNS) hierarchy for the curvature and the torsion, including the nonlinear Schrödinger (NLS) equation as a typical example. In particular, if the torsion is a constant, the integrable deformation is governed by the mKdV equation and its hierarchy [10, 25, 29]. The connections are also investigated in a discrete setting, where curves are discretised as piecewise linear (polygonal) curves. For example, continuous deformations of discrete plane and spatial curves have been considered in [6], where the deformations are governed by the semi-discrete analogue of the mKdV and the NLS equations. Discrete deformations of the discrete plane and spatial curves have been studied in [18]. In principle, for a given solution to the nonlinear differential/difference equation for the curvature and torsion, one can reconstruct the curve by solving the associated system of linear differential/difference equations for the Frenet frame and integrating the tangent vector.
The theta functions naturally appear in so-called quasi-periodic solutions to the nonlinear integrable equations, and the corresponding solutions to the associated linear problem are nothing but the Baker-Akhiezer functions. Reconstruction of the deformation of the spatial curve with constant torsion described by the theta function solutions to the mKdV equation has been studied in this framework in [4, 26]. Also, an explicit formula of the spatial curves in terms of the $\tau$ functions of the two-component KP hierarchy based on Hirota’s bilinear formalism has been presented in [12].
3.1. Expression of a curve by the $\tau$ functions
We introduce a special class of discrete curves $\gamma_{n}$ and their Frenet frames $\Phi_{n}=[T_{n},\widetilde{N}_{n},\widetilde{B}_{n}]$ indexed by $n\in\mathbb{Z}$ using a modification of the $\tau$ functions presented in [12]. Let $F_{n},\widetilde{G}_{n}:\mathbb{C}\to\mathbb{R}$ and $f_{n},g_{n},G_{n},H_{n}:\mathbb{C}\to\mathbb{C}$ be functions indexed by $n\in\mathbb{Z}$ with a complex variable $z$ that satisfy the following equations
(3.1) $f_{n}f_{n}^{*}+g_{n}g_{n}^{*}=F_{n+1}F_{n},$
(3.2) $\beta_{1}F_{n}f_{n}+G_{n}g_{n}^{*}=F_{n+1}f_{n-1},$
(3.3) $\beta_{1}F_{n}g_{n}-G_{n}f_{n}^{*}=F_{n+1}g_{n-1},$
(3.4) $H_{n+1}F_{n}-H_{n}F_{n+1}=f_{n}^{*}g_{n},$
(3.5) $D_{z}F_{n+1}\cdot F_{n}:=\frac{dF_{n+1}}{dz}F_{n}-F_{n+1}\frac{dF_{n}}{dz}=g_{n}g_{n}^{*},$
(3.6) $\left|G_{n}\right|^{2}=\widetilde{G}_{n}^{2},$
where $\beta_{1}\in\mathbb{R}$ is some constant and $D_{z}$ is Hirota’s bilinear differential operator [13], and the symbol * denotes the complex conjugate. These functions are called the $\tau$ functions. Let $\varrho_{n}=\dfrac{1}{2}\dfrac{G_{n}}{\widetilde{G}_{n}}$ and
(3.10) $T_{n}$ $=\left.\frac{1}{f_{n}f_{n}^{*}+g_{n}g_{n}^{*}}\left(\begin{array}[]{c}\vskip 5.69054ptf_{n}^{*}g_{n}+f_{n}g_{n}^{*}\\ \vskip 5.69054pt\frac{1}{i}\left(f_{n}^{*}g_{n}-f_{n}g_{n}^{*}\right)\\ f_{n}f_{n}^{*}-g_{n}g_{n}^{*}\end{array}\right)\right|_{z=0},$
(3.14) $\widetilde{N}_{n}$ $=\left.\frac{1}{f_{n}f_{n}^{*}+g_{n}g_{n}^{*}}\left(\begin{array}[]{c}\vskip 5.69054pt\left\{(f_{n}^{*})^{2}-(g_{n}^{*})^{2}\right\}\varrho_{n}+\left\{(f_{n})^{2}-(g_{n})^{2}\right\}\varrho_{n}^{*}\\ \vskip 5.69054pt\frac{1}{i}\left[\left\{(f_{n}^{*})^{2}+(g_{n}^{*})^{2}\right\}\varrho_{n}-\left\{(f_{n})^{2}+(g_{n})^{2}\right\}\varrho_{n}^{*}\right]\\ -2\left(f_{n}^{*}g_{n}^{*}\varrho_{n}+f_{n}g_{n}\varrho_{n}^{*}\right)\end{array}\right)\right|_{z=0},$
(3.18) $\widetilde{B}_{n}$ $=\left.\frac{1}{f_{n}f_{n}^{*}+g_{n}g_{n}^{*}}\left(\begin{array}[]{c}\vskip 5.69054pt\frac{1}{i}\left[-\left\{(f_{n}^{*})^{2}-(g_{n}^{*})^{2}\right\}\varrho_{n}+\left\{(f_{n})^{2}-(g_{n})^{2}\right\}\varrho_{n}^{*}\right]\\ \vskip 5.69054pt\left\{(f_{n}^{*})^{2}+(g_{n}^{*})^{2}\right\}\varrho_{n}+\left\{(f_{n})^{2}+(g_{n})^{2}\right\}\varrho_{n}^{*}\\ \frac{2}{i}\left(f_{n}^{*}g_{n}^{*}\varrho_{n}-f_{n}g_{n}\varrho_{n}^{*}\right)\end{array}\right)\right|_{z=0},$
(3.22) $\gamma_{n}$ $=\left.\left(\begin{array}[]{c}\vskip 5.69054pt\dfrac{H_{n}+H_{n}^{*}}{F_{n}}\\ \vskip 5.69054pt\dfrac{1}{i}\dfrac{H_{n}-H_{n}^{*}}{F_{n}}\\ n-2\dfrac{\partial\log F_{n}}{\partial z}\end{array}\right)\right|_{z=0}.$
Similarly to [12, §3], we obtain:
Proposition 3.1.
$\Phi_{n}=[T_{n},\widetilde{N}_{n},\widetilde{B}_{n}]$ defines the discrete Frenet frame for the curve $\gamma$ satisfying the discrete Frenet-Serret formula (2.5). Its curvature and torsion angles are given by
(3.23) $\lambda_{n}$ $=\left.\arccos\left(\frac{G_{n}^{*}G_{n-1}+G_{n}G_{n-1}^{*}}{2\widetilde{G}_{n}\widetilde{G}_{n-1}}\right)\right|_{z=0},$
$\kappa_{n}$ $=\left.2\arctan\left(\frac{\widetilde{G}_{n}}{\beta_{1}F_{n}}\right)\right|_{z=0},$
where we take the principal value for $\arccos$ in $0,\pi]$ and for $\arctan$ in $(-\pi/2,\pi/2)$.
3.2. Motion of a Kaleidocycle on the osculating plane
Let $\gamma$ be a Kaleidocycle, that is, a closed curve with a constant torsion angle and a unit segment length. A motion of a Kaleidocycle is identified with a time-parametrised family $\gamma:\mathbb{Z}\times\mathbb{R}_{\geq 0}\to\mathbb{R}^{3}$ such that $\gamma(t)$ is a Kaleidocycle for all $t\geq 0$. In [21], a particular motion of a Kaleidocycle satisfying the following conditions is considered:
(3.24) $\left\langle\dot{\gamma}_{n},\widetilde{B}_{n}\right\rangle=0,\quad\left\langle\dot{\gamma}_{n},\dot{\gamma}_{n}\right\rangle=\tilde{\rho}^{2},\quad|{\gamma}_{n+1}-{\gamma}_{n}|=1\quad(0\leq\forall n\leq k),$
where $\rho\neq 0$ is a constant and $\dot{\gamma}_{n}=\dfrac{d\gamma_{n}}{dt}$. These conditions imply that the velocity of each vertex is restricted to the osculating plane and has a constant magnitude. It is shown in [21] that the Frenet-Serret formula (2.5), together with (3.24), requires that there exists a function $\Theta_{n}$, called the potential function, with $\kappa_{n}=\dfrac{\Theta_{n+1}-\Theta_{n-1}}{2}$ that satisfies either the semi-discrete potential mKdV equation:
(3.25) $\frac{d}{dt}\left(\frac{\Theta_{n+1}+\Theta_{n}}{2}\right)=\left(1+\cos\lambda\right)\rho\sin\left(\frac{\Theta_{n+1}-\Theta_{n}}{2}\right),$
or the semi-discrete sine-Gordon equation:
(3.26) $\frac{d}{dt}\left(\frac{\Theta_{n+1}-\Theta_{n}}{2}\right)=\left(1-\cos\lambda\right)\rho\sin\left(\frac{\Theta_{n+1}+\Theta_{n}}{2}\right).$
The motions corresponding to these equations differ in general. However, we will construct a motion in §5 that satisfies both equations simultaneously.
Remark 3.2.
As noted in §2.1, the dimension of the configurations of a Kaleidocycle defined by (1.1) is generally k6k-6. However, for the Möbius Kaleidocycles, which exhibit critical torsion angles, the dimension of the configurations is conjectured to be one, regardless of $k$. An extensive numerical simulation suggests that the motion of the Möbius Kaleidocycle satisfies (3.24). Therefore, (3.24) specifies a particular yet essential motion of Kaleidocycles.
4. An explicit formula of a discrete curve with a constant torsion angle in terms of elliptic theta functions
Based on the representation of curves discussed in §3.1, we define certain $\tau$ functions satisfying (3.1)-(3.6) using elliptic theta functions to construct a family of discrete curves with a constant torsion angle in Theorem 4.2.
4.1. A family of curves with a constant torsion angle
Throughout this section, $j$ runs through $\{1,2,3,4\}$. Fix $w\in\mathbb{C}$ with $\mathrm{Im}(w)>0$. The elliptic theta functions $\vartheta_{j}:\mathbb{C}\to\mathbb{C}$, are defined as follows [28, 24]:
(4.1) $\vartheta_{1}(u;w)$ $=-\sum_{n=-\infty}^{\infty}e^{\pi iw\left(n+\frac{1}{2}\right)^{2}+2\pi i\left(n+\frac{1}{2}\right)\left(u+\frac{1}{2}\right)},$
$\vartheta_{2}(u;w)$ $=\sum_{n=-\infty}^{\infty}e^{\pi iw\left(n+\frac{1}{2}\right)^{2}+2\pi i\left(n+\frac{1}{2}\right)u},$
$\vartheta_{3}(u;w)$ $=\sum_{n=-\infty}^{\infty}e^{\pi iwn^{2}+2\pi inu},$
$\vartheta_{4}(u;w)$ $=\sum_{n=-\infty}^{\infty}e^{\pi iwn^{2}+2\pi in\left(u+\frac{1}{2}\right)}.$
We often denote $\vartheta_{j}(u;w)$ by $\vartheta_{j}(u)$ when $w$ is fixed. These functions satisfy various identities, collected in Appendix A, that are convenient to construct the $\tau$ functions satisfying (3.1)–(3.6). We see the torsion angle given by Proposition 3.1 is a constant.
The following property is useful for examining whether a $\tau$ function constructed in this paper is real-valued.
Lemma 4.1.
The following holds for any $u\in\mathbb{C}$ when $w=iy$ $(y>0)$.
(4.2) $\vartheta_{j}(u)^{*}=\vartheta_{j}(u^{*}).$
Proof.
Consider the following theta function with characteristics.
(4.3) $\vartheta_{a,b}(u):=\sum_{n\in\mathbb{Z}}\exp\left(2\pi i(n+a)(u+b)+\pi i(n+a)^{2}w\right),\quad a,b\in\left\{0,\frac{1}{2}\right\},$
then,
(4.4) $\vartheta_{1}(u)=-\vartheta_{\frac{1}{2},\frac{1}{2}}(u),\quad\vartheta_{2}(u)=\vartheta_{\frac{1}{2},0}(u),\quad\vartheta_{3}(u)=\vartheta_{0,0}(u),\quad\vartheta_{4}(u)=\vartheta_{0,\frac{1}{2}}(u),$
hold. Since $w^{*}=-w$, we have
(4.5) $\begin{split}\vartheta_{a,b}(u)^{*}=&\sum_{n\in\mathbb{Z}}\exp\left(-2\pi i(n+a)(u^{*}+b)-\pi i(n+a)^{2}w^{*}\right)\\ =&\sum_{n\in\mathbb{Z}}\exp\left(2\pi i(-n-a)(u^{*}+b)+\pi i(n+a)^{2}w\right)\\ =&\sum_{n\in\mathbb{Z}}\exp\left(2\pi i(n+a)(u^{*}+b)+\pi i(n+a)^{2}w\right)\\ =&\vartheta_{a,b}(u^{*}).\end{split}$
For $u\in\mathbb{C}$ and $y>0$, we define $d_{j}(u;y)=\frac{\vartheta^{{}^{\prime}}_{j}(u;iy)}{\vartheta_{j}(u;iy)}$, where $^{{}^{\prime}}$ denotes the derivative with respect to the first variable of elliptic theta functions. For $y>0$, $r\in\mathbb{R}\setminus{(1/2\mathbb{Z})}$ and $v\in\mathbb{R}\setminus{(y\mathbb{Z})}$, we define
(4.6) $u_{1}=\vartheta_{1}\left(-\frac{1}{2}iv+r\right)\vartheta_{3}\left(-\frac{1}{2}iv+r\right)\vartheta_{1}\left(\frac{1}{2}iv+r\right)\vartheta_{3}\left(\frac{1}{2}iv+r\right),$
$\Delta_{j}=d_{j}\left(-\frac{1}{2}iv+r;y\right)-d_{j}\left(\frac{1}{2}iv+r;y\right),\quad R_{j}=\frac{\vartheta_{j}\left(-\frac{1}{2}iv+r\right)}{\vartheta_{j}\left(\frac{1}{2}iv+r\right)},$
where $\vartheta_{j}(u;iy)$ is denoted by $\vartheta_{j}(u)$. Using Lemma4.1, we see
(4.7) $\begin{split}u_{1}>0,\quad\Delta_{j}^{*}=-\Delta_{j},\quad R_{j}^{*}=R_{j}^{-1}.\end{split}$
Note that from (4.7), we see $R_{1}R_{3}^{*}=R_{1}R_{3}^{-1}$ and is a complex number of magnitude $1$. Thus
(4.8) $\begin{split}\frac{R_{1}R_{3}^{*}+(R_{1}R_{3}^{*})^{*}}{2}=\frac{R_{1}R_{3}^{-1}+R_{1}^{-1}R_{3}}{2},\end{split}$
is real and lies in [1,1][-1,1]. Also, from (4.7) and
(4.9) $\Delta_{3}-\Delta_{1}=-\frac{\vartheta_{3}\left(2r\right)\vartheta_{3}\left(0\right)\vartheta_{1}\left(iv\right)\vartheta_{1}^{{}^{\prime}}\left(0\right)}{u_{1}},$
we see that $\Delta_{3}-\Delta_{1}\neq 0$.
Theorem 4.2.
Put $\mu_{n}=ivn+\frac{z}{\Delta_{3}-\Delta_{1}}$ and define
(4.10) $F_{n}(z)=\alpha_{1}\exp\left(\frac{n\Delta_{3}}{\Delta_{3}-\Delta_{1}}z\right)\vartheta_{2}\left(\mu_{n}-\frac{1}{2}iv\right),$
$f_{n}(z)=\alpha_{2}R_{3}^{n}\exp\left(\frac{\left(n+\frac{1}{2}\right)\Delta_{3}}{\Delta_{3}-\Delta_{1}}z\right)\vartheta_{2}\left(\mu_{n}+r\right),$
$g_{n}(z)=\alpha_{3}R_{1}^{-n}\exp\left(\frac{\left(n+\frac{1}{2}\right)\Delta_{3}}{\Delta_{3}-\Delta_{1}}z\right)\vartheta_{4}\left(\mu_{n}-r\right),$
$G_{n}(z)=\alpha_{4}R_{1}^{-n}R_{3}^{n}\exp\left(\frac{n\Delta_{3}}{\Delta_{3}-\Delta_{1}}z\right)\vartheta_{4}\left(\mu_{n}-\frac{1}{2}iv\right),$
$H_{n}(z)=\alpha_{5}R_{1}^{-n}R_{3}^{-n}\exp\left(\frac{n\Delta_{3}}{\Delta_{3}-\Delta_{1}}z\right)\vartheta_{4}\left(\mu_{n}-\frac{1}{2}iv-2r\right),$
$\widetilde{G}_{n}(z)=-i\alpha_{4}\exp\left(\frac{n\Delta_{3}}{\Delta_{3}-\Delta_{1}}z\right)\vartheta_{4}\left(\mu_{n}-\frac{1}{2}iv\right),$
where
$\alpha_{1}=\sqrt{\vartheta_{3}(2r)\vartheta_{3}(0)},\quad\alpha_{2}=\vartheta_{3}\left(-\frac{1}{2}iv+r\right),\quad\alpha_{3}=\vartheta_{1}\left(\frac{1}{2}iv+r\right),$
$\alpha_{4}=\frac{\vartheta_{1}(iv)\sqrt{\vartheta_{3}(2r)\vartheta_{3}(0)}}{\vartheta_{3}(0)},\quad\alpha_{5}=\frac{u_{1}}{\vartheta_{3}(2r)\vartheta_{1}(iv)\sqrt{\vartheta_{3}(2r)\vartheta_{3}(0)}}.$
Then, they satisfy (3.1)-(3.6) with
(4.11) $\beta_{1}=\frac{\vartheta_{3}(iv)}{\vartheta_{3}(0)},$
and form $\tau$ functions. The corresponding discrete curve $\gamma$ defined by (3.10) has a unit segment length $\left|\gamma_{n+1}-\gamma_{n}\right|=1$. The torsion angle is given by
(4.12) $\begin{split}\lambda_{n}&=\arccos\left(\frac{G_{n}^{*}(0)G_{n-1}(0)+G_{n}(0)G_{n-1}(0)^{*}}{2\widetilde{G}_{n}(0)\widetilde{G}_{n-1}(0)}\right)\\ &=\arccos\left(\frac{R_{1}R_{3}^{-1}+R_{1}^{-1}R_{3}}{2}\right).\end{split}$
In particular, this does not depend on $n$, and $\gamma$ has a constant torsion angle $\lambda_{n}=\lambda_{0}$. Moreover, the signed curvature angle is given by
(4.13) $\begin{split}\kappa_{n}&=2\arctan\left(\frac{\widetilde{G}_{n}(0)}{\beta_{1}F_{n}(0)}\right)\\ &=2\arctan\left(-i\frac{\vartheta_{1}(iv)\vartheta_{4}\left(ivn-\frac{1}{2}iv\right)}{\vartheta_{3}(iv)\vartheta_{2}\left(ivn-\frac{1}{2}iv\right)}\right).\end{split}$
This theorem follows as a corollary of Theorem 5.1 and will be proven in §5.1.
Remark 4.3.
Note that $2\tan\frac{\kappa_{n}}{2}$ is often considered as the discritised curvature (see, for example, [14]), which is expressed using Jacobi $\mathrm{cn}$-function as follows:
[…]
where $\tilde{K}=\pi\vartheta_{3}\left(0;\frac{i}{y}\right)^{2}/2$ is the complete elliptic integral of the first kind and $\tilde{\xi}=\vartheta_{2}\left(0;\frac{i}{y}\right)^{2}/\vartheta_{3}\left(0;\frac{i}{y}\right)^{2}$ is the elliptic modulus.
5. Deformation of a discrete curve with a constant torsion angle
By introducing the deformation parameter $t$ to the curve constructed in §4.1, we construct an explicit formula for the motion of a Kaleidocycle in Theorem $5.1$). We then derive the evolution equations for the vertex positions of the curve and the (potential of the) curvature (§5.2, §5.3, §5.4). Connections to the previously-known isoperimetric deformation of a curve by the semi-discrete mKdV equation [18] and by the semi-discrete potential mKdV equation and the semi-discrete sine-Gordon equation [21] are discussed. This is also summarised in Table 1 in §6.
5.1. Explicit formula of the curve with deformation parameter
We introduce the deformation parameter $t$ so that only the signed curvature angles change while the torsion angle is kept constant. Additionally, the real parameters $C$ and $\Gamma$ are introduced to control the global motion of the curve. The following theorem describes a Kaleidocycle together with its motion explicitly in terms of elliptic theta functions.
Theorem 5.1.
Denote $\mu_{n}=ivn+\frac{z}{\Delta_{3}-\Delta_{1}}$. Consider the following functions parameterised by $t\in\mathbb{R}$:
(5.1) $F_{n}(t,z)=\alpha_{1}\exp\left(\frac{n\Delta_{3}}{\Delta_{3}-\Delta_{1}}z+\frac{C}{2}tz\right)\vartheta_{2}\left(\mu_{n}-\frac{1}{2}iv+it\right),$
$f_{n}(t,z)=\alpha_{2}R_{3}^{n}\exp\left(\frac{\left(n+\frac{1}{2}\right)\Delta_{3}}{\Delta_{3}-\Delta_{1}}z+\frac{C}{2}tz-\frac{\Gamma}{2}it\right)\vartheta_{2}\left(\mu_{n}+r+it\right),$
$g_{n}(t,z)=\alpha_{3}R_{1}^{-n}\exp\left(\frac{\left(n+\frac{1}{2}\right)\Delta_{3}}{\Delta_{3}-\Delta_{1}}z+\frac{C}{2}tz+\frac{\Gamma}{2}it\right)\vartheta_{4}\left(\mu_{n}-r+it\right),$
$G_{n}(t,z)=\alpha_{4}R_{1}^{-n}R_{3}^{n}\exp\left(\frac{n\Delta_{3}}{\Delta_{3}-\Delta_{1}}z+\frac{C}{2}tz\right)\vartheta_{4}\left(\mu_{n}-\frac{1}{2}iv+it\right),$
$H_{n}(t,z)=\alpha_{5}R_{1}^{-n}R_{3}^{-n}\exp\left(\frac{n\Delta_{3}}{\Delta_{3}-\Delta_{1}}z+\frac{C}{2}tz+\Gamma it\right)\vartheta_{4}\left(\mu_{n}-\frac{1}{2}iv-2r+it\right),$
$\widetilde{G}_{n}(t,z)=-i\alpha_{4}\exp\left(\frac{n\Delta_{3}}{\Delta_{3}-\Delta_{1}}z+\frac{C}{2}tz\right)\vartheta_{4}\left(\mu_{n}-\frac{1}{2}iv+it\right),$
where $C,\Gamma\in\mathbb{R}$ are constants. Then, they satisfy (3.1)-(3.6) with $\beta_{1}$ given by (4.11) and form $\tau$ functions. For each $t$, the corresponding discrete curve $\gamma$ (3.10) has a unit segment length $\left|\gamma_{n+1}-\gamma_{n}\right|=1$. The torsion angle is given by
(5.2) $\begin{split}\lambda_{n}&=\arccos\left(\frac{G_{n}^{*}(t,0)G_{n-1}(t,0)+G_{n}(t,0)G_{n-1}^{*}(t,0)}{2\widetilde{G}_{n}(t,0)\widetilde{G}_{n-1}(t,0)}\right)\\ &=\arccos\left(\frac{R_{1}R_{3}^{-1}+R_{1}^{-1}R_{3}}{2}\right).\end{split}$
In particular, this depends neither on $n$ nor on $t$, and $\gamma$ has a constant torsion angle $\lambda_{n}(t)=\lambda_{0}(0)$. Moreover, the signed curvature angle is given by
(5.3) $\begin{split}\kappa_{n}\left(t\right)&=\kappa_{n}\left(t;v,y\right)=2\arctan\left(\frac{\widetilde{G}_{n}(t,0)}{\beta_{1}F_{n}(t,0)}\right)\\ &=2\arctan\left(-i\frac{\vartheta_{1}(iv)\vartheta_{4}\left(ivn-\frac{1}{2}iv+it\right)}{\vartheta_{3}(iv)\vartheta_{2}\left(ivn-\frac{1}{2}iv+it\right)}\right).\end{split}$
Proof.
Putting
[…]
in (A.2) yields (3.1). Putting
[…]
in (A.2) yields (3.2). Putting
[…]
in (A.3) yields (3.3). Putting
[…]
in (A.4) yields
(5.24) $\begin{split}&\frac{\vartheta_{2}\left(0\right)\vartheta_{4}\left(0\right)}{\vartheta_{1}\left(-\frac{1}{2}iv+r\right)\vartheta_{3}\left(-\frac{1}{2}iv+r\right)}\vartheta_{2}\left(\mu_{n}-\frac{1}{2}iv+it\right)\vartheta_{4}\left(\mu_{n+1}-\frac{1}{2}iv-2r+it\right)\\ &=\frac{\vartheta_{2}\left(-\frac{1}{2}iv+r\right)\vartheta_{4}\left(-\frac{1}{2}iv+r\right)}{\vartheta_{1}\left(-\frac{1}{2}iv+r\right)\vartheta_{3}\left(-\frac{1}{2}iv+r\right)}\vartheta_{2}\left(\mu_{n}-r+it\right)\vartheta_{4}\left(\mu_{n}-r+it\right)\\ &-\vartheta_{3}\left(\mu_{n}-r+it\right)\vartheta_{1}\left(\mu_{n}-r+it\right).\end{split}$
Also, putting
[…]
in (A.4) yields
(5.30) $\begin{split}&\frac{\vartheta_{2}\left(0\right)\vartheta_{4}\left(0\right)}{\vartheta_{1}\left(\frac{1}{2}iv+r\right)\vartheta_{3}\left(\frac{1}{2}iv+r\right)}\vartheta_{2}\left(\mu_{n+1}-\frac{1}{2}iv+it\right)\vartheta_{4}\left(\mu_{n}-\frac{1}{2}iv-2r+it\right)\\ &=\frac{\vartheta_{2}\left(\frac{1}{2}iv+r\right)\vartheta_{4}\left(\frac{1}{2}iv+r\right)}{\vartheta_{1}\left(\frac{1}{2}iv+r\right)\vartheta_{3}\left(\frac{1}{2}iv+r\right)}\vartheta_{2}\left(\mu_{n}-r+it\right)\vartheta_{4}\left(\mu_{n}-r+it\right)\\ &-\vartheta_{3}\left(\mu_{n}-r+it\right)\vartheta_{1}\left(\mu_{n}-r+it\right).\end{split}$
Subtracting (5.30) from (5.24) and substituting
(5.31) $\begin{split}&\vartheta_{3}\left(2r\right)\vartheta_{4}\left(0\right)\vartheta_{2}\left(0\right)\vartheta_{1}\left(iv\right)\\ &=\vartheta_{1}\left(\frac{1}{2}iv+r\right)\vartheta_{3}\left(\frac{1}{2}iv+r\right)\vartheta_{2}\left(-\frac{1}{2}iv+r\right)\vartheta_{4}\left(-\frac{1}{2}iv+r\right)\\ &-\vartheta_{1}\left(-\frac{1}{2}iv+r\right)\vartheta_{3}\left(-\frac{1}{2}iv+r\right)\vartheta_{2}\left(\frac{1}{2}iv+r\right)\vartheta_{4}\left(\frac{1}{2}iv+r\right),\end{split}$
in the equation, we obtain
(5.32) $\begin{split}&\frac{\vartheta_{1}\left(-\frac{1}{2}iv+r\right)\vartheta_{3}\left(-\frac{1}{2}iv+r\right)\vartheta_{1}\left(\frac{1}{2}iv+r\right)\vartheta_{3}\left(\frac{1}{2}iv+r\right)}{\vartheta_{3}(2r)\vartheta_{1}(iv)}\\ &\times\left\{\frac{\vartheta_{1}\left(\frac{1}{2}iv+r\right)\vartheta_{3}\left(\frac{1}{2}iv+r\right)}{\vartheta_{1}\left(-\frac{1}{2}iv+r\right)\vartheta_{3}\left(-\frac{1}{2}iv+r\right)}\vartheta_{2}\left(\mu_{n}-\frac{1}{2}iv+it\right)\vartheta_{4}\left(\mu_{n+1}-\frac{1}{2}iv-2r+it\right)\right.\\ &\left.-\vartheta_{2}\left(\mu_{n+1}-\frac{1}{2}iv+it\right)\vartheta_{4}\left(\mu_{n}-\frac{1}{2}iv-2r+it\right)\right\}\\ &=\vartheta_{3}\left(\frac{1}{2}iv+r\right)\vartheta_{2}\left(\mu_{n}-r+it\right)\vartheta_{1}\left(\frac{1}{2}iv+r\right)\vartheta_{4}\left(\mu_{n}-r+it\right),\end{split}$
from which we see (3.4) holds. Putting
[…]
in (A.2) yields
(5.38) $\begin{split}&\vartheta_{2}\left(\mu_{n+1}-\frac{1}{2}iv+it+h\right)\vartheta_{2}\left(\mu_{n}-\frac{1}{2}iv+it-h\right)\vartheta_{3}(2r)\vartheta_{3}(0)\\ &=\vartheta_{2}\left(\mu_{n}+r+it\right)\vartheta_{2}\left(\mu_{n}-r+it\right)\vartheta_{3}\left(\frac{1}{2}iv+r+h\right)\vartheta_{3}\left(\frac{1}{2}iv-r+h\right)\\ &-\vartheta_{4}\left(\mu_{n}+r+it\right)\vartheta_{4}\left(\mu_{n}-r+it\right)\vartheta_{1}\left(\frac{1}{2}iv+r+h\right)\vartheta_{1}\left(\frac{1}{2}iv-r+h\right).\end{split}$
Differentiating (5.38) with $h$ and then putting $h=0$, we obtain
(5.39) $\begin{split}&\vartheta_{3}(2r)\vartheta_{3}(0)\left(\Delta_{3}-\Delta_{1}\right)D_{z}\vartheta_{2}\left(\mu_{n+1}-\frac{1}{2}iv+it\right)\cdot\vartheta_{2}\left(\mu_{n}-\frac{1}{2}iv+it\right)\\ &=-\Delta_{3}\vartheta_{3}\left(-\frac{1}{2}iv+r\right)\vartheta_{3}\left(\frac{1}{2}iv+r\right)\vartheta_{2}\left(\mu_{n}+r+it\right)\vartheta_{2}\left(\mu_{n}-r+it\right)\\ &-\Delta_{1}\vartheta_{1}\left(-\frac{1}{2}iv+r\right)\vartheta_{1}\left(\frac{1}{2}iv+r\right)\vartheta_{4}\left(\mu_{n}+r+it\right)\vartheta_{4}\left(\mu_{n}-r+it\right).\end{split}$
Using the equation (5.38) with $h=0$ and substituting (5.39), we obtain
(5.40) $\begin{split}&\vartheta_{3}(2r)\vartheta_{3}(0)\left\{\left(\Delta_{3}-\Delta_{1}\right)D_{z}\vartheta_{2}\left(\mu_{n+1}-\frac{1}{2}iv+it\right)\cdot\vartheta_{2}\left(\mu_{n}-\frac{1}{2}iv+it\right)\right.\\ &\left.+\Delta_{3}\vartheta_{2}\left(\mu_{n+1}-\frac{1}{2}iv+it\right)\vartheta_{2}\left(\mu_{n}-\frac{1}{2}iv+it\right)\right\}\\ &=\left(\Delta_{3}-\Delta_{1}\right)\vartheta_{1}\left(-\frac{1}{2}iv+r\right)\vartheta_{4}\left(\mu_{n}+r+it\right)\vartheta_{1}\left(\frac{1}{2}iv+r\right)\vartheta_{4}\left(\mu_{n}-r+it\right).\end{split}$
From this we see (3.5) holds. By the definition of $G$ and $\widetilde{G}$, (3.6) follows. Theorem 4.2 follows from Theorem 5.1 by setting $t=0$. ∎
Remark 5.2.
The parameter $C$ contributes only to the parallel translation of the curve along the $z$-axis, whereas $\Gamma$ contributes only to the rotation of the curve around the $z$-axis.
In Theorem 5.11, we will make convenient choices of $C$ and $\Gamma$ to derive the deformation equation considered in the previous studies presented in §3.2.
5.2. Motion of vertex positions
The construction in Theorem 5.1 involves parameters $C$ and $\Gamma$ that contribute only to the global rigid transformation of the curve. In general, the velocity of the vertex position $\dot{\gamma_{n}}$ has a binormal component, and its magnitude is not constant as described in Theorem 5.8. However, with the right choice of $C$ and $\Gamma$, curves that satisfy (3.24) are obtained as described in Theorem 5.11.
We first derive equations satisfied by the $\tau$ functions given by Theorem 5.1 to calculate $\dot{\gamma_{n}}$.
Lemma 5.3 ([12, §3]).
Using (3.2) and (3.3), we see that the following equations hold.
(5.41) $f_{n}f_{n-1}^{*}+g_{n}g_{n-1}^{*}=f_{n}^{*}f_{n-1}+g_{n}^{*}g_{n-1}=\beta_{1}F_{n}F_{n},$
(5.42) $g_{n}f_{n-1}-f_{n}g_{n-1}=G_{n}F_{n}.$
Proposition 5.4.
The $\tau$ functions (5.1) satisfy the following equations.
(5.43) $f_{n}f_{n-1}^{*}+g_{n}^{*}g_{n-1}=\beta_{1}F_{n}F_{n}+\beta_{2}G_{n}G_{n}^{*},$
(5.44) $g_{n}^{*}g_{n-1}-g_{n}g_{n-1}^{*}=f_{n}f_{n-1}^{*}-f_{n}^{*}f_{n-1}=\beta_{2}G_{n}G_{n}^{*},$
(5.45) $\beta_{4}f_{n}^{*}g_{n-1}-\beta_{4}^{*}f_{n-1}^{*}g_{n}=H_{n}F_{n},$
(5.46) $\beta_{3}F_{n}F_{n}+\beta_{4}G_{n}G_{n}^{*}=f_{n}f_{n-1}^{*},$
(5.47) $\left(\beta_{1}-\beta_{3}\right)F_{n}F_{n}+\left(\beta_{2}-\beta_{4}\right)G_{n}G_{n}^{*}=g_{n}^{*}g_{n-1},$
(5.48) $D_{t}H_{n}\cdot F_{n}=\left(\beta_{5}+i\Gamma\beta_{4}\right)f_{n}^{*}g_{n-1}+\left(\beta_{5}^{*}-i\Gamma\beta_{4}^{*}\right)f_{n-1}^{*}g_{n},$
(5.49) $D_{t}D_{z}F_{n}\cdot F_{n}=\left(\beta_{6}+C\right)F_{n}F_{n}+\beta_{7}G_{n}G_{n}^{*}.$
where
(5.50) $\beta_{2}=\frac{\vartheta_{1}\left(2r\right)\vartheta_{3}\left(0\right)}{\vartheta_{3}\left(2r\right)\vartheta_{1}\left(iv\right)},$
$\beta_{3}=\frac{\vartheta_{3}\left(\frac{1}{2}iv+r\right)^{2}\vartheta_{3}\left(-\frac{1}{2}iv+r\right)^{2}}{\vartheta_{3}\left(2r\right)\vartheta_{3}\left(0\right)^{3}},\quad\beta_{4}=\frac{\vartheta_{1}\left(\frac{1}{2}iv+r\right)^{2}\vartheta_{3}\left(-\frac{1}{2}iv+r\right)^{2}}{\vartheta_{1}\left(iv\right)^{2}\vartheta_{3}\left(2r\right)\vartheta_{3}\left(0\right)},$
$\beta_{5}=\frac{-iu_{1}}{\vartheta_{3}(0)\vartheta_{3}(2r)\vartheta_{1}(iv)^{2}}\left[\frac{\vartheta_{1}^{{}^{\prime}}\left(\frac{1}{2}iv+r\right)\vartheta_{3}\left(-\frac{1}{2}iv+r\right)}{\vartheta_{1}\left(-\frac{1}{2}iv+r\right)\vartheta_{3}\left(\frac{1}{2}iv+r\right)}+\frac{\vartheta_{1}\left(\frac{1}{2}iv+r\right)\vartheta_{3}^{{}^{\prime}}\left(-\frac{1}{2}iv+r\right)}{\vartheta_{1}\left(-\frac{1}{2}iv+r\right)\vartheta_{3}\left(\frac{1}{2}iv+r\right)}\right],$
$\beta_{6}=\frac{2i}{\Delta_{3}-\Delta_{1}}\frac{\vartheta_{3}^{{}^{\prime\prime}}(0)}{\vartheta_{3}(0)},\quad\beta_{7}=\frac{2i}{\Delta_{3}-\Delta_{1}}\frac{\vartheta_{1}^{{}^{\prime}}(0)^{2}}{\vartheta_{1}(iv)^{2}}.$
Proof.
Putting
[…]
in (A.8) yields (5.43). Also, (5.44) follows from (5.41) and (5.43). Putting
[…]
in (A.3) yields
(5.61) $\begin{split}&\vartheta_{4}\left(\mu_{n}-\frac{1}{2}iv-2r+it+h\right)\vartheta_{3}\left(0\right)\vartheta_{1}\left(iv\right)\vartheta_{2}\left(\mu_{n}-\frac{1}{2}iv+it-h\right)\\ &=\vartheta_{4}\left(\mu_{n}-r+it\right)\vartheta_{3}\left(-\frac{1}{2}iv-r+h\right)\vartheta_{1}\left(\frac{1}{2}iv-r+h\right)\vartheta_{2}\left(\mu_{n-1}-r+it\right)\\ &-\vartheta_{2}\left(\mu_{n}-r+it\right)\vartheta_{1}\left(-\frac{1}{2}iv-r+h\right)\vartheta_{3}\left(\frac{1}{2}iv-r+h\right)\vartheta_{4}\left(\mu_{n-1}-r+it\right),\end{split}$
After differentiating (5.61) by $h$ and putting $h=0$, we obtain (5.48) by rearrangement. Also, rearranging the equation with $h=0$ in (5.61) yields (5.45). Putting
[…]
in (A.5) yields (5.46). Also, (5.47) follows from (5.43) and (5.46). Putting
[…]
in (A.2) yields
(5.72) $\begin{split}&\vartheta_{2}\left(\mu_{n}-\frac{1}{2}iv+it+h\right)\vartheta_{2}\left(\mu_{n}-\frac{1}{2}iv+it-h\right)\\ &=-\vartheta_{4}\left(\mu_{n}-\frac{1}{2}iv+it\right)^{2}\frac{\vartheta_{1}\left(h\right)^{2}}{\vartheta_{3}\left(0\right)^{2}}+\vartheta_{2}\left(\mu_{n}-\frac{1}{2}iv+it\right)^{2}\frac{\vartheta_{3}\left(h\right)^{2}}{\vartheta_{3}\left(0\right)^{2}}.\end{split}$
After differentiating (5.72) twice with $h$ and putting $h=0$, we obtain (5.49). ∎
Coefficients $\beta_{2}$ and $\beta_{4}$ in Proposition 5.4 satisfy the following relation.
Lemma 5.5.
The following relation holds.
(5.73) $\beta_{2}=\beta_{4}-\beta_{4}^{*}.$
Proof.
Putting
[…]
in (A.9) yields (5.73). ∎
Furthermore, to simplify the calculation of $\dot{\gamma_{n}}$, we define $P$ and $Q$ as follows
(5.79) $P_{n}=g_{n}^{*}f_{n-1}-f_{n}g_{n-1}^{*},$
(5.80) $Q_{n}=g_{n}^{*}f_{n-1}+f_{n}g_{n-1}^{*}.$
Then the following expressions hold
(5.81) $D_{t}H_{n}\cdot F_{n}=\eta_{1}Q_{n}^{*}+i\eta_{2}P_{n}^{*},$
(5.82) $\tan\frac{\kappa_{n}}{2}\widetilde{N}_{n}=T_{n}-\frac{1}{2\beta_{1}F_{n}^{2}}\left(\begin{array}[]{c}\vskip 5.69054ptQ_{n}^{*}+Q_{n}\\ \vskip 5.69054pt\frac{1}{i}\left(Q_{n}^{*}-Q_{n}\right)\\ -2\left(\beta_{1}-2\beta_{3}\right)F_{n}^{2}+2\left(\beta_{4}+\beta_{4}^{*}\right)|G_{n}|^{2}\end{array}\right),$
(5.83) $\tan\frac{\kappa_{n}}{2}\widetilde{B}_{n}=\frac{1}{2\beta_{1}F_{n}^{2}}\left(\begin{array}[]{c}\vskip 5.69054pt\frac{1}{i}\left(P_{n}-P_{n}^{*}\right)\\ \vskip 5.69054ptP_{n}+P_{n}^{*}\\ -\frac{2}{i}\beta_{2}|G_{n}|^{2}\end{array}\right),$
where
(5.84) $\eta_{1}=\frac{1}{2}\left(\beta_{5}+\beta_{5}^{*}\right)+\frac{1}{2}i\Gamma\left(\beta_{4}-\beta_{4}^{*}\right),\quad\eta_{2}=\frac{1}{2}i\left(\beta_{5}-\beta_{5}^{*}\right)-\frac{1}{2}\Gamma\left(\beta_{4}+\beta_{4}^{*}\right).$
We have used (3.2) and (3.3) for (5.83). Then we have the following lemmas.
Lemma 5.6.
The following relations hold.
(5.85) $2\eta_{1}\left(\beta_{4}+\beta_{4}^{*}\right)+2i\beta_{2}\eta_{2}+\beta_{7}=2\beta_{4}\beta_{5}^{*}+2\beta_{4}^{*}\beta_{5}+\beta_{7}=0,$
(5.86) $4\eta_{1}^{2}\left(\beta_{4}+\beta_{4}^{*}\right)^{2}+8i\eta_{1}\eta_{2}\beta_{2}\left(\beta_{4}+\beta_{4}^{*}\right)-4\eta_{2}^{2}\beta_{2}^{2}-\beta_{7}^{2}=0.$
Proof.
By the definition of $\eta_{1}$ and $\eta_{2}$ and Lemma 5.5, we can see that the leftmost and middle part of (5.85) match. Putting
[…]
in (A.9) yields
(5.92) $\begin{split}&\vartheta_{3}\left(\frac{1}{2}iv+r+h\right)\vartheta_{3}\left(-\frac{1}{2}iv+r-h\right)\vartheta_{1}\left(\frac{1}{2}iv+r-h\right)\vartheta_{1}\left(-\frac{1}{2}iv+r+h\right)\\ &-\vartheta_{1}\left(\frac{1}{2}iv+r+h\right)\vartheta_{1}\left(-\frac{1}{2}iv+r-h\right)\vartheta_{3}\left(\frac{1}{2}iv+r-h\right)\vartheta_{3}\left(-\frac{1}{2}iv+r+h\right)\\ &=\vartheta_{3}\left(2r\right)\vartheta_{3}\left(0\right)\vartheta_{1}\left(iv\right)\vartheta_{1}\left(2h\right).\end{split}$
After differentiating (5.92) by $h$ and setting $h=0$, we obtain (5.85). From
(5.93) $\begin{split}&4\eta_{1}^{2}\left(\beta_{4}+\beta_{4}^{*}\right)^{2}+8i\eta_{1}\eta_{2}\beta_{2}\left(\beta_{4}+\beta_{4}^{*}\right)-4\eta_{2}^{2}\beta_{2}^{2}-\beta_{7}^{2}\\ =&\left\{2\eta_{1}\left(\beta_{4}+\beta_{4}^{*}\right)+2i\beta_{2}\eta_{2}+\beta_{7}\right\}\left\{2\eta_{1}\left(\beta_{4}+\beta_{4}^{*}\right)+2i\beta_{2}\eta_{2}-\beta_{7}\right\},\end{split}$
we obtain (5.86). ∎
Lemma 5.7.
The following relations hold.
(5.94) $P_{n}P_{n}^{*}=F_{n}^{2}|G_{n}|^{2}+\beta_{2}^{2}|G_{n}|^{4},$
(5.95) $\begin{split}P_{n}Q_{n}^{*}-P_{n}^{*}Q_{n}=2\left(2\beta_{3}-\beta_{1}\right)\beta_{2}F_{n}^{2}|G_{n}|^{2}+2\beta_{2}\left(\beta_{4}+\beta_{4}^{*}\right)|G_{n}|^{4},\end{split}$
(5.96) $\begin{split}&Q_{n}Q_{n}^{*}=4\beta_{3}\left(\beta_{1}-\beta_{3}\right)F_{n}^{4}+\left\{1+2\left(\beta_{1}-2\beta_{3}\right)\left(\beta_{4}+\beta_{4}^{*}\right)\right\}F_{n}^{2}|G_{n}|^{2}\\ &-\left(\beta_{4}+\beta_{4}^{*}\right)^{2}|G_{n}|^{4}.\end{split}$
Proof.
By definition of $P$, we see that
(5.97) $\begin{split}P_{n}P_{n}^{*}&=\left(g_{n}^{*}f_{n-1}-f_{n}g_{n-1}^{*}\right)\left(g_{n}f_{n-1}^{*}-f_{n}^{*}g_{n-1}\right)\\ &=\left(f_{n}^{*}g_{n-1}^{*}-g_{n}^{*}f_{n-1}^{*}\right)\left(f_{n}g_{n-1}-g_{n}f_{n-1}\right)\\ &+\left(f_{n}f_{n-1}^{*}-f_{n}^{*}f_{n-1}\right)\left(g_{n}^{*}g_{n-1}-g_{n}g_{n-1}^{*}\right),\end{split}$
holds. Substituting (5.42) and (5.44) into (5.97) yields (5.94). By definition of $P$ and $Q$, we see that
[…]
holds. On the other hand, from (5.41) and (5.44), the following holds.
[…]
Using (5.99), we obtain
(5.100) $\begin{split}&P_{n}Q_{n}^{*}-P_{n}^{*}Q_{n}\\ &=f_{n}f_{n-1}^{*}f_{n}f_{n-1}^{*}-f_{n}^{*}f_{n-1}f_{n}^{*}f_{n-1}+g_{n}g_{n-1}^{*}g_{n}g_{n-1}^{*}-g_{n}^{*}g_{n-1}g_{n}^{*}g_{n-1}\\ &=\left(f_{n}f_{n-1}^{*}\right)^{2}-\left(f_{n}^{*}f_{n-1}\right)^{2}+\left(g_{n}g_{n-1}^{*}\right)^{2}-\left(g_{n}^{*}g_{n-1}\right)^{2}\\ &=\left(f_{n}f_{n-1}^{*}-f_{n}^{*}f_{n-1}\right)\left(f_{n}f_{n-1}^{*}+f_{n}^{*}f_{n-1}\right)\\ &+\left(g_{n}g_{n-1}^{*}-g_{n}^{*}g_{n-1}\right)\left(g_{n}g_{n-1}^{*}+g_{n}^{*}g_{n-1}\right).\end{split}$
Substituting (5.44), (5.46) and (5.47) into (5.100) gives
(5.101) $\begin{split}&P_{n}Q_{n}^{*}-P_{n}^{*}Q_{n}\\ &=\beta_{2}|G_{n}|^{2}\left(f_{n}f_{n-1}^{*}+f_{n}^{*}f_{n-1}-g_{n}g_{n-1}^{*}-g_{n}^{*}g_{n-1}\right)\\ &=2\left(2\beta_{3}-\beta_{1}\right)\beta_{2}F_{n}^{2}|G_{n}|^{2}+2\beta_{2}\left(\beta_{4}+\beta_{4}^{*}\right)|G_{n}|^{4},\end{split}$
thus (5.95). Also from (5.42), (5.46) and (5.47), we see that
(5.102) $\begin{split}&Q_{n}Q_{n}^{*}-F_{n}^{2}|G_{n}|^{2}\\ &=\left(g_{n}^{*}f_{n-1}+f_{n}g_{n-1}^{*}\right)\left(g_{n}f_{n-1}^{*}+f_{n}^{*}g_{n-1}\right)\\ &-\left(g_{n}f_{n-1}-f_{n}g_{n-1}\right)\left(g_{n}^{*}f_{n-1}^{*}-f_{n}^{*}g_{n-1}^{*}\right)\\ &=\left(f_{n}f_{n-1}^{*}+f_{n}^{*}f_{n-1}\right)\left(g_{n}g_{n-1}^{*}+g_{n}^{*}g_{n-1}\right)\\ &=\left(2\beta_{3}F_{n}^{2}+\left(\beta_{4}+\beta_{4}^{*}\right)|G_{n}|^{2}\right)\left(2\left(\beta_{1}-\beta_{3}\right)F_{n}^{2}-\left(\beta_{4}+\beta_{4}^{*}\right)|G_{n}|^{2}\right)\\ &=4\beta_{3}\left(\beta_{1}-\beta_{3}\right)F_{n}^{4}+2\left(\beta_{1}-2\beta_{3}\right)\left(\beta_{4}+\beta_{4}^{*}\right)F_{n}^{2}|G_{n}|^{2}-\left(\beta_{4}+\beta_{4}^{*}\right)^{2}|G_{n}|^{4},\end{split}$
holds. Thus we get (5.96). ∎
By using Proposition 5.4, Lemma 5.5 Lemma 5.6 and Lemma 5.7, we obtain the evolution equation of the vertex positions of the curve.
Theorem 5.8.
The curve $\gamma$ constructed in Theorem 5.1 satisfies the following deformation equation.
(5.103) $\begin{split}\dot{\gamma}_{n}(t)=&2\beta_{1}\left[\eta_{1}T_{n}-\eta_{1}\tan\frac{\kappa_{n}}{2}\widetilde{N}_{n}+\eta_{2}\tan\frac{\kappa_{n}}{2}\widetilde{B}_{n}\right]\\ &+\left(\begin{array}[]{c}0\\ 0\\ 2\eta_{1}\left(\beta_{1}-2\beta_{3}\right)-\beta_{6}-C\end{array}\right).\end{split}$
Proof.
By differentiating $\gamma$ with respect to $t$, we get
(5.104) $\begin{split}\dot{\gamma}_{n}(t)&=\frac{1}{F_{n}^{2}}\left(\begin{array}[]{c}\vskip 5.69054pt\displaystyle D_{t}H_{n}\cdot F_{n}+D_{t}H_{n}^{*}\cdot F_{n}\\ \vskip 5.69054pt\displaystyle\frac{1}{i}\left(D_{t}H_{n}\cdot F_{n}-D_{t}H_{n}^{*}\cdot F_{n}\right)\\ -D_{t}D_{z}F_{n}\cdot F_{n}\end{array}\right)\\ &=\frac{1}{F_{n}^{2}}\left(\begin{array}[]{c}\vskip 5.69054pt\displaystyle\eta_{1}Q_{n}^{*}+i\eta_{2}P_{n}^{*}+\eta_{1}Q_{n}-i\eta_{2}P_{n}\\ \vskip 5.69054pt\displaystyle\frac{1}{i}\left(\eta_{1}Q_{n}^{*}+i\eta_{2}P_{n}^{*}-\eta_{1}Q_{n}+i\eta_{2}P_{n}\right)\\ -\left(\beta_{6}+C\right)F_{n}^{2}-\beta_{7}|G_{n}|^{2}\end{array}\right).\end{split}$
On the other hand, from (5.82) and (5.83), we see
(5.105) $\begin{split}&\eta_{1}\left(T_{n}-\tan\frac{\kappa_{n}}{2}\widetilde{N}_{n}\right)+\eta_{2}\tan\frac{\kappa_{n}}{2}\widetilde{B}_{n}\\ &=\frac{1}{2\beta_{1}F_{n}^{2}}\left(\begin{array}[]{c}\vskip 5.69054pt\displaystyle\eta_{1}Q_{n}^{*}+i\eta_{2}P_{n}^{*}+\eta_{1}Q_{n}-i\eta_{2}P_{n}\\ \vskip 5.69054pt\displaystyle\frac{1}{i}\left(\eta_{1}Q_{n}^{*}+i\eta_{2}P_{n}^{*}-\eta_{1}Q_{n}+i\eta_{2}P_{n}\right)\\ -2\eta_{1}\left(\beta_{1}-2\beta_{3}\right)F_{n}^{2}+2\eta_{1}\left(\beta_{4}+\beta_{4}^{*}\right)|G_{n}|^{2}+2i\eta_{2}\beta_{2}|G_{n}|^{2}\end{array}\right)\end{split}$
holds. From (5.85), (5.105) can be rewritten as follows
(5.106) $\begin{split}&\eta_{1}\left(T_{n}-\tan\frac{\kappa_{n}}{2}\widetilde{N}_{n}\right)+\eta_{2}\tan\frac{\kappa_{n}}{2}\widetilde{B}_{n}\\ &=\frac{1}{2\beta_{1}F_{n}^{2}}\left(\begin{array}[]{c}\vskip 5.69054pt\displaystyle\eta_{1}Q_{n}^{*}+i\eta_{2}P_{n}^{*}+\eta_{1}Q_{n}-i\eta_{2}P_{n}\\ \vskip 5.69054pt\displaystyle\frac{1}{i}\left(\eta_{1}Q_{n}^{*}+i\eta_{2}P_{n}^{*}-\eta_{1}Q_{n}+i\eta_{2}P_{n}\right)\\ -2\eta_{1}\left(\beta_{1}-2\beta_{3}\right)F_{n}^{2}-\beta_{7}|G_{n}|^{2}\end{array}\right).\end{split}$
Thus (5.103) is obtained. ∎
This is a deformation equation similar to the one presented in [18]. Starting from (5.103), the compatibility condition of the Frenet frame yields the semi-discrete mKdV equation.
Proposition 5.9.
When the curve with a unit segment length and a constant torsion angle deforms according to the deformation equation (5.103), the following holds.
  1. (1)
    The segment length is preserved.
  2. (2)
    $\kappa$ satisfies the following semi-discrete mKdV equation:
    (5.107) $\dot{\kappa}_{n}=2\beta_{1}\left(\eta_{1}\cos\lambda+\eta_{2}\sin\lambda\right)\left(\tan\frac{\kappa_{n+1}}{2}-\tan\frac{\kappa_{n-1}}{2}\right).$
Furthermore, if the left-hand-side of the equation (5.136) does not depend on $n$, the following holds:
  1. (3)
    The torsion angle is preserved.
Proof.
The proofs of (i) and (ii) follow from the same calculation as in Section 5 of [18]. We prove (iii). From the deformation equation (5.103) and the discrete Frenet-Serret formula (2.5), we see that
(5.108) $\begin{split}&-\dot{\lambda}\sin\lambda=\left\langle\dot{\widetilde{B}}_{n},\widetilde{B}_{n+1}\right\rangle+\left\langle\widetilde{B}_{n},\dot{\widetilde{B}}_{n+1}\right\rangle\\ &=-\sin\lambda\left(2\beta_{1}\eta_{1}\sin\lambda-2\beta_{1}\eta_{2}\cos\lambda\right)\times\left(\frac{1}{\sin\kappa_{n+1}}\tan\frac{\kappa_{n+2}}{2}\right.\\ &\left.+\frac{1}{\tan\kappa_{n+1}}\tan\frac{\kappa_{n}}{2}-\frac{1}{\tan\kappa_{n}}\tan\frac{\kappa_{n+1}}{2}-\frac{1}{\sin\kappa_{n}}\tan\frac{\kappa_{n-1}}{2}\right),\end{split}$
holds, which vanishes because of (5.136). Therefore $\dot{\lambda}=0$.
Remark 5.10.
Proposition 5.9 (iii) does not hold in general without the assumption on (5.136). Indeed, there is a motion of a Kaleidocycle satisfying (5.103) such that the left-hand side of (5.136) depends both on $n$ and $t$.
By adjusting the global motion of $\gamma$ by $C$ and $\Gamma$, the deformation equation of $\gamma$ can be made to satisfy the equation considered in the previous studies presented in §3.2.
Theorem 5.11.
Set $C$ and $\Gamma$ to be
(5.109) $\begin{split}\Gamma_{+}\left(r,y\right)=&\frac{\vartheta_{1}^{{}^{\prime}}(2r)\vartheta_{3}(0)+\vartheta_{3}\left(2r\right)\vartheta_{1}^{{}^{\prime}}(0)}{\vartheta_{1}\left(2r\right)\vartheta_{3}\left(0\right)},\\ C_{+}\left(v,r,y\right)=&-i\Gamma_{+}\left(\beta_{2}\left(2\beta_{3}-\beta_{1}\right)-\frac{\beta_{4}+\beta_{4}^{*}}{\beta_{2}}\right)-\beta_{6}-\left(\beta_{5}+\beta_{5}^{*}\right)\left(2\beta_{3}-\beta_{1}\right)+\frac{\beta_{5}-\beta_{5}^{*}}{\beta_{2}}=0,\\ \rm or\\ \Gamma_{-}\left(r,y\right)=&\frac{\vartheta_{1}^{{}^{\prime}}(2r)\vartheta_{3}(0)-\vartheta_{3}\left(2r\right)\vartheta_{1}^{{}^{\prime}}(0)}{\vartheta_{1}\left(2r\right)\vartheta_{3}\left(0\right)},\\ C_{-}\left(v,r,y\right)=&-i\Gamma_{-}\left(\beta_{2}\left(2\beta_{3}-\beta_{1}\right)-\frac{\beta_{4}+\beta_{4}^{*}}{\beta_{2}}\right)-\beta_{6}-\left(\beta_{5}+\beta_{5}^{*}\right)\left(2\beta_{3}-\beta_{1}\right)+\frac{\beta_{5}-\beta_{5}^{*}}{\beta_{2}}=0.\end{split}$
Then, the curve constructed in Theorem 5.1 deforms according to the equation (3.24). The magnitude of $\dot{\gamma}_{n}$ is a constant and is given by
(5.110) $\rho^{2}=\rho_{+}^{2}:=\left(i\frac{\vartheta_{1}^{{}^{\prime}}(0)}{\vartheta_{1}(iv)}\frac{u_{1}}{\vartheta_{3}(r)^{2}\vartheta_{1}(r)^{2}}\right)^{2},$
when $(C,\Gamma)=(C_{+},\Gamma_{+})$ and
(5.111) $\rho^{2}=\rho_{-}^{2}:=\left(-i\frac{\vartheta_{1}^{{}^{\prime}}(0)}{\vartheta_{1}(iv)}\frac{u_{1}}{\vartheta_{4}(r)^{2}\vartheta_{2}(r)^{2}}\right)^{2},$
when $(C,\Gamma)=(C_{-},\Gamma_{-})$.
Proof.
We determine the necessary and sufficient condition for $C$ and $\Gamma$ to satisfy
(5.112) $\left\langle\dot{\gamma}_{n},\widetilde{B}_{n}\right\rangle=0,$
(5.113) $\left\langle\dot{\gamma}_{n},\dot{\gamma}_{n}\right\rangle=\rho^{2},$
where $\rho\in\mathbb{R}$ is a constant. From (5.83) and (5.104) we see
(5.114) $\begin{split}2iF_{n}^{3}\widetilde{G}_{n}\left\langle\dot{\gamma}_{n},\widetilde{B}_{n}\right\rangle=&\left(P_{n}-P_{n}^{*}\right)\left(\eta_{1}Q_{n}^{*}+i\eta_{2}P_{n}^{*}+\eta_{1}Q_{n}-i\eta_{2}P_{n}\right)\\ &+\left(P_{n}+P_{n}^{*}\right)\left(\eta_{1}Q_{n}^{*}+i\eta_{2}P_{n}^{*}-\eta_{1}Q_{n}+i\eta_{2}P_{n}\right)\\ &+2\beta_{2}|G_{n}|^{2}\left\{\left(\beta_{6}+C\right)F_{n}^{2}+\beta_{7}|G_{n}|^{2}\right\}\\ =&2F_{n}^{2}|G_{n}|^{2}\left\{2\eta_{1}\left(2\beta_{3}-\beta_{1}\right)\beta_{2}+2i\eta_{2}+\beta_{2}\left(\beta_{6}+C\right)\right\}\\ &+2\beta_{2}|G_{n}|^{4}\left\{2\eta_{1}\left(\beta_{4}+\beta_{4}^{*}\right)+2i\beta_{2}\eta_{2}+\beta_{7}\right\}\\ =&2F_{n}^{2}|G_{n}|^{2}\left\{2\eta_{1}\left(2\beta_{3}-\beta_{1}\right)\beta_{2}+2i\eta_{2}+\beta_{2}\left(\beta_{6}+C\right)\right\},\end{split}$
where we have used (5.85). Therefore, we obtain the following as a necessary and sufficient condition for (5.112) to hold.
(5.115) $2\eta_{1}\left(2\beta_{3}-\beta_{1}\right)\beta_{2}+2i\eta_{2}+\beta_{2}\left(\beta_{6}+C\right)=0,\\$
Using (5.84), we rewrite this in terms of $C$ and $\Gamma$ to obtain
(5.116) $i\Gamma\left(\beta_{2}\left(2\beta_{3}-\beta_{1}\right)-\frac{\beta_{4}+\beta_{4}^{*}}{\beta_{2}}\right)+\beta_{6}+C+\left(\beta_{5}+\beta_{5}^{*}\right)\left(2\beta_{3}-\beta_{1}\right)-\frac{\beta_{5}-\beta_{5}^{*}}{\beta_{2}}=0.$
Next is the calculation of (5.113). The following holds
(5.117) $\begin{split}&F_{n}^{4}\langle\dot{\gamma}_{n},\dot{\gamma}_{n}\rangle\\ =&\left(\eta_{1}Q_{n}^{*}+i\eta_{2}P_{n}^{*}+\eta_{1}Q_{n}-i\eta_{2}P_{n}\right)^{2}\\ &-\left(\eta_{1}Q_{n}^{*}+i\eta_{2}P_{n}^{*}-\eta_{1}Q_{n}+i\eta_{2}P_{n}\right)^{2}\\ &+\left\{\left(\beta_{6}+C\right)F_{n}^{2}+\beta_{7}|G_{n}|^{2}\right\}^{2}\\ =&F_{n}^{4}\left\{16\eta_{1}^{2}\beta_{3}\left(\beta_{1}-\beta_{3}\right)+\left(\beta_{6}+C\right)^{2}\right\}\\ &+F_{n}^{2}|G_{n}|^{2}\left\{4\eta_{1}^{2}\left(1+2\left(\beta_{1}-2\beta_{3}\right)\left(\beta_{4}+\beta_{4}^{*}\right)\right)-8i\eta_{1}\eta_{2}\beta_{2}\left(2\beta_{3}-\beta_{1}\right)+4\eta_{2}^{2}+2\beta_{7}\left(\beta_{6}+C\right)\right\}\\ &+|G_{n}|^{4}\left(-4\eta_{1}^{2}\left(\beta_{4}+\beta_{4}^{*}\right)^{2}-8i\eta_{1}\eta_{2}\beta_{2}\left(\beta_{4}+\beta_{4}^{*}\right)+4\eta_{2}^{2}\beta_{2}^{2}+\beta_{7}^{2}\right)\\ =&F_{n}^{4}\left\{16\eta_{1}^{2}\beta_{3}\left(\beta_{1}-\beta_{3}\right)+\left(\beta_{6}+C\right)^{2}\right\}\\ &+F_{n}^{2}|G_{n}|^{2}\left\{4\eta_{1}^{2}\left(1+2\left(\beta_{1}-2\beta_{3}\right)\left(\beta_{4}+\beta_{4}^{*}\right)\right)-8i\eta_{1}\eta_{2}\beta_{2}\left(2\beta_{3}-\beta_{1}\right)+4\eta_{2}^{2}+2\beta_{7}\left(\beta_{6}+C\right)\right\},\end{split}$
where we have used (5.86). Therefore, we obtain the following as a necessary and sufficient condition for (5.113) to hold.
(5.118) $4\eta_{1}^{2}\left(1+2\left(\beta_{1}-2\beta_{3}\right)\left(\beta_{4}+\beta_{4}^{*}\right)\right)-8i\eta_{1}\eta_{2}\beta_{2}\left(2\beta_{3}-\beta_{1}\right)+4\eta_{2}^{2}+2\beta_{7}\left(\beta_{6}+C\right)=0.$
By using (5.85) and (5.116), (5.118) can be written in terms of $\Gamma$ as follows.
(5.119) $i\beta_{4}\beta_{4}^{*}\left(\beta_{4}-\beta_{4}^{*}\right)\Gamma^{2}+2\beta_{4}\beta_{4}^{*}\left(\beta_{5}+\beta_{5}^{*}\right)\Gamma+i\left(\beta_{4}{\beta_{5}^{*}}^{2}-\beta_{4}^{*}\beta_{5}^{2}\right)=0.$
This is a quadratic equation in $\Gamma$ and has the following solutions.
(5.120) $\Gamma_{\pm}=\frac{-\beta_{4}\beta_{4}^{*}\left(\beta_{5}+\beta_{5}^{*}\right)\pm\sqrt{\beta_{4}\beta_{4}^{*}\left(\beta_{4}\beta_{5}^{*}+\beta_{4}^{*}\beta_{5}\right)^{2}}}{i\beta_{4}\beta_{4}^{*}\left(\beta_{4}-\beta_{4}^{*}\right)}.$
Since
(5.121) $\begin{split}&\beta_{4}\beta_{4}^{*}\left(\beta_{4}\beta_{5}^{*}+\beta_{4}^{*}\beta_{5}\right)^{2}=\left(\beta_{4}\beta_{4}^{*}\right)^{2}\left(i\frac{\vartheta_{1}^{{}^{\prime}}(0)}{\vartheta_{1}(iv)}\right)^{2},\\ &\beta_{4}-\beta_{4}^{*}=\frac{\vartheta_{1}(2r)\vartheta_{3}(0)}{\vartheta_{3}(2r)\vartheta_{1}(iv)},\\ &\beta_{5}+\beta_{5}^{*}=-i\frac{\vartheta_{1}^{{}^{\prime}}(2r)\vartheta_{3}(0)}{\vartheta_{3}(2r)\vartheta_{1}(iv)},\end{split}$
holds, $\Gamma_{\pm}$ can be rewritten as
(5.122) $\Gamma_{\pm}=\frac{\vartheta_{1}^{{}^{\prime}}(2r)\vartheta_{3}(0)\pm\vartheta_{3}\left(2r\right)\vartheta_{1}^{{}^{\prime}}(0)}{\vartheta_{1}\left(2r\right)\vartheta_{3}\left(0\right)}.$
Substituting this into (5.116) yields the expression for $C_{\pm}$ as well. From the above calculations, we see that when $C_{\pm}$ and $\Gamma_{\pm}$ are chosen as in (5.109), the following holds
(5.123) $\langle\dot{\gamma}_{n},\dot{\gamma}_{n}\rangle=16\eta_{1}^{2}\beta_{3}\left(\beta_{1}-\beta_{3}\right)+\left(\beta_{6}+C\right)^{2}.$
We calculate this to compute $\rho^{2}$. From (5.84), we see that when $(C,\Gamma)=(C_{+},\Gamma_{+})$, the following holds
(5.124) $16\eta_{1}^{2}\beta_{3}\left(\beta_{1}-\beta_{3}\right)=4\left(i\frac{\vartheta_{1}^{{}^{\prime}}(0)}{\vartheta_{1}(iv)}\right)^{2}\frac{u_{1}^{2}}{\vartheta_{3}(2r)^{2}\vartheta_{3}(0)^{6}}.$
From (5.115) and (5.85), we see that when $(C,\Gamma)=(C_{+},\Gamma_{+})$, the following holds
(5.125) $\left(\beta_{6}+C\right)^{2}=\left(i\frac{\vartheta_{1}^{{}^{\prime}}(0)}{\vartheta_{1}(iv)}\right)^{2}\frac{u_{1}^{2}}{\vartheta_{3}(2r)^{2}\vartheta_{3}(0)^{6}}\frac{\left(\vartheta_{3}(r)^{4}-\vartheta_{1}(r)^{4}\right)^{2}}{\vartheta_{3}(r)^{4}\vartheta_{1}(r)^{4}}.$
Adding the above equations yields
(5.126) $\rho^{2}=16\eta_{1}^{2}\beta_{3}\left(\beta_{1}-\beta_{3}\right)+\left(\beta_{6}+C\right)^{2}=\left(i\frac{\vartheta_{1}^{{}^{\prime}}(0)}{\vartheta_{1}(iv)}\right)^{2}\frac{u_{1}^{2}}{\vartheta_{3}(r)^{4}\vartheta_{1}(r)^{4}}=\rho_{+}^{2},$
where we have used
(5.127) $\vartheta_{3}(r)^{4}+\vartheta_{1}(r)^{4}=\vartheta_{3}(2r)\vartheta_{3}(0)^{3}.$
In the same way, we see that when $(C,\Gamma)=(C_{-},\Gamma_{-})$,
(5.128) $\rho^{2}=16\eta_{1}^{2}\left(\beta_{1}-\beta_{3}\right)\beta_{3}+\left(\beta_{6}+C\right)^{2}=\left(i\frac{\vartheta_{1}^{{}^{\prime}}(0)}{\vartheta_{1}(iv)}\right)^{2}\frac{u_{1}^{2}}{\vartheta_{4}(r)^{4}\vartheta_{2}(r)^{4}}=\rho_{-}^{2},$
holds. ∎
Remark 5.12.
A discrete analogue of the constant negative curvature surface, called the discrete K-surface, is introduced in [2]. It is discussed that certain discrete-time deformation of discrete curves with a constant torsion angle generates discrete K-surfaces in [18]. Deformation equation (3.24) can be seen as a continuous-time analogue of the discrete-time deformation. Theorem 5.11 shows that specialising $(C,\Gamma)$ to $(C_{-},\Gamma_{-})$ or $(C_{+},\Gamma_{+})$, the curve constructed in Theorem 5.1 deforms according to (3.24). Therefore, its sweeping surface forms a semi-discrete analogue of the discrete K-surface (see figures in §7).
5.3. Evolution of curvature
The curve constructed in Theorem 5.1 reveals an interesting recurrence equation for $\kappa$. To obtain the recurrence equation, we use the following Lemma.
Lemma 5.13.
The $\tau$ functions (5.1) satisfy the following equations.
(5.129) $\widetilde{G}_{n}^{2}+\beta_{1}^{2}F_{n}^{2}=F_{n+1}F_{n-1},$
(5.130) $\beta_{1}^{2}\widetilde{G}_{n}^{2}-\frac{\vartheta_{1}\left(iv\right)^{4}}{\vartheta_{3}\left(iv\right)^{4}}\beta_{1}^{4}F_{n}^{2}=\widetilde{G}_{n+1}\widetilde{G}_{n-1}.$
Proof.
Using (3.1)-(3.3), we see that (5.129) hold. Putting
[…]
in (A.10), we obtain (5.130). ∎
Proposition 5.14.
The signed curvature angle of the curve obtained in Theorem 5.1 satisfies the following recurrence equation for any $t$ and $n$:
(5.136) $\tan\frac{\kappa_{n+1}}{2}\tan\frac{\kappa_{n-1}}{2}-\tan^{2}\frac{\kappa_{n}}{2}+\tan\frac{\kappa_{n+1}}{2}\tan^{2}\frac{\kappa_{n}}{2}\tan\frac{\kappa_{n-1}}{2}=-\frac{\vartheta_{1}\left(iv\right)^{4}}{\vartheta_{3}\left(iv\right)^{4}}.$
The right-hand side of this equation depends neither on $n$ nor on $t$. Therefore, the quantity on the left-hand side can be considered a conserved quantity of the signed curvature angle.
Proof.
Using (5.129) and (5.130), we see that (5.136) holds. ∎
The evolution of the signed curvature angle satisfies the semi-discrete mKdV equation in Proposition 5.9. To obtain the semi-discrete mKdV equation, we use the following Lemma.
Lemma 5.15.
The $\tau$ functions (5.1) satisfy the following equation.
(5.137) $D_{t}\widetilde{G}_{n}\cdot F_{n}=\beta_{8}\left(\widetilde{G}_{n+1}F_{n-1}-\widetilde{G}_{n-1}F_{n+1}\right),$
where
(5.138) $\beta_{8}=i\frac{\vartheta_{1}\left(iv\right)\vartheta_{3}\left(iv\right)}{\vartheta_{1}\left(2iv\right)\vartheta_{3}\left(0\right)}\left(d_{1}(iv;y)-d_{3}(iv;y)\right).$
Proof.
After differentiating the equation (A.3) with
[…]
by $h$ and putting $h=0$, we obtain
(5.144) $\begin{split}&D_{t}\left\{-i\vartheta_{1}(iv)\vartheta_{4}\left(\mu_{n}-\frac{1}{2}iv+it\right)\right\}\cdot\left\{\vartheta_{3}(iv)\vartheta_{2}\left(\mu_{n}-\frac{1}{2}iv+it\right)\right\}\\ &=\frac{\vartheta_{1}(iv)\vartheta_{3}(iv)}{\vartheta_{1}(2iv)\vartheta_{3}(0)}\frac{\vartheta^{{}^{\prime}}_{1}\left(iv\right)\vartheta_{3}\left(iv\right)-\vartheta_{1}\left(iv\right)\vartheta^{{}^{\prime}}_{3}\left(iv\right)}{i\vartheta_{1}\left(iv\right)\vartheta_{3}\left(iv\right)}\\ &\times\left\{-i\vartheta_{1}(iv)\vartheta_{4}\left(\mu_{n-1}-\frac{1}{2}iv+it\right)\right\}\left\{\vartheta_{3}(iv)\vartheta_{2}\left(\mu_{n+1}-\frac{1}{2}iv+it\right)\right\}\\ &-\frac{\vartheta_{1}(iv)\vartheta_{3}(iv)}{\vartheta_{1}(2iv)\vartheta_{3}(0)}\frac{\vartheta^{{}^{\prime}}_{1}\left(iv\right)\vartheta_{3}\left(iv\right)-\vartheta_{1}\left(iv\right)\vartheta^{{}^{\prime}}_{3}\left(iv\right)}{i\vartheta_{1}\left(iv\right)\vartheta_{3}\left(iv\right)}\\ &\times\left\{-i\vartheta_{1}(iv)\vartheta_{4}\left(\mu_{n+1}-\frac{1}{2}iv+it\right)\right\}\left\{\vartheta_{3}(iv)\vartheta_{2}\left(\mu_{n-1}-\frac{1}{2}iv+it\right)\right\}.\end{split}$
After rearranging (5.144), we obtain (5.137). ∎
Theorem 5.16.
The signed curvature angle (5.3) satisfies
(5.145) $\dot{\kappa}_{n}=2\beta_{1}\left(\eta_{1}\cos\lambda+\eta_{2}\sin\lambda\right)\left(\tan\frac{\kappa_{n+1}}{2}-\tan\frac{\kappa_{n-1}}{2}\right).$
Proof.
Differentiating the signed curvature angle yields the following.
(5.146) $\dot{\kappa}_{n}=2\frac{d}{dt}\arctan\left(\frac{\widetilde{G}_{n}}{\beta_{1}F_{n}}\right)=2\beta_{1}\frac{D_{t}\widetilde{G}_{n}\cdot F_{n}}{\widetilde{G}_{n}^{2}+\beta_{1}^{2}F_{n}^{2}}.$
Using (5.129) and (5.137), we see that the signed curvature angle satisfies
(5.147) $\dot{\kappa}_{n}=2\beta_{1}^{2}\beta_{8}\left(\tan\frac{\kappa_{n+1}}{2}-\tan\frac{\kappa_{n-1}}{2}\right).$
Since
(5.148) $\beta_{1}\beta_{8}=\eta_{1}\cos\lambda+\eta_{2}\sin\lambda,$
holds, this is nothing but the semi-discrete mKdV equation (5.145). ∎
5.4. Potential function of curvature
A family of functions $\Theta_{n}$ parameterised by $n\in\mathbb{Z}$ is said to be a potential of the curvature $\kappa$ if the equation $2\kappa_{n}=\Theta_{n+1}-\Theta_{n-1}$ holds. When $\gamma$ deforms by (3.24), the potential function $\Theta$ is known to satisfy the semi-discrete potential mKdV equation or the semi-discrete sine-Gordon equation [21]. Here, we derive the evolution equations satisfied by the potential function for the curve constructed by Theorem 5.1.
We construct a potential function which simultaneously satisfy (3.25) and (3.26) in terms of elliptic theta functions.
Theorem 5.17.
For fixed $v$ and $y$, define the function $\Theta$ as follows.
(5.149) $\Theta_{n}\left(t\right)=\Theta_{n}\left(t;v,y\right)=4\arctan\frac{\vartheta_{1}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y};\frac{2i}{y}\right)}{\vartheta_{4}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y};\frac{2i}{y}\right)}.$
Then we see that $\kappa$ defined in (5.3) satisfies
(5.150) $\kappa_{n}=\frac{\Theta_{n+1}-\Theta_{n-1}}{2}.$
Proof.
From (A.17), (A.18) and (A.19), we see that
(5.151) $\begin{split}&\kappa_{n}=2\arctan\left(-i\frac{\vartheta_{1}(iv;iy)\vartheta_{4}\left(iv\left(n-\frac{1}{2}\right)+it;iy\right)}{\vartheta_{3}(iv;iy)\vartheta_{2}\left(iv\left(n-\frac{1}{2}\right)+it;iy\right)}\right)\\ &=2\arctan\left(-i\frac{\frac{\vartheta_{1}\left(i\frac{v}{2}\left(n+\frac{1}{2}\right)+\frac{it}{2};\frac{iy}{2}\right)}{\vartheta_{2}\left(i\frac{v}{2}\left(n+\frac{1}{2}\right)+\frac{it}{2};\frac{iy}{2}\right)}-\frac{\vartheta_{1}\left(i\frac{v}{2}\left(n-\frac{3}{2}\right)+\frac{it}{2};\frac{iy}{2}\right)}{\vartheta_{2}\left(i\frac{v}{2}\left(n-\frac{3}{2}\right)+\frac{it}{2};\frac{iy}{2}\right)}}{1-\frac{\vartheta_{1}\left(i\frac{v}{2}\left(n+\frac{1}{2}\right)+\frac{it}{2};\frac{iy}{2}\right)\vartheta_{1}\left(i\frac{v}{2}\left(n-\frac{3}{2}\right)+\frac{it}{2};\frac{iy}{2}\right)}{\vartheta_{2}\left(i\frac{v}{2}\left(n+\frac{1}{2}\right)+\frac{it}{2};\frac{iy}{2}\right)\vartheta_{2}\left(i\frac{v}{2}\left(n-\frac{3}{2}\right)+\frac{it}{2};\frac{iy}{2}\right)}}\right)\\ &=2\arctan\left(\frac{\frac{\vartheta_{1}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y};\frac{2i}{y}\right)}{\vartheta_{4}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y};\frac{2i}{y}\right)}-\frac{\vartheta_{1}\left(\frac{v}{y}\left(n-\frac{3}{2}\right)+\frac{t}{y};\frac{2i}{y}\right)}{\vartheta_{4}\left(\frac{v}{y}\left(n-\frac{3}{2}\right)+\frac{t}{y};\frac{2i}{y}\right)}}{1+\frac{\vartheta_{1}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y};\frac{2i}{y}\right)\vartheta_{1}\left(\frac{v}{y}\left(n-\frac{3}{2}\right)+\frac{t}{y};\frac{2i}{y}\right)}{\vartheta_{4}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y};\frac{2i}{y}\right)\vartheta_{4}\left(\frac{v}{y}\left(n-\frac{3}{2}\right)+\frac{t}{y};\frac{2i}{y}\right)}}\right)\\ &=2\arctan\left(\frac{\vartheta_{1}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y};\frac{2i}{y}\right)}{\vartheta_{4}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y};\frac{2i}{y}\right)}\right)\\ &-2\arctan\left(\frac{\vartheta_{1}\left(\frac{v}{y}\left(n-\frac{3}{2}\right)+\frac{t}{y};\frac{2i}{y}\right)}{\vartheta_{4}\left(\frac{v}{y}\left(n-\frac{3}{2}\right)+\frac{t}{y};\frac{2i}{y}\right)}\right)\\ &=\frac{\Theta_{n+1}-\Theta_{n-1}}{2},\end{split}$
holds. ∎
The potential function simultaneously solves the semi-discrete potential mKdV equation and the semi-discrete sine-Gordon equation, establishing an explicit relation between our construction and the previous work [21] reviewed in §3.2. Since the semi-discrete sine-Gordon equation can be solved by almost the same calculation, we will solve only the semi-discrete potential mKdV equation here. We construct a solution to the equation using the theory of $\tau$ functions [17]. For this purpose, we first define the following real-valued functions for fixed $v$ and $y$.
(5.152) $\begin{split}&A_{n}\left(t\right)=\vartheta_{4}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y};\frac{2i}{y}\right),\\ &B_{n}\left(t\right)=\vartheta_{1}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y};\frac{2i}{y}\right).\end{split}$
Then the potential function can be written as follows
(5.153) $\Theta_{n}\left(t\right)=4\arctan\frac{B_{n}}{A_{n}}=\frac{2}{i}\log\frac{A_{n}+iB_{n}}{A_{n}-iB_{n}}=\frac{2}{i}\log\frac{\widetilde{\tau}_{n}}{\widetilde{\tau}^{*}_{n}},$
where $\widetilde{\tau}_{n}=A_{n}+iB_{n}$. $\widetilde{\tau}$ corresponds to the $\tau$ functions of the one-component KP hierarchy, while $A$ and $B$ correspond to the $\tau$ functions of the two-component KP hierarchy. Then we have the following lemma.
Lemma 5.18.
The functions $A$ and $B$ defined by (5.152) satisfy the following equations.
(5.154) $D_{t}A_{n+1}\cdot A_{n}+D_{t}B_{n+1}\cdot B_{n}=\frac{L_{1}}{2y}\left(\beta_{9}-1\right)\left(A_{n+1}A_{n}+B_{n+1}B_{n}\right),$
(5.155) $D_{t}A_{n+1}\cdot B_{n}-D_{t}B_{n+1}\cdot A_{n}=\frac{L_{1}}{2y}\left(\beta_{9}+1\right)\left(A_{n+1}B_{n}-B_{n+1}A_{n}\right),$
where
(5.156) $J_{1}(v,y)=\frac{\vartheta^{{}^{\prime}}_{1}\left(0;\frac{2i}{y}\right)\vartheta_{1}\left(\frac{v}{y};\frac{2i}{y}\right)}{\vartheta_{4}\left(0;\frac{2i}{y}\right)\vartheta_{4}\left(\frac{v}{y};\frac{2i}{y}\right)},\quad J_{2}(v,y)=\frac{\vartheta^{{}^{\prime}}_{1}\left(0;\frac{2i}{y}\right)\vartheta_{4}\left(\frac{v}{y};\frac{2i}{y}\right)}{\vartheta_{1}\left(\frac{v}{y};\frac{2i}{y}\right)\vartheta_{4}\left(0;\frac{2i}{y}\right)},$
(5.157) $L_{1}(v,y)=d_{1}\left(\frac{v}{y};\frac{2}{y}\right)-d_{4}\left(\frac{v}{y};\frac{2}{y}\right)+J_{1}(v,y)+J_{2}(v,y),$
(5.158) $\beta_{9}=\frac{d_{1}\left(\frac{v}{y};\frac{2}{y}\right)+d_{4}\left(\frac{v}{y};\frac{2}{y}\right)-J_{1}(v,y)+J_{2}(v,y)}{L_{1}(v,y)}.$
Proof.
In this proof, all elliptic theta functions have $w=2i/y$ and we denote $\vartheta_{j}(v;2i/y)$ by $\vartheta_{j}(v)$. After differentiating the equation in (A.6) with
[…]
by $h$ and putting $h=0$, we obtain
(5.164) $\begin{split}&D_{t}\vartheta_{4}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\cdot\vartheta_{4}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &=\frac{1}{y}\frac{\vartheta^{{}^{\prime}}_{4}\left(\frac{v}{y}\right)}{\vartheta_{4}\left(\frac{v}{y}\right)}\vartheta_{4}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{4}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &-\frac{1}{y}\frac{\vartheta_{1}\left(\frac{v}{y}\right)\vartheta^{{}^{\prime}}_{1}\left(0\right)}{\vartheta_{4}\left(\frac{v}{y}\right)\vartheta_{4}\left(0\right)}\vartheta_{1}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{1}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &=\frac{1}{y}d_{4}\left(\frac{v}{y};\frac{2}{y}\right)\vartheta_{4}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{4}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &-\frac{1}{y}J_{1}(v,y)\vartheta_{1}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{1}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right).\end{split}$
Also, after differentiating the equation in (A.7) with
[…]
by $h$ and putting $h=0$, we obtain
(5.170) $\begin{split}&D_{t}\vartheta_{1}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\cdot\vartheta_{1}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &=\frac{1}{y}\frac{\vartheta^{{}^{\prime}}_{4}\left(\frac{v}{y}\right)}{\vartheta_{4}\left(\frac{v}{y}\right)}\vartheta_{1}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{1}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &-\frac{1}{y}\frac{\vartheta_{1}\left(\frac{v}{y}\right)\vartheta^{{}^{\prime}}_{1}\left(0\right)}{\vartheta_{4}\left(\frac{v}{y}\right)\vartheta_{4}\left(0\right)}\vartheta_{4}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{4}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &=\frac{1}{y}d_{4}\left(\frac{v}{y};\frac{2}{y}\right)\vartheta_{1}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{1}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &-\frac{1}{y}J_{1}(v,y)\vartheta_{4}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{4}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right).\end{split}$
Also, after differentiating the equation in (A.7) with
[…]
by $h$ and putting $h=0$, we obtain
(5.176) $\begin{split}&D_{t}\vartheta_{1}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\cdot\vartheta_{4}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &=\frac{1}{y}\frac{\vartheta^{{}^{\prime}}_{1}\left(\frac{v}{y}\right)}{\vartheta_{1}\left(\frac{v}{y}\right)}\vartheta_{1}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{4}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &-\frac{1}{y}\frac{\vartheta_{4}\left(\frac{v}{y}\right)\vartheta^{{}^{\prime}}_{1}\left(0\right)}{\vartheta_{1}\left(\frac{v}{y}\right)\vartheta_{4}\left(0\right)}\vartheta_{4}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{1}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &=\frac{1}{y}d_{1}\left(\frac{v}{y};\frac{2}{y}\right)\vartheta_{1}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{4}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &-\frac{1}{y}J_{2}(v,y)\vartheta_{4}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{1}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right).\end{split}$
Also, after differentiating the equation in (A.7) with
[…]
by $h$ and putting $h=0$, we obtain
(5.182) $\begin{split}&D_{t}\vartheta_{4}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\cdot\vartheta_{1}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &=\frac{1}{y}\frac{\vartheta^{{}^{\prime}}_{1}\left(\frac{v}{y}\right)}{\vartheta_{1}\left(\frac{v}{y}\right)}\vartheta_{4}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{1}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &-\frac{1}{y}\frac{\vartheta_{4}\left(\frac{v}{y}\right)\vartheta^{{}^{\prime}}_{1}\left(0\right)}{\vartheta_{1}\left(\frac{v}{y}\right)\vartheta_{4}\left(0\right)}\vartheta_{1}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{4}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &=\frac{1}{y}d_{1}\left(\frac{v}{y};\frac{2}{y}\right)\vartheta_{4}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{1}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &-\frac{1}{y}J_{2}(v,y)\vartheta_{1}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{4}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right).\end{split}$
Adding (5.164) and (5.170) yields (5.154), and subtracting (5.176) from (5.182) yields (5.155). ∎
Then we have the following theorem.
Theorem 5.19.
Let $y>0$ and $v\in\mathbb{R}\setminus{(y\mathbb{Z})}$. The potential function $\Theta$ defined by (5.149) simultaneously satisfies the semi-discrete potential mKdV and sine-Gordon equations
(5.183) $\frac{d}{dt}\left(\frac{\Theta_{n+1}+\Theta_{n}}{2}\right)$ $=\left(1+\cos\lambda\right)\rho_{+}\sin\left(\frac{\Theta_{n+1}-\Theta_{n}}{2}\right),$
(5.184) $\frac{d}{dt}\left(\frac{\Theta_{n+1}-\Theta_{n}}{2}\right)$ $=\left(1-\cos\lambda\right)\rho_{-}\sin\left(\frac{\Theta_{n+1}+\Theta_{n}}{2}\right).$
Proof.
We first show that the potential function satisfies
(5.185) $\frac{d}{dt}\left(\frac{\Theta_{n+1}+\Theta_{n}}{2}\right)$ $=\frac{L_{1}(v,y)}{y}\sin\left(\frac{\Theta_{n+1}-\Theta_{n}}{2}\right),$
using Lemma 5.18. Then we show that
(5.186) $\frac{L_{1}(v,y)}{y}\frac{1}{1+\cos\lambda}=i\frac{\vartheta_{1}^{{}^{\prime}}(0)}{\vartheta_{1}(iv)}\frac{u_{1}}{\vartheta_{3}(r)^{2}\vartheta_{1}(r)^{2}}=\rho_{+}.$
Suppose that a solution to the equation (5.185) is expressed as (5.153). Then both sides of the equation can be rewritten as follows.
(5.187) $\begin{split}&\frac{d}{dt}\left(\frac{\Theta_{n+1}+\Theta_{n}}{2}\right)=\frac{1}{i}\frac{d}{dt}\left(\log\frac{\widetilde{\tau}_{n+1}}{\widetilde{\tau}^{*}_{n+1}}+\log\frac{\widetilde{\tau}_{n}}{\widetilde{\tau}^{*}_{n}}\right)\\ &=\frac{1}{i}\frac{d}{dt}\left(\log\frac{\widetilde{\tau}_{n+1}}{\widetilde{\tau}^{*}_{n}}-\log\frac{\widetilde{\tau}^{*}_{n+1}}{\widetilde{\tau}_{n}}\right)\\ &=\frac{1}{i}\left(\frac{D_{t}\widetilde{\tau}_{n+1}\cdot\widetilde{\tau}^{*}_{n}}{\widetilde{\tau}_{n+1}\widetilde{\tau}^{*}_{n}}-\frac{D_{t}\widetilde{\tau}^{*}_{n+1}\cdot\widetilde{\tau}_{n}}{\widetilde{\tau}^{*}_{n+1}\widetilde{\tau}_{n}}\right),\end{split}$
and
(5.188) $\begin{split}&\sin\left(\frac{\Theta_{n+1}-\Theta_{n}}{2}\right)=\frac{1}{2i}\left(\frac{\widetilde{\tau}_{n+1}\widetilde{\tau}^{*}_{n}}{\widetilde{\tau}^{*}_{n+1}\widetilde{\tau}_{n}}-\frac{\widetilde{\tau}^{*}_{n+1}\widetilde{\tau}_{n}}{\widetilde{\tau}_{n+1}\widetilde{\tau}^{*}_{n}}\right).\end{split}$
Then it can be seen that a sufficient condition for (5.153) to be a solution of (5.185) is that
(5.189) $\begin{split}&D_{t}\widetilde{\tau}_{n+1}\cdot\widetilde{\tau}^{*}_{n}=-\frac{L_{1}}{2y}\widetilde{\tau}^{*}_{n+1}\widetilde{\tau}_{n}+\frac{L_{1}}{2y}\beta_{9}\widetilde{\tau}_{n+1}\widetilde{\tau}_{n}^{*},\end{split}$
is satisfied. Substituting $\widetilde{\tau}_{n}=A_{n}+iB_{n}$ for this, we get
(5.190) $\begin{split}&D_{t}\left(A_{n+1}\cdot A_{n}+B_{n+1}\cdot B_{n}\right)-iD_{t}\left(A_{n+1}\cdot B_{n}-B_{n+1}\cdot A_{n}\right)\\ &=\frac{L_{1}}{2y}\left(\beta_{9}-1\right)\left(A_{n+1}A_{n}+B_{n+1}B_{n}\right)-i\frac{L_{1}}{2y}\left(\beta_{9}+1\right)\left(A_{n+1}B_{n}-B_{n+1}A_{n}\right).\end{split}$
The real part of this equation yields (5.154), and the imaginary part yields (5.155). Thus, Lemma 5.18 shows that the potential function satisfies the equation (5.185). We then calculate $L_{1}(v,y)/y$. From (A.15), (A.16) and (A.20), we see
(5.191) $\begin{split}\frac{L_{1}(v,y)}{y}&=\frac{1}{y}\frac{\vartheta_{1}^{{}^{\prime}}\left(\frac{v}{y};\frac{2i}{y}\right)\vartheta_{4}\left(\frac{v}{y};\frac{2i}{y}\right)-\vartheta_{1}\left(\frac{v}{y};\frac{2i}{y}\right)\vartheta_{4}^{{}^{\prime}}\left(\frac{v}{y};\frac{2i}{y}\right)}{\vartheta_{1}\left(\frac{v}{y};\frac{2i}{y}\right)\vartheta_{4}\left(\frac{v}{y};\frac{2i}{y}\right)}\\ &+\frac{1}{y}\frac{\vartheta_{1}^{{}^{\prime}}\left(0;\frac{2i}{y}\right)\vartheta_{1}\left(\frac{v}{y};\frac{2i}{y}\right)}{\vartheta_{4}\left(0;\frac{2i}{y}\right)\vartheta_{4}\left(\frac{v}{y};\frac{2i}{y}\right)}+\frac{1}{y}\frac{\vartheta_{1}^{{}^{\prime}}\left(0;\frac{2i}{y}\right)\vartheta_{4}\left(\frac{v}{y};\frac{2i}{y}\right)}{\vartheta_{4}\left(0;\frac{2i}{y}\right)\vartheta_{1}\left(\frac{v}{y};\frac{2i}{y}\right)}\\ &=i\frac{\vartheta_{1}^{{}^{\prime}}\left(0;iy\right)}{\vartheta_{4}\left(0;iy\right)}\frac{\vartheta_{4}\left(iv;iy\right)}{\vartheta_{1}\left(iv;iy\right)}+i\frac{\vartheta_{1}^{{}^{\prime}}\left(0;iy\right)}{\vartheta_{2}\left(0;iy\right)}\frac{\vartheta_{2}\left(iv;iy\right)}{\vartheta_{1}\left(iv;iy\right)}\\ &=i\frac{\vartheta_{1}^{{}^{\prime}}\left(0;iy\right)}{\vartheta_{1}\left(iv;iy\right)}\frac{\vartheta_{4}\left(iv;iy\right)\vartheta_{2}\left(0;iy\right)+\vartheta_{2}\left(iv;iy\right)\vartheta_{4}\left(0;iy\right)}{\vartheta_{4}\left(0;iy\right)\vartheta_{2}\left(0;iy\right)}.\end{split}$
On the other hand, we see
(5.192) $\begin{split}&1+\cos\lambda\\ &=\frac{\left\{\vartheta_{1}\left(\frac{iv}{2}+r;iy\right)\vartheta_{3}\left(-\frac{iv}{2}+r;iy\right)+\vartheta_{1}\left(-\frac{iv}{2}+r;iy\right)\vartheta_{3}\left(\frac{iv}{2}+r;iy\right)\right\}^{2}}{2u_{1}}.\end{split}$
We can see that (5.186) holds by rewriting (5.191) and (5.192) using the following.
(5.193) $\begin{split}&\vartheta_{2}\left(iv/2\right)\vartheta_{4}\left(iv/2\right)\left\{\vartheta_{1}\left(iv/2+r\right)\vartheta_{3}\left(-iv/2+r\right)+\vartheta_{1}\left(-iv/2+r\right)\vartheta_{3}\left(iv/2+r\right)\right\}\\ &=\vartheta_{3}\left(r\right)\vartheta_{1}\left(r\right)\left\{\vartheta_{4}\left(iv\right)\vartheta_{2}\left(0\right)+\vartheta_{2}\left(iv\right)\vartheta_{4}\left(0\right)\right\},\\ &\vartheta_{3}\left(r\right)\vartheta_{1}\left(r\right)\left\{\vartheta_{1}\left(iv/2+r\right)\vartheta_{3}\left(-iv/2+r\right)+\vartheta_{1}\left(-iv/2+r\right)\vartheta_{3}\left(iv/2+r\right)\right\}\\ &=\vartheta_{2}\left(iv/2\right)\vartheta_{4}\left(iv/2\right)\left\{\vartheta_{4}\left(2r\right)\vartheta_{2}\left(0\right)-\vartheta_{2}\left(2r\right)\vartheta_{4}\left(0\right)\right\},\\ &\vartheta_{2}\left(0\right)\vartheta_{4}\left(0\right)\left\{\vartheta_{4}\left(2r\right)\vartheta_{2}\left(0\right)-\vartheta_{2}\left(2r\right)\vartheta_{4}\left(0\right)\right\}\\ &=2\vartheta_{3}\left(r\right)^{2}\vartheta_{1}\left(r\right)^{2}.\end{split}$
Remark 5.20.
Rewriting (5.145) in terms of the potential function $\Theta$, we obtain
(5.194) $\frac{d}{dt}\Theta_{n}=4\beta_{1}\left(\eta_{1}\cos\lambda+\eta_{2}\sin\lambda\right)\tan\left(\frac{\Theta_{n+1}-\Theta_{n-1}}{4}\right),$
which is also called the semi-discrete potential mKdV equation [16, 15, 7]. In other words, the potential function $\Theta$ defined by (5.149) is a simultaneous solution of three differential-difference equations, (5.183), (5.184) and (5.194). It does not usually happen that these equations are simultaneously solved. For example, it can be checked that the 2- and 3-soliton solutions of the semi-discrete sine-Gordon equation constructed in [7, 8] are not solutions of the semi-discrete potential mKdV equation.
Note the difference in approaches of [21] and this work; [21] starts with a geometric property observed in the motion of Kaleidocycles, and derive evolution equations, whereas we aim to construct a particular motion of Kaleidocycles and investigate the equations satisfied by them. They turn out to arrive at the same place; the latter gives an explicit solution to the equations derived in the former.
6. Existence of Kaleidocycles
The constant torsion-angle and unit segment-length curves given in Theorem 5.1 are not necessarily closed. In this section, we derive a necessary and sufficient condition on the parameters $v,r,y$ for the curve $\gamma$ to be closed to define a Kaleidocycle. Furthermore, we prove that there exist parameters that satisfy the closure condition when $k\geq 6$. The overall approach is as follows: For the curve $\gamma$ to be closed, the absolute value of the signed curvature angle has to have period $k$. From this, we see that $m:=kv/y$ is an integer (§6.1). Then, rewriting the closedness condition $\gamma_{n+k}(t)=\gamma_{n}(t)$ assuming $m$ is an integer, we obtain two equations for v,r,yv,r,y6.2). By rearranging one of the equations, we verify that $r=r(y)$ is uniquely determined when $y$ is in a certain range (§6.3). Substituting this $r(y)$ into the other equation, we find that if $m\geq 3$, then there exists $y$ satisfying it (§6.4). By considering two transformations of $m$ and $n$ that preserve the shape of a curve, we see that it is sufficient to consider $0<m\leq k/2$. Finally, we show that under this range of $m$ and $k\geq 6$, there always exists a set of parameters with which the curve Theorem 5.1 is closed, and hence, gives a $1$-dimensional periodic motion of a Kaleidocycle (§6.5).
6.1. Period of the curvature angle
We say a curve $\gamma$ is closed with a period $k\in\mathbb{N}$ when $\gamma_{n+k}(t)=\gamma_{n}(t)$ for all $n$ and $t$. In this case, the absolute value of the curvature angles also has a period $k$. Hence, we obtain the necessary condition for the curve to be closed that
(6.1) $\begin{split}\tan^{2}\frac{\kappa_{n+k}(t)}{2}-\tan^{2}\frac{\kappa_{n}(t)}{2}=0,\end{split}$
holds for all $n$ and $t$. Note that there is no Kaleidocycle with period $1$, hence $k\geq 2$. Rewriting this condition using Jacobi $\mathrm{cn}$-function, we obtain
(6.2) $\begin{split}&\tan^{2}\frac{\kappa_{n+k}(t)}{2}-\tan^{2}\frac{\kappa_{n}(t)}{2}\\ &=\frac{\vartheta_{1}\left(\frac{v}{y};\frac{i}{y}\right)^{2}\vartheta_{2}\left(0;\frac{i}{y}\right)^{2}}{\vartheta_{3}\left(\frac{v}{y};\frac{i}{y}\right)^{2}\vartheta_{4}\left(0;\frac{i}{y}\right)^{2}}\left\{\mathrm{cn}^{2}\left[2\tilde{K}\left(\tilde{\mu}_{n}(t)+\frac{v}{y}k\right)\right]-\mathrm{cn}^{2}\left[2\tilde{K}\tilde{\mu}_{n}(t)\right]\right\}=0,\end{split}$
where we note $\tilde{\mu}_{n}(t)=\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}$. Since $\tilde{\mu}_{n}(t)$ can be regarded as a linear function of $t$ when $v,y$ and $n$ are fixed, $\tilde{\mu}_{n}(t)$ attains all reals as $t$ varies. Thus, we only need to check that (6.2) holds for $n=0$. Since $\mathrm{cn}^{2}$ has a real period $2\tilde{K}$, $m=kv/y$ must be an integer. Hereafter, let $v$ satisfy
(6.3) $v=\frac{m}{k}y.$
6.2. Closure conditions
Here we obtain the necessary and sufficient conditions satisfied by $v,r,y$ for the curve to be closed. We first rewrite $\gamma_{n+k}(t)=\gamma_{n}(t)$ in terms of $\tau$ functions. We have
(6.4) $\frac{H_{n+k}}{F_{n+k}}=\frac{H_{n}}{F_{n}},$
$n+k-2\displaystyle\frac{\partial\log F_{n+k}}{\partial z}=n-2\displaystyle\frac{\partial\log F_{n}}{\partial z}.$
We rewrite the left-hand side further with (A.12) under the condition that $m$ is an integer. Then we obtain
(6.5) $\frac{H_{n+k}}{F_{n+k}}=\frac{H_{n}}{F_{n}}\left(R_{1}R_{3}\right)^{-k}\exp\left(4mr\pi i\right)(-1)^{m},$
$F_{n+k}=F_{n}\exp\left(k\frac{\Delta_{3}}{\Delta_{3}-\Delta_{1}}z\right)\exp\left[-\pi i\left\{2m\left(\mu_{n}-\frac{1}{2}iv\right)+m^{2}w\right\}\right],$
where $\mu_{n}=ivn+\frac{z}{\Delta_{3}-\Delta_{1}}$. Using (6.5), (6.4) can be rewritten as
(6.6) $\Lambda(v,r,y):=\exp\left\{\pi im\left(4r+1\right)\right\}\left(R_{1}R_{3}\right)^{-k}-1=0,$
(6.7) $\Xi(v,r,y):=\frac{\left(\Delta_{3}+\Delta_{1}\right)k-4\pi im}{\Delta_{3}-\Delta_{1}}=0.$
Note that from the definitions of $R_{j}$ and $\Delta_{j}$, we see that
(6.8) $\frac{\partial}{\partial r}\log R_{j}=\Delta_{j}\quad(j=1,2,3,4),$
holds. From this, we have
(6.9) $\frac{\partial}{\partial r}\log(\Lambda+1)=4\pi im-k\left(\Delta_{3}+\Delta_{1}\right)=-\left(\Delta_{3}-\Delta_{1}\right)\Xi.$
Since $\Delta_{3}-\Delta_{1}\neq 0$ holds, (6.7) can be written as
(6.10) $\frac{\partial}{\partial r}\log(\Lambda+1)=0.$
On the other hand, the closure condition (6.4) holds for any $n\in\mathbb{Z}$ and $t\in\mathbb{R}$ when the parameters $v,r,y$ satisfy $kv=my$, (6.6) and (6.10). Thus, they give necessary and sufficient condition for the curve given by Theorem 5.1 to be a closed curve for every $t\in\mathbb{R}$.
6.3. An explicit solution to the condition (6.10)
Here we solve (6.10) with respect to $r$. From the period of elliptic theta function, we see the period and symmetry of $\Lambda(v,r,y)$ with respect to $r$ as follows: $\Lambda(v,r+1,y)=\Lambda(v,r,y)$, $\Lambda(v,-r,y)=\Lambda(v,r,y)^{*}$. Therefore, when searching for $(v,r,y)$ such that $\Lambda(v,r,y)=0$, we do not lose generality by assuming $0\leq r\leq 1/2$. We exclude the case $r=0,1/2$ because $\gamma$ forms a plane curve. Hereafter, we assume $0<r<1/2$. Since
[…]
holds, we have
(6.12) $\begin{split}-&\frac{u_{1}}{k}\frac{\partial}{\partial r}\log(\Lambda+1)=-u_{1}\left(\Delta_{3}+\Delta_{1}\right)+4\pi i\frac{v}{y}u_{1}\\ &=\vartheta_{4}\left(0\right)\vartheta_{2}\left(0\right)\left(\vartheta_{4}\left(2r\right)\vartheta_{2}^{{}^{\prime}}\left(iv\right)-\vartheta_{2}\left(2r\right)\vartheta_{4}^{{}^{\prime}}\left(iv\right)\right)\\ &+2\pi i\frac{v}{y}\vartheta_{4}\left(0\right)\vartheta_{2}\left(0\right)\left(\vartheta_{4}\left(2r\right)\vartheta_{2}\left(iv\right)-\vartheta_{2}\left(2r\right)\vartheta_{4}\left(iv\right)\right)=0.\end{split}$
Arranging this we obtain
(6.13) $\frac{\vartheta_{4}\left(0\right)\vartheta_{2}\left(2r\right)}{\vartheta_{2}\left(0\right)\vartheta_{4}\left(2r\right)}=\frac{\vartheta_{4}\left(0\right)\left(2\pi i\frac{v}{y}\vartheta_{2}\left(iv\right)+\vartheta_{2}^{{}^{\prime}}\left(iv\right)\right)}{\vartheta_{2}\left(0\right)\left(2\pi i\frac{v}{y}\vartheta_{4}\left(iv\right)+\vartheta_{4}^{{}^{\prime}}\left(iv\right)\right)}.$
The right-hand side of this equation does not contain $r$. The left-hand side is nothing but Jacobi $\mathrm{cn}$-function.
(6.14) $\mathrm{cn}\left[4Kr,\xi\right]=\frac{\vartheta_{4}\left(0\right)\vartheta_{2}\left(2r\right)}{\vartheta_{2}\left(0\right)\vartheta_{4}\left(2r\right)},$
where $\xi=\vartheta_{2}\left(0\right)^{2}/\vartheta_{3}\left(0\right)^{2}$ is an elliptic modulus and $K=\pi\vartheta_{3}\left(0\right)^{2}/2$ is a complete elliptic integral of the first kind. We define the following function $\varphi$ for the ratio of $m$ and $k$ and for $y>0$.
(6.15) $\begin{split}\varphi\left(\frac{m}{k},y\right):=\frac{\vartheta_{4}\left(0\right)\left(2\pi i\frac{m}{k}\vartheta_{2}\left(i\frac{m}{k}y\right)+\vartheta_{2}^{{}^{\prime}}\left(i\frac{m}{k}y\right)\right)}{\vartheta_{2}\left(0\right)\left(2\pi i\frac{m}{k}\vartheta_{4}\left(i\frac{m}{k}y\right)+\vartheta_{4}^{{}^{\prime}}\left(i\frac{m}{k}y\right)\right)}&=\frac{\vartheta_{2}\left(0;\frac{i}{y}\right)\vartheta_{4}^{{}^{\prime}}\left(\frac{m}{k};\frac{i}{y}\right)}{\vartheta_{4}\left(0;\frac{i}{y}\right)\vartheta_{2}^{{}^{\prime}}\left(\frac{m}{k};\frac{i}{y}\right)}.\end{split}$
When $1<\varphi\left(m/k,y\right)<1$ is satisfied, $r=r(y)$ is uniquely determined because the inverse function of $\mathrm{cn}$ exists.
6.3.1. Existence of a inverse function
Here we show that for a certain range of $y$, $1<\varphi\left(m/k,y\right)\leq 0$ holds when $0<m\leq k/2$. Then we find $r=r(y)$ explicitly using elliptic integrals. First, from the zeros and evenness of elliptic theta function we see the following: For any $y>0$, $\varphi(1/2,y)=0$ and $\varphi(m/k,y)\neq 0\quad(0<m/k<1/2)$. Therefore, when $m/k=1/2$, $r=1/4$. Next, consider the limit of $\varphi(m/k,y)$. From the definition of elliptic theta function, when $u$ is fixed and $y\rightarrow+0$, we see that
(6.16) $\vartheta_{2}\left(u;\frac{i}{y}\right)$ $=2q^{\frac{1}{4}}\cos\pi u+\mathcal{O}(q^{\frac{9}{4}}),$
$\vartheta_{4}\left(u;\frac{i}{y}\right)$ $=1+\mathcal{O}(q),$
$\vartheta_{2}^{{}^{\prime}}\left(u;\frac{i}{y}\right)$ $=-2\pi q^{\frac{1}{4}}\sin\pi u+\mathcal{O}(q^{\frac{9}{4}}),$
$\vartheta_{4}^{{}^{\prime}}\left(u;\frac{i}{y}\right)$ $=4\pi q\sin 2\pi u+\mathcal{O}(q^{4}),$
where we denote $q=e^{-\frac{\pi}{y}}$. From this we have
(6.17) $\lim_{y\to+0}\varphi\left(\frac{m}{k},y\right)=\lim_{y\to+0}\frac{\vartheta_{2}\left(0;\frac{i}{y}\right)\vartheta_{4}^{{}^{\prime}}\left(\frac{m}{k};\frac{i}{y}\right)}{\vartheta_{4}\left(0;\frac{i}{y}\right)\vartheta_{2}^{{}^{\prime}}\left(\frac{m}{k};\frac{i}{y}\right)}=0\quad(0<\frac{m}{k}<1/2).$
On the other hand, the following hold when $y\rightarrow\infty$
(6.18) $p^{\left(\frac{m}{k}\right)^{2}}\vartheta_{2}\left(i\frac{m}{k}y;iy\right)$ $=p^{\left(\frac{1}{2}-\frac{m}{k}\right)^{2}}+\mathcal{O}(p^{\left(\frac{1}{2}+\frac{m}{k}\right)^{2}}),$
$p^{\left(\frac{m}{k}\right)^{2}}\vartheta_{4}\left(i\frac{m}{k}y;iy\right)$ $=p^{\left(\frac{m}{k}\right)^{2}}+\mathcal{O}(p^{\left(1-\frac{m}{k}\right)^{2}}),$
$ip^{\left(\frac{m}{k}\right)^{2}}\vartheta_{2}^{{}^{\prime}}\left(i\frac{m}{k}y;iy\right)$ $=\pi p^{\left(\frac{1}{2}-\frac{m}{k}\right)^{2}}+\mathcal{O}(p^{\left(\frac{1}{2}+\frac{m}{k}\right)^{2}}),$
$ip^{\left(\frac{m}{k}\right)^{2}}\vartheta_{4}^{{}^{\prime}}\left(i\frac{m}{k}y;iy\right)$ $=-2\pi p^{\left(1-\frac{m}{k}\right)^{2}}+\mathcal{O}(p^{\left(1+\frac{m}{k}\right)^{2}})\quad(0<\frac{m}{k}<1/2),$
where we denote $p=e^{-\pi y}$. From this we have
(6.19) $\lim_{y\to\infty}\varphi\left(\frac{m}{k},y\right)=\lim_{y\to\infty}\frac{\vartheta_{4}\left(0\right)\left(2\pi i\frac{m}{k}\vartheta_{2}\left(i\frac{m}{k}y\right)+\vartheta_{2}^{{}^{\prime}}\left(i\frac{m}{k}y\right)\right)}{\vartheta_{2}\left(0\right)\left(2\pi i\frac{m}{k}\vartheta_{4}\left(i\frac{m}{k}y\right)+\vartheta_{4}^{{}^{\prime}}\left(i\frac{m}{k}y\right)\right)}=-\infty\quad(0<\frac{m}{k}<1/2).$
Hence, when $m$ is fixed within $0<m<k/2$, there exists $y$ such that $\varphi\left(m/k,y\right)=-1$. We denote the smallest one among them by $y_{0}(m/k)$. Summarizing the above, when $m$ is fixed within $0<m<k/2$ and $y$ moves in the range $0<y<y_{0}(m/k)$, $r=r(y)$ satisfying (6.10) is uniquely determined within the range of $0<r<1/2$. By taking the inverse function of $\mathrm{cn}$ in (6.13), we obtain
(6.20) $\begin{split}&r\left(y;\frac{m}{k}\right)=\frac{1}{4K}\int^{1}_{\varphi(m/k,y)}\frac{dx}{\sqrt{(1-x^{2})(1-\xi^{2}+\xi^{2}x^{2})}}.\end{split}$
This also holds when $m/k=1/2$, hence we see that $1/4\leq r<1/2$. If we assume 1/2<r<0-1/2<r<0, the same calculation yields
(6.21) $\begin{split}r\left(y;\frac{m}{k}\right)=-\frac{1}{4K}\int^{1}_{\varphi(m/k,y)}\frac{dx}{\sqrt{(1-x^{2})(1-\xi^{2}+\xi^{2}x^{2})}}.\end{split}$
From this, we see $1/2<r\leq-1/4$ holds when $0<m\leq k/2$. The sign of $r$ is chosen so that $\sin\lambda=\langle\widetilde{B}_{n},\widetilde{N}_{n-1}\rangle>0$ holds.
6.4. Existence of a solution to the condition $\Lambda=0$
By substituting $r=r(y)$ obtained in the previous section into $\Lambda(v,r,y)$, we prove that $\Lambda(v,r(y),y)=0$ has a solution in $0<y<y_{0}$. Note that $\Lambda+1$ can also be written as
(6.22) $\begin{split}\Lambda+1&=\exp\left\{\pi im\left(4r+1\right)\right\}\left[\frac{\vartheta_{1}\left(\frac{1}{2}iv+r;iy\right)\vartheta_{3}\left(\frac{1}{2}iv+r;iy\right)}{\vartheta_{1}\left(-\frac{1}{2}iv+r;iy\right)\vartheta_{3}\left(-\frac{1}{2}iv+r;iy\right)}\right]^{k}\\ &=\exp{\left(\pi im\right)}\left[\frac{\vartheta_{1}\left(\frac{1}{2}\frac{m}{k}+\frac{r}{iy};\frac{i}{y}\right)\vartheta_{3}\left(\frac{1}{2}\frac{m}{k}+\frac{r}{iy};\frac{i}{y}\right)}{\vartheta_{1}\left(-\frac{1}{2}\frac{m}{k}+\frac{r}{iy};\frac{i}{y}\right)\vartheta_{3}\left(-\frac{1}{2}\frac{m}{k}+\frac{r}{iy};\frac{i}{y}\right)}\right]^{k},\end{split}$
by transforming the parameter $y\rightarrow 1/y$. Since $\Lambda+1$ is a complex number of magnitude $1$, one may denote
(6.23) $\Lambda+1=\exp{\left(i\Psi\right)},$
using a real-valued function $\Psi=\Psi(v,r,y)$. When
(6.24) $\left|\Psi(v,r(y_{0}),y_{0})-\lim_{y\to+0}\Psi(v,r(y),y)\right|,$
is greater than $2\pi$, then there exists a solution to $\Lambda(v,r(y),y)=0$ in the range $0<y<y_{0}(m/k)$. Here, we compute the above quantities and show that such a $y$ exists if $m\geq 3$. From a representation by an infinite product of elliptic theta functions,
(6.25) $\vartheta_{1}\left(u;\frac{i}{y}\right)$ $=2q^{\frac{1}{4}}\sin(\pi u)\prod_{n=1}^{\infty}\left(1-q^{2n}\right)\left(1-q^{2n}e^{2\pi iu}\right)\left(1-q^{2n}e^{-2\pi iu}\right),$
$\vartheta_{3}\left(u;\frac{i}{y}\right)$ $=\prod_{n=1}^{\infty}\left(1-q^{2n}\right)\left(1+q^{2n-1}e^{2\pi iu}\right)\left(1+q^{2n-1}e^{-2\pi iu}\right),$
we obtain
(6.26) $\begin{split}&\frac{\vartheta_{1}\left(\frac{r}{iy}+\frac{m}{2k};\frac{i}{y}\right)}{\vartheta_{1}\left(\frac{r}{iy}-\frac{m}{2k};\frac{i}{y}\right)}\\ &=\frac{\sin\left(\pi\frac{r}{iy}+\pi\frac{m}{2k}\right)}{\sin\left(\pi\frac{r}{iy}-\pi\frac{m}{2k}\right)}\exp{\left(-2i\sum_{n=1}^{\infty}\frac{1}{n}\frac{\sin(\pi n\frac{m}{k})}{1-q^{2n}}\left(q^{2n(1-r)}-q^{2n(1+r)}\right)\right)},\\ &\frac{\vartheta_{3}\left(\frac{r}{iy}+\frac{m}{2k};\frac{i}{y}\right)}{\vartheta_{3}\left(\frac{r}{iy}-\frac{m}{2k};\frac{i}{y}\right)}\\ &=\exp{\left(-2i\sum_{n=1}^{\infty}\frac{(-1)^{n}}{n}\frac{\sin(\pi n\frac{m}{k})}{1-q^{2n}}\left(q^{n(1-2r)}-q^{n(1+2r)}\right)\right)}.\end{split}$
Furthermore, from
[…]
we obtain
(6.28) $\lim_{y\to+0}\frac{\vartheta_{3}\left(\frac{r}{iy}+\frac{m}{2k};\frac{i}{y}\right)}{\vartheta_{3}\left(\frac{r}{iy}-\frac{m}{2k};\frac{i}{y}\right)}=1.$
Similarly, we obtain
(6.29) $\lim_{y\to+0}\frac{\vartheta_{1}\left(\frac{r}{iy}+\frac{m}{2k};\frac{i}{y}\right)}{\vartheta_{1}\left(\frac{r}{iy}-\frac{m}{2k};\frac{i}{y}\right)}=\exp{\left(\pi i\frac{m}{k}\right)}.$
From the above, we have
(6.30) $\lim_{y\to+0}\Psi(v,r(y),y)=2\pi m.$
From the pseudoperiod of elliptic theta function, we have
(6.31) $\Psi(v,r(y_{0}),y_{0})=3\pi m.$
Summarizing the above discussion, we have
(6.32) $\left|\Psi(v,r(y_{0}),y_{0})-\lim_{y\to+0}\Psi(v,r(y),y)\right|=\pi m.$
Therefore, when $m\geq 3$, there is always a solution in $\Lambda(v,r(y),y)=0$.
6.5. Transformations of a parameter and a variable
We show that the set of our explicit solutions admits two types of self-bijections that transform
  • $m\rightarrow m+k$, $n\rightarrow k-n$,
  • $m\rightarrow-m$, $n\rightarrow n$.
A finite number of combinations of these transformations can make any $m\in\mathbb{Z}\setminus(k\mathbb{Z})$ fall in the interval $0<m\leq k/2$. Combining with the previous subsection, this proves the existence of Kaleidocycles for $k\geq 6$.
In the following, we will construct the self-bijections and prove that the Kaleidocycles obtained by the self-bijections are essentially identical to the Kaleidocycles before the transformation. A combination of the two bijection exchanges a solution of the semi-discrete potential mKdV equation with one of the semi-discrete sine-Gordon equation, and the results are summarised in Table 1.
6.5.1. Transformation $m\rightarrow m+k$, $n\rightarrow k-n$
Consider transformation $v\rightarrow v+y$. Then, $m$ changes to $k+m$. The period of elliptic theta function $\varphi(m/k+1,y)=-\varphi(m/k,y)$ and (6.21) implies the change $r\rightarrow-r-1/2$. From the period of elliptic theta function, we see that $\Lambda(v+y,-r-1/2,y)=\Lambda(v,r,y)^{*}$ holds. Thus, this transformation preserves the closure conditions. From the original binormal vectors $\widetilde{B}_{n}=\widetilde{B}_{n}(t;v,r,y)$, the new binormal vectors are computed as
(6.33) $\begin{split}(-1)^{k-n}\widetilde{B}_{k-n}\left(t-y/2;v,r,y\right).\end{split}$
That is, the new Kaleidocycle is obtained by shifting the time parameter, reversing the orientation of the central curve, and flipping every other hinge, which is equivalently encoded by the torsion angle update $\lambda\mapsto\pi-\lambda$.
6.5.2. Transformation $m\rightarrow-m$, $n\rightarrow n$
Consider the transformation $v\rightarrow-v$ and $r\rightarrow-r$. Then, $m$ changes to m-m. From the period of elliptic theta function, we see that $\Lambda(-v,-r,y)=\Lambda(v,r,y)$ holds. Thus, this transformation preserves the closure conditions. From the original binormal vectors $\widetilde{B}_{n}=\widetilde{B}_{n}(t;v,r,y)$, the new binormal vectors are computed as
(6.34) $\begin{split}-\widetilde{B}_{n}\left(-t;v,r,y\right).\end{split}$
That is, the new Kaleidocycle is obtained by reversing the time parameter and flipping all hinges.
Summarising the results of this section, we have the following theorem.
Theorem 6.1.
For a fixed $k\geq 6$, let $0<m\leq k/2$. Then there exist $v\in\mathbb{R}\setminus{(y\mathbb{Z})}$, $r\in\mathbb{R}\setminus{(1/2\mathbb{Z})}$ and $y>0$ satisfying $kv=my$, (6.6) and (6.10). Then, the curve defined by Theorem 5.1 is closed for every $t\in\mathbb{R}$ and forms a $k$-Kaleidocycle with the torsion angle given by (5.2).
Remark 6.2.
When a curve is deformed according to (3.24), a combination of the two transformations described above exchanges the deformation by the semi-discrete potential mKdV equation and the deformation by the semi-discrete sine-Gordon equation. Consider the transformation $v\rightarrow-v-y$, $r\rightarrow r+1/2$ and $n\rightarrow k-n$. Then the new $\Gamma$ and $C$ are computed as
(6.35) $\begin{split}&\Gamma_{\pm}\left(r+1/2,y\right)=\Gamma_{\mp}\left(r,y\right),\\ &C_{\pm}\left(-v-y,r+1/2,y\right)=C_{\mp}\left(v,r,y\right).\end{split}$
The new potential function is computed as
(6.36) $(-1)^{k-n+1}\Theta_{k-n}\left(-t+y/2;v,y\right).$
The coefficients of the semi-discrete potential mKdV equation and of the semi-discrete sine-Gordon equation are transformed as
(6.37) $\left(1+\cos\lambda\right)\rho_{+}\rightarrow-\left(1-\cos\lambda\right)\rho_{-},\quad\left(1-\cos\lambda\right)\rho_{-}\rightarrow-\left(1+\cos\lambda\right)\rho_{+}.$
The transformation $\Theta_{n}\rightarrow\widetilde{\Theta}_{n}:=(-1)^{k-n+1}\Theta_{k-n}$ exchanges a solution of the semi-discrete potential mKdV equation (5.183) with one of the semi-discrete sine-Gordon equation (5.184). We see that this equation-level transformation is also achieved at the solution level by transforming parameters and a variable. This is also consistent with the result of the transformation of $\Gamma$ and $C$.
Table 1 summarises the relationships among the three deformation equations.
Table 1. Relation between the three deformation equations
Deformation equation of the curve Equation for the potential function
$C,\Gamma\in\mathbb{R}$
$\dot{\gamma}_{n}=\left(\begin{array}[]{c}0\\ 0\\ 2\eta_{1}\left(\beta_{1}-2\beta_{3}\right)-\beta_{6}-C\end{array}\right)$
$2\beta_{1}\left(\eta_{1}T_{n}-\eta_{1}\tan\frac{\kappa_{n}}{2}\widetilde{N}_{n}+\eta_{2}\tan\frac{\kappa_{n}}{2}\widetilde{B}_{n}\right)$
$\dot{\Theta}_{n}=4\beta_{1}\left(\eta_{1}\cos\lambda+\eta_{2}\sin\lambda\right)$
$\times\tan\left(\frac{\Theta_{n+1}-\Theta_{n-1}}{4}\right)$
$\downarrow$ Specialise the deformation equation by choosing
rigid transformation parameters $C$ and $\Gamma$ as in Theorem 5.11.
$C_{+},\Gamma_{+}$ $\left\langle\dot{\gamma}_{n},\widetilde{B}_{n}\right\rangle=0,\quad\langle\dot{\gamma}_{n},\dot{\gamma}_{n}\rangle=\rho^{2}_{+}$
$\frac{\dot{\Theta}_{n+1}+\dot{\Theta}_{n}}{2}=\left(1+\cos\lambda\right)\rho_{+}$
$\times\sin\left(\frac{\Theta_{n+1}-\Theta_{n}}{2}\right)$
$\updownarrow$ As in Remark 6.2, these equations exchange with
the transformation of parameters $v,r$ and a variable $n$ of the solution.
$C_{-},\Gamma_{-}$ $\left\langle\dot{\gamma}_{n},\widetilde{B}_{n}\right\rangle=0,\quad\langle\dot{\gamma}_{n},\dot{\gamma}_{n}\rangle=\rho^{2}_{-}$
$\frac{\dot{\Theta}_{n+1}-\dot{\Theta}_{n}}{2}=\left(1-\cos\lambda\right)\rho_{-}$
$\times\sin\left(\frac{\Theta_{n+1}+\Theta_{n}}{2}\right)$
Example 6.3.
The classical 6-Kaleidocycle (Figure 1 Left) with $\lambda=\pi/2$ (e.g., see [9]) is obtained by $m=3$ and $(v,r,y)\approx(0.6400,-0.2500,1.279)$.
Remark 6.4.
By using (6.6) and (A.12), we see that $\widetilde{B}_{0}=(-1)^{m}\widetilde{B}_{k}$. That is, the parity of $m$ corresponds to the orientability of the Kaleidocycle.
7. Examples
In this section, we present examples of Kaleidocycles constructed by Theorem 6.1 and the semi-discrete K-surfaces obtained as their motion trajectory as explained in Remark 5.12.
Numerical experiments using our computer codes [19] suggest that the $8$-, 99- and $15$-Kaleidocycles in Figs. 2, 3, and 4 have the minimum torsion angles among non-orientable $8$-, 99-, and $15$-Kaleidocycles, respectively [20]. In other words, they are the Möbius Kaleidocycles [22] (see Remark 2.1).
For Fig. 5, the torsion angle is larger than the above minimum one. However, it has a self-linking number $7.5$. As the self-linking number of a Kaleidocycle is a half-integer, these $15$-Kaleidocycles depicted in Figs. 4 and 5 belong to different connected components of the configuration space of $15$-Kaleidocycles. We numerically find that the latter attains the local minima of the torsion angle. We present the following conjecture based on our extensive numerical experiments:
Conjecture 7.1.
Any Kaleidocycle obtained by Theorem 6.1 has a minimal or maximal torsion angle. In other words, the projection from the space defined by (1.1) onto the $\lambda$-coordinate attains a critical value.
As briefly discussed in §1 and Remark 2.1, these Kaleidocycles with critical torsion angles numerically exhibit special properties, including the single degree of freedom [20, 22, 31]. Our construction produces this interesting family of Kaleidocycles and may shed light on the rigorous proof for some of their unique properties.
Figure 2. (Left) An example of 8-Kaleidocycle with $(v,r,y)\approx(0.4000,-0.2834,1.067)$ having $m=3$ and $\cos\lambda\approx 0.4700$. (Middle) 8-Kaleidocycle as a discrete curve. (Right) the semi-discrete K-surface obtained as its trajectory.
Figure 3. (Left) An example of 9-Kaleidocycle with $(v,r,y)\approx(0.3422,-0.2890,1.027)$ having $m=3$ and $\cos\lambda\approx 0.5852$. (Middle) 9-Kaleidocycle as a discrete curve. (Right) the semi-discrete K-surface obtained as its trajectory.
Figure 4. (Left) An example of 15-Kaleidocycle with $(v,r,y)\approx(0.1894,-0.2996,0.9470)$ having $m=3$ and $\cos\lambda\approx 0.8533$. (Middle) 15-Kaleidocycle as a discrete curve. (Right) the semi-discrete K-surface obtained as its trajectory.
Figure 5. (Left) An example of 15-Kaleidocycle with $(v,r,y)\approx(0.4012,-0.3113,1.204)$ having $m=5$ and $\cos\lambda\approx 0.6497$. (Middle) 15-Kaleidocycle as a discrete curve. (Right) the semi-discrete K-surface obtained as its trajectory.
Appendix A Functional identities of elliptic theta functions
In the appendix, we collect some properties of elliptic theta functions which we rely heavily on our calculations.
A.1. Four term identities
Let
(A.1) $\begin{split}&X_{1}=\frac{1}{2}(X+Y+U+V),\quad Y_{1}=\frac{1}{2}(X+Y-U-V),\\ &U_{1}=\frac{1}{2}(X-Y+U-V),\quad V_{1}=\frac{1}{2}(X-Y-U+V),\end{split}$
then for any $X,Y,U,V\in\mathbb{C}$, elliptic theta functions satisfy the following identities [28, 24].
(A.2) $\vartheta_{2}(X_{1})\vartheta_{2}(Y_{1})\vartheta_{3}(U_{1})\vartheta_{3}(V_{1})-\vartheta_{4}(X_{1})\vartheta_{4}(Y_{1})\vartheta_{1}(U_{1})\vartheta_{1}(V_{1})$
$=\vartheta_{2}(X)\vartheta_{2}(Y)\vartheta_{3}(U)\vartheta_{3}(V)-\vartheta_{4}(X)\vartheta_{4}(Y)\vartheta_{1}(U)\vartheta_{1}(V),$
(A.3) $\vartheta_{4}(X_{1})\vartheta_{3}(Y_{1})\vartheta_{1}(U_{1})\vartheta_{2}(V_{1})-\vartheta_{2}(X_{1})\vartheta_{1}(Y_{1})\vartheta_{3}(U_{1})\vartheta_{4}(V_{1})$
$=\vartheta_{4}(X)\vartheta_{3}(Y)\vartheta_{1}(U)\vartheta_{2}(V)-\vartheta_{2}(X)\vartheta_{1}(Y)\vartheta_{3}(U)\vartheta_{4}(V),$
(A.4) $\vartheta_{2}(X_{1})\vartheta_{2}(Y_{1})\vartheta_{4}(U_{1})\vartheta_{4}(V_{1})+\vartheta_{1}(X_{1})\vartheta_{1}(Y_{1})\vartheta_{3}(U_{1})\vartheta_{3}(V_{1})$
$=\vartheta_{4}(X)\vartheta_{4}(Y)\vartheta_{2}(U)\vartheta_{2}(V)-\vartheta_{3}(X)\vartheta_{3}(Y)\vartheta_{1}(U)\vartheta_{1}(V),$
(A.5) $\vartheta_{3}(X_{1})\vartheta_{3}(Y_{1})\vartheta_{2}(U_{1})\vartheta_{2}(V_{1})+\vartheta_{1}(X_{1})\vartheta_{1}(Y_{1})\vartheta_{4}(U_{1})\vartheta_{4}(V_{1})$
$=\vartheta_{3}(X)\vartheta_{3}(Y)\vartheta_{2}(U)\vartheta_{2}(V)+\vartheta_{1}(X)\vartheta_{1}(Y)\vartheta_{4}(U)\vartheta_{4}(V),$
(A.6) $\vartheta_{4}(X_{1})\vartheta_{4}(Y_{1})\vartheta_{4}(U_{1})\vartheta_{4}(V_{1})-\vartheta_{1}(X_{1})\vartheta_{1}(Y_{1})\vartheta_{1}(U_{1})\vartheta_{1}(V_{1})$
$=\vartheta_{4}(X)\vartheta_{4}(Y)\vartheta_{4}(U)\vartheta_{4}(V)-\vartheta_{1}(X)\vartheta_{1}(Y)\vartheta_{1}(U)\vartheta_{1}(V),$
(A.7) $\vartheta_{4}(X_{1})\vartheta_{4}(Y_{1})\vartheta_{1}(U_{1})\vartheta_{1}(V_{1})-\vartheta_{1}(X_{1})\vartheta_{1}(Y_{1})\vartheta_{4}(U_{1})\vartheta_{4}(V_{1})$
$=\vartheta_{4}(X)\vartheta_{4}(Y)\vartheta_{1}(U)\vartheta_{1}(V)-\vartheta_{1}(X)\vartheta_{1}(Y)\vartheta_{4}(U)\vartheta_{4}(V),$
(A.8) $\vartheta_{4}(X_{1})\vartheta_{4}(Y_{1})\vartheta_{1}(U_{1})\vartheta_{1}(V_{1})+\vartheta_{2}(X_{1})\vartheta_{2}(Y_{1})\vartheta_{3}(U_{1})\vartheta_{3}(V_{1})$
$=\vartheta_{3}(X)\vartheta_{3}(Y)\vartheta_{2}(U)\vartheta_{2}(V)-\vartheta_{1}(X)\vartheta_{1}(Y)\vartheta_{4}(U)\vartheta_{4}(V),$
(A.9) $\vartheta_{3}(X_{1})\vartheta_{3}(Y_{1})\vartheta_{1}(U_{1})\vartheta_{1}(V_{1})-\vartheta_{1}(X_{1})\vartheta_{1}(Y_{1})\vartheta_{3}(U_{1})\vartheta_{3}(V_{1})$
$=\vartheta_{3}(X)\vartheta_{3}(Y)\vartheta_{1}(U)\vartheta_{1}(V)-\vartheta_{1}(X)\vartheta_{1}(Y)\vartheta_{3}(U)\vartheta_{3}(V),$
(A.10) $\vartheta_{3}(X_{1})\vartheta_{3}(Y_{1})\vartheta_{4}(U_{1})\vartheta_{4}(V_{1})-\vartheta_{1}(X_{1})\vartheta_{1}(Y_{1})\vartheta_{2}(U_{1})\vartheta_{2}(V_{1})$
$=\vartheta_{3}(X)\vartheta_{3}(Y)\vartheta_{4}(U)\vartheta_{4}(V)-\vartheta_{1}(X)\vartheta_{1}(Y)\vartheta_{2}(U)\vartheta_{2}(V).$
A.2. Shifts by periods and half-periods
For any $X\in\mathbb{C}$, $n\in\mathbb{Z}$ and $y>0$, the following holds [28, 24].
(A.11) $\vartheta_{1}\left(X+n\right)=(-1)^{n}\vartheta_{1}\left(X\right),\quad\vartheta_{4}\left(X+n\right)=\vartheta_{4}\left(X\right).$
(A.12) $\begin{split}&\vartheta_{1}\left(X+niy;iy\right)=(-1)^{n}\exp\{-\pi i\left(2nX+n^{2}iy\right)\}\vartheta_{1}\left(X;iy\right),\\ &\vartheta_{2}\left(X+niy;iy\right)=\exp\{-\pi i\left(2nX+n^{2}iy\right)\}\vartheta_{2}\left(X;iy\right),\\ &\vartheta_{3}\left(X+niy;iy\right)=\exp\{-\pi i\left(2nX+n^{2}iy\right)\}\vartheta_{3}\left(X;iy\right),\\ &\vartheta_{4}\left(X+niy;iy\right)=(-1)^{n}\exp\{-\pi i\left(2nX+n^{2}iy\right)\}\vartheta_{4}\left(X;iy\right).\end{split}$
(A.13) $\begin{split}&\vartheta_{1}\left(X+\frac{1}{2}\right)=\vartheta_{2}\left(X\right),\quad\vartheta_{2}\left(X+\frac{1}{2}\right)=-\vartheta_{1}\left(X\right),\\ &\vartheta_{3}\left(X+\frac{1}{2}\right)=\vartheta_{4}\left(X\right),\quad\vartheta_{4}\left(X+\frac{1}{2}\right)=\vartheta_{3}\left(X\right).\end{split}$
(A.14) $\begin{split}&\vartheta_{1}\left(X+\frac{iy}{2};iy\right)=i\exp\left\{-\pi i\left(X+\frac{iy}{4}\right)\right\}\vartheta_{4}\left(X;iy\right),\\ &\vartheta_{2}\left(X+\frac{iy}{2};iy\right)=\exp\left\{-\pi i\left(X+\frac{iy}{4}\right)\right\}\vartheta_{3}\left(X;iy\right),\\ &\vartheta_{3}\left(X+\frac{iy}{2};iy\right)=\exp\left\{-\pi i\left(X+\frac{iy}{4}\right)\right\}\vartheta_{2}\left(X;iy\right),\\ &\vartheta_{4}\left(X+\frac{iy}{2};iy\right)=i\exp\left\{-\pi i\left(X+\frac{iy}{4}\right)\right\}\vartheta_{1}\left(X;iy\right).\end{split}$
A.3. Identity for theta constants
For any $y>0$, elliptic theta functions satisfy the following identity [28, 24].
(A.15) $\vartheta_{1}^{{}^{\prime}}\left(0;iy\right)=\pi\vartheta_{2}\left(0;iy\right)\vartheta_{3}\left(0;iy\right)\vartheta_{4}\left(0;iy\right).$
A.4. Modular transformations
For any $X\in\mathbb{C}$ and $y>0$, elliptic theta functions satisfy the following identities [28, 24].
(A.16) $\begin{split}&\vartheta_{1}\left(\frac{X}{y};i\frac{2}{y}\right)=-i\sqrt{\frac{y}{2}}\exp\left(-\pi\frac{X^{2}}{2y}\right)\vartheta_{1}\left(\frac{iX}{2};i\frac{y}{2}\right),\\ &\vartheta_{2}\left(\frac{X}{y};i\frac{2}{y}\right)=\sqrt{\frac{y}{2}}\exp\left(-\pi\frac{X^{2}}{2y}\right)\vartheta_{4}\left(\frac{iX}{2};i\frac{y}{2}\right),\\ &\vartheta_{3}\left(\frac{X}{y};i\frac{2}{y}\right)=\sqrt{\frac{y}{2}}\exp\left(-\pi\frac{X^{2}}{2y}\right)\vartheta_{3}\left(\frac{iX}{2};i\frac{y}{2}\right),\\ &\vartheta_{4}\left(\frac{X}{y};i\frac{2}{y}\right)=\sqrt{\frac{y}{2}}\exp\left(-\pi\frac{X^{2}}{2y}\right)\vartheta_{2}\left(\frac{iX}{2};i\frac{y}{2}\right).\end{split}$
(A.17) $\frac{\vartheta_{1}\left(X;iy\right)}{\vartheta_{2}\left(X;iy\right)}=-i\frac{\vartheta_{1}\left(\frac{iX}{y};\frac{i}{y}\right)}{\vartheta_{4}\left(\frac{iX}{y};\frac{i}{y}\right)}.$
A.5. Bilinear identities connecting theta functions with $y$ and $y/2$
For any $X,Y\in\mathbb{C}$, and $y>0$, elliptic theta functions satisfy the following identities [28, 24].
$2\vartheta_{1}\left(X+Y;iy\right)\vartheta_{4}\left(X-Y;iy\right)$
(A.18) $=\vartheta_{1}\left(X;\frac{iy}{2}\right)\vartheta_{2}\left(Y;\frac{iy}{2}\right)+\vartheta_{2}\left(X;\frac{iy}{2}\right)\vartheta_{1}\left(Y;\frac{iy}{2}\right),$
$2\vartheta_{2}\left(X+Y;iy\right)\vartheta_{3}\left(X-Y;iy\right)$
(A.19) $=\vartheta_{2}\left(X;\frac{iy}{2}\right)\vartheta_{2}\left(Y;\frac{iy}{2}\right)-\vartheta_{1}\left(X;\frac{iy}{2}\right)\vartheta_{1}\left(Y;\frac{iy}{2}\right).$
[…]